A Complete Bibliography of Journal of the Royal Statistical Society, Series D (The Statistician): 1990–1999

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

             E-mail: beebe@math.utah.edu, beebe@acm.org,
                       beebe@computer.org (Internet)
           WWW URL: http://www.math.utah.edu/~beebe/

13 October 2017 Version 1.01

Title word cross-reference

2 [814, 677]. 24 [59, 58]. 2^{3} [886]. 2 × 2 [643]. 48 [59]. X̄_{Q} [210]. C(α) [4, 334]. C_{p} [669]. χ^{2} [643, 467, 434]. F [674]. k [672]. L [1029]. P [641, 742, 374, 494]. P (X_{2} < X_{1}) [774]. R_{Q} [210]. t [1002, 104, 674, 1156]. τ [1085]. U [635]. z [309].

- [643, 674]. -Error [1002]. -Factorials [886]. -h [58]. -Moments [1029]. -Nearest-Neighbour [672]. -Statistic [1156, 669]. -Statistics [635]. -Test [434]. -Tests [674]. -Transformation [309]. -Value [494]. -Values [374, 742]. -Water [677].

100 [556]. 1750 [248]. 1900 [1169]. 1984 [700]. 1988 [486]. 1990s [846].

2 [1111, 1101]. 21st [583].

3 [1024].

5 [932].

Abelson [832]. Abortion [131, 132]. ABPM [59]. Absorption [1052]. Abstracts [575, 668, 745]. Abuse [397]. Accident [3, 818]. Accidental [409]. Accountants [685]. Accounts [98].

Accuracy [425, 619, 364]. Achieved [540]. Achievement [376]. Achieving [678]. Acknowledgement [751]. Action [36]. Active [244]. Activities [954]. Activity [841]. Actualité [1035]. Actuarial [743, 748]. Actuaries [603]. Adaptive [1137, 433, 667, 403]. Address [373, 34, 593, 37]. Address-Based [373]. Adhikari [318]. Adil [191]. Adjusted [882]. Adjustment [647, 125, 1095, 494]. Admission [369]. Advanced [459, 81, 896]. Advances [735, 968, 1071, 1181, 247, 1030]. Advantage [649, 855, 1087]. Advantages [354]. Advertisements [365]. Advice [1122]. Affected [704]. Aflatoxin [469]. Africa [824]. After [369, 409, 1047, 1081]. Against [461, 306]. Age [8]. Age-Prevalence [8]. Ageing [46, 47]. Agglomeration [103]. Aggregate [291]. Agreement [1121, 1158]. Agresti [778, 183]. Agriculture [524]. Ahead [811]. AHP [259]. AI [479]. AIC [529]. Aid [203]. AIDS [174]. Aike [508]. Air [51]. Ajit [26]. Alan [148, 1091]. Alexander [1010]. Alfredo [247]. Algebra [1148]. Algorithm [276, 826, 988, 719]. Algorithms [272]. All-Equity [1123]. Allan [550]. Allison [972]. Allocation [830, 775, 597]. Alternative [466, 544]. Althoen [680, 680]. Among [650, 1127]. Analyse [265]. Analyses [396, 632, 979]. Analysing [470, 504, 736]. Analysis [293, 184, 1096, 121, 137, 1122, 665, 412, 880, 73, 434, 926, 130, 953, 909, 839, 1045, 963, 209, 607, 435, 860, 512, 541, 813, 117, 490, 395, 3, 560, 1076, 969, 974, 904, 964, 580, 659, 423, 861, 1168, 981, 1018, 442, 648, 962, 257, 127, 929, 683, 150, 825, 561, 399, 763, 370, 319, 264, 1149, 332, 995, 529, 606, 162, 1050, 728, 886, 139, 109, 366, 147, 322, 284, 810, 1097, 952, 950, 989, 1174, 1029, 252, 180, 279, 795, 510, 558, 923, 993, 375, 428, 8, 503, 703, 285, 814, 579, 468, 1055]. Analysis [2, 542, 376, 334, 310, 958, 126, 1060, 746, 770, 421, 132, 961, 317, 474, 576, 1171, 168, 429, 905, 972, 487, 1020, 1066, 1073, 711, 218, 381, 186, 760, 778, 837, 1057, 735, 19, 477, 1098, 1146, 455, 759, 1152, 183, 495, 250, 688, 1054, 116, 710, 1067, 692, 835, 1072, 559, 653, 1059, 658, 454, 249]. Analysis-Challenges [257]. Analysis-of-Variance [770]. Analytic [527]. Analyzing [161, 339]. Ancient [372]. Andersen [727, 929]. Anderson [728]. Andrew [351]. Andrich [118]. Angles [97]. Anguish [1055]. Annealing [598]. Annette [159]. Announcement [418]. ANOVA [902]. Ansell [435]. Answering [1114]. Answers [1014]. Antihypertensive [60]. Applebaum [888, 930]. Applicability [596]. Application [309, 850, 879, 394, 212, 379, 769, 265, 365, 292, 1144, 722, 671, 139, 366, 116, 383, 579, 402, 680]. Applications [155, 910, 23, 125, 147, 989, 876, 569, 248, 1148, 82, 885, 115, 1147, 182, 680, 514, 1025, 971, 87, 523, 533, 1071, 1136, 1135, 951, 440, 80, 789, 215, 1165, 1063, 973, 352, 942, 239, 1179, 156, 510, 563, 1012, 221, 712, 1074, 680, 909]. Applied [575, 668, 745, 1040, 840, 572, 160, 348, 1054, 902, 458, 381, 347, 638, 969, 904, 185, 759, 561, 177, 1149, 688, 1067, 1174, 1026, 970, 1059, 1060]. Appliquée [513]. Applying [984]. Appraisal [361]. Approach [590, 665, 1121, 73, 924, 95, 969, 1011, 259, 423, 800, 1144, 466, 101, 369, 1029, 232, 410, 833, 570, 420, 701, 398, 758, 560, 222, 687, 518, 652, 160, 217, 158, 108, 930]. Approaches [280, 394, 1143]. Appropriate [919, 921, 468]. Approximating [827]. Approximation [1156, 511]. Aptitudes [66]. APU [142]. Arbitrariness [171]. ARCH [1025]. Archaeological [833]. Archaeostatistics [372]. Area [541]. Areas [32, 31, 42, 69, 41, 33]. Argument [832]. Arithmetic [1141]. ARMA [439]. Armitage [968]. Armstrong [1134]. Arnold [1163, 473]. Arrays [627]. Art [1038, 558]. Arterial [56]. Artificial [479].

ASP [680]. Aspects [1164, 225, 996, 911]. Assess [366]. Assessing [365, 672, 337]. Assessment [435, 32, 414, 1028, 931, 60]. Assessments [430, 25]. Assisted [585]. Association [963, 964, 594, 102, 967, 894]. Assurance [498]. Astronomical [405]. Astronomy [909, 327]. Asymmetric [742, 284]. Asymmetry [622, 694]. Asymptotic [523, 1155, 15, 584, 976, 611]. Asymptotically [821]. Asymptotics [898, 600, 631]. Athletic [707, 966]. Atmospheric [409]. Attitudes [131, 132]. Attribute [507]. Attributes [362]. Attribution [767]. Attrition [700, 145]. Audiences [197]. Audit [202, 191]. Auditing [427]. Audits [1004]. Australia [708]. Australians [708]. Authorship [767]. Autocorrelated [295, 2]. Autocovariance [847]. Automated [569, 597]. Automatic [922]. Autoregressive [622, 694, 624]. Availability [637]. Average [148, 501]. Avoiding [114]. Ayton [693]. Azzalini [969, 856].

B [1040, 439, 732, 909, 486, 711, 516, 607, 436, 381, 780, 523, 730, 785, 240, 842, 659, 840, 867, 1111, 1135, 1018, 442, 342, 477, 1146, 929, 554, 160, 912, 1163, 551, 632, 162, 838, 441, 1099, 1117, 217, 237, 473, 1093, 585, 241, 220, 1072, 488, 1140, 1148, 1101, 970, 859, 1142, 731, 474, 691, 972]. Baba [1103]. Babyfood [579]. Back [29, 48, 63, 91, 120, 134, 164, 198, 277, 223, 254, 287, 301, 390, 406, 416, 462, 328, 358, 496, 535, 566, 587, 616, 639, 662, 695, 715, 739, 765, 791, 816, 843, 872, 914, 955, 997, 1042, 1079, 1118, 1153, 1183, 398]. Back-Calculation [398]. Bain [195]. Bakeman [652]. Balakrishnan [1163, 473, 657, 1071, 1030]. Balanced [800]. Baldev [380]. Balls [1086]. Baltagi [1093]. Bank [365]. Banks [1111]. Bapat [1135]. Barlow [549]. Barnard [315]. Barndorff [600, 15, 631]. Barndorff-Nielsen [600, 15, 631]. Barnett [1164, 356, 931, 677, 84]. Barra [343]. Bartell [217]. Bartholomew [1011]. Bartle [752]. Bartlett [885]. Bartlett-Type [885]. Based [1121, 373, 924, 1125, 780, 727, 742, 841, 315, 768, 546, 796, 1029, 753, 856, 420, 465, 494]. Bases [513]. Basic [867, 537, 245, 476, 243, 450, 1134, 787]. Basics [983, 902, 1069]. Basin [683]. Basis [62]. Basu [1165]. Bayes [27, 351, 75, 665, 401, 800, 774, 413, 628, 604]. Bayesian [1020, 937, 865, 1021, 90, 688, 604, 108, 182, 280, 330, 590, 396, 228, 1004, 412, 747, 880, 434, 394, 499, 395, 3, 423, 432, 272, 397, 411, 564, 592, 803, 399, 763, 403, 101, 748, 922, 279, 591, 795, 415, 410, 818, 409, 427, 719, 801, 398, 402, 802, 429, 422, 711, 1022, 12, 838, 13, 994, 833, 932]. be [1001, 921, 206]. Beaumont [857]. Beautiful [142]. Beaver [456]. Before [708, 248]. Beginner [947, 1024]. Behalf [34, 37]. Behavioral [236, 555]. Behaviour [594, 1083]. Behavioural [323, 1113]. Behaviours [850]. Behrens [670, 595]. Being [35, 44]. Bejar [552]. Belfast [42]. Beliefs [419, 397, 876]. Belsley [653]. Bench [1004]. Bench-Mark [1004]. Bendat [1136, 322]. Bennett [581, 509, 319]. Beran [1106]. Berger [932, 590]. Bergman [242]. Bernardo [13, 932]. Bernoulli [431, 226, 770]. Berry [1112]. Bertin [889]. Best [651]. Beta [825, 168]. Betteley [654]. Better [583]. Betting [1130, 1131, 371, 471]. Betty [88]. Between [293, 1002, 708, 803, 472, 903, 804, 1121, 1052]. Beveridge [999, 1000]. Beyond [902]. Bhattacharjee [1165]. Bhattacharya [239]. Bhattacharyya [938, 451]. Bhatti [753]. Bias [593, 327, 304]. Bibliography [105]. Bickel [1137]. Big [142]. Bijaoui [1152]. Bilingual [684]. Billard [453]. Billington [550]. Bills [364]. Binary [1049, 335, 423, 313, 168]. Binomial [280, 717, 75, 71, 297, 589, 819]. Binormal [379, 769]. Bio [901]. Bio/ [901]. Bioequivalence [28]. Biography [315]. Biological [43, 235, 1126]. Biology [815, 149, 524, 241, 1094]. Biometric [968]. Biometry [968, 211]. Biostatistics [262, 780, 901, 66, 787]. Biplots [863]. Birth [967]. Bissell [601]. Biswas [388]. Bivariate [823, 774, 215, 268, 1003, 384, 977]. Biweight [170, 307]. Bladder [265]. Blocks [290]. Blood [58]. Bock [1103, 510]. Bolanowski [324]. Bole [291]. Bonferroni [973]. Bonferroni-Type [973]. Book [581, 808, 184, 1096, 1141, 444, 353, 1064, 891, 1040, 1010, 439, 1039, 815, 487, 492, 732, 895, 900, 866, 1020, 1164, 985, 1066, 483, 752, 242, 940, 1019, 1112, 608, 514, 148, 157, 953, 1034, 1105, 1109, 909, 1025, 1073, 152, 992, 839, 937, 486, 534, 387, 437, 458, 615, 679, 686, 711, 781, 971, 982, 325, 484, 636, 655, 526, 733, 516, 87, 489, 790, 892, 1023, 1095, 509, 607, 110, 118, 435, 436, 11, 382, 657, 689, 758, 860, 218, 512, 381, 913, 88, 186, 345, 936, 1056, 780, 727, 813, 865]. Book [939, 1022, 1145, 600, 117, 24, 347, 490, 945, 15, 523, 614, 730, 528, 533, 638, 738, 557, 785, 933, 114, 240, 560, 658, 1075, 1076, 27, 969, 1150, 907, 974, 979, 1011, 190, 678, 1070, 904, 163, 343, 968, 1177, 760, 778, 788, 811, 841, 842, 1180, 1021, 493, 556, 610, 612, 659, 222, 236, 840, 861, 1168, 1172, 149, 693, 899, 1071, 836, 837, 867, 897, 943, 980, 1065, 1100, 1106, 1111, 1136, 1137, 191, 346, 981, 633, 1135, 687, 518, 521, 25, 456, 457, 584, 635, 681, 951, 1057, 652, 735, 894, 1018]. Book [159, 238, 442, 1028, 342, 713, 315, 530, 19, 782, 219, 447, 477, 517, 552, 555, 870, 946, 1104, 888, 1098, 151, 1107, 479, 84, 243, 17, 520, 1146, 12, 90, 929, 1144, 631, 1181, 947, 16, 554, 683, 107, 160, 194, 440, 455, 313, 321, 634, 912, 150, 155, 185, 323, 450, 80, 901, 348, 759, 320, 561, 714, 789, 356, 438, 86, 154, 177, 215, 459, 319, 941, 1163, 1165, 1178, 20, 906, 910, 1063, 23, 452, 551, 632, 893, 1015, 1062, 1149, 83, 973, 995, 1016, 529, 245, 511, 524, 606, 976, 1103, 1152, 1058, 1036, 522, 1061]. Book [162, 316, 478, 898, 187, 247, 253, 183, 454, 761, 812, 838, 868, 728, 784, 26, 119, 1077, 1170, 441, 519, 79, 249, 495, 1068, 214, 246, 385, 352, 532, 445, 354, 942, 983, 1099, 1117, 485, 448, 179, 217, 22, 109, 192, 783, 834, 944, 934, 216, 250, 349, 147, 237, 153, 239, 322, 527, 562, 688, 757, 949, 1108, 810, 938, 1054, 1097, 1179, 116, 996, 1175, 948, 710, 1067, 952, 451, 460, 950, 1151, 978, 989, 14, 111, 386, 473, 604, 1174, 1029, 314, 903, 480, 692, 835, 755, 113, 682, 1093, 81]. Book [252, 1026, 585, 13, 161, 312, 994, 180, 241, 453, 637, 340, 156, 193, 220, 508, 510, 734, 558, 1072, 389, 251, 85, 344, 753, 935, 1115, 987, 188, 488, 993, 1024, 1110, 158, 318, 986, 196, 248, 189, 244, 890, 902, 1173, 18, 563, 684, 1176, 990, 1140, 449, 1012, 559, 446, 515, 609, 889, 908, 1030, 1148, 582, 814, 833, 858, 931, 984, 1013, 1101, 1113, 1134, 1166, 1017, 586, 737, 178, 221, 832, 970, 384, 549, 654, 380, 311, 327, 613, 10, 601, 108, 195, 656, 383, 182, 351, 685, 954, 1038, 525, 611]. Book [513, 324, 550, 653, 677, 712, 809, 1055, 1114, 491, 388, 1031, 1059, 1169, 21, 82, 89, 112, 786, 857, 859, 756, 1032, 930, 932, 1033, 1035, 1138, 475, 583, 690, 350, 482, 1069, 494, 762, 869, 988, 1142, 977, 355, 1094, 1139, 856, 1074, 779, 1060, 603, 181, 197, 911, 481, 531, 680, 341, 602, 736, 754, 863, 991, 787, 729, 864, 731, 896, 1014, 115, 443, 317, 326, 474, 476, 553, 605, 691, 862, 1037, 1167, 357, 975, 1027, 1102, 1116, 1143, 1147, 1171, 905, 972]. Bookstein [1094]. Boolean [945]. Bootstrap [502, 626, 756, 551, 453, 609, 971]. Bootstrapping [28]. Borg [1012]. Borgan [727]. Borovkov [1138]. Borovskich [635]. Bosq [890]. Bossche [637]. Both [806]. Boundaries [505]. Bounds [707, 724, 499, 202]. Bourque [436].

Bouvier [975]. Bowerman [160]. Bowman [969]. Bown [892]. Box [1095, 826, 229]. Boyle [1010]. Brand [426]. Brandt [237]. Brazil [174]. Breaking [917]. Breakthroughs [982]. Breast [721]. Breckling [147]. Breen [809]. Brenner [297]. Brewery [401]. Brief [641]. Brillinger [1096]. Briscoe [784]. Britain [350]. British [697]. Britt [891]. Brockwell [314, 858]. Bromek [387]. Brown [437, 729]. Brownian [996]. BS [498]. Buck [833]. Building [1124, 110]. Built [43]. Buja [382]. Bulmer [532]. Bunday [859]. Burden [218]. Burdick [1026]. Bureau [122]. Burman [1161]. Business [105, 456, 179, 216, 970, 718, 939]. Bust [604]. Buying [156].

C [581, 487, 1019, 1109, 1025, 1073, 839, 679, 971, 790, 689, 860, 512, 557, 760, 610, 837, 1100, 1111, 1137, 447, 1144, 450, 714, 438, 215, 1163, 1165, 893, 511, 1103, 247, 868, 784, 26, 79, 1068, 22, 239, 562, 949, 948, 1067, 473, 585, 180, 340, 156, 389, 189, 446, 1148, 833, 1166, 1017, 384, 549, 182, 491, 475, 583, 1074, 603, 863, 787, 729, 115, 862]. C.-E [585]. C.-L [1166]. Cacoullos [119]. Caerphilly [70]. Calculating [330]. Calculation [398]. Calculations [397, 589, 805, 577, 801]. Calculus [155, 1169]. Calibration [465, 422]. Cambridge [1099]. Campbell [158]. Can [1001, 419]. Cancer [396, 721, 265, 264]. Capability [747, 626, 148]. Capitalist [353]. Capture [851]. Capture-Recapture [851]. Cardiovascular [56, 62]. Cardoso [1109]. Caribbean [211]. Carlin [711]. Carlo [781, 1021, 241, 623, 879]. Carlo-Stochastic [1021]. Carlos [327]. Carlson [970]. Carroll [1054, 779]. Cartographic [964]. Case [396, 1122, 338, 370, 1015, 245, 767, 923, 1128, 469, 819, 972]. Case-Control [396, 923]. Case-Study [1122, 469]. Casebook [710]. Casella [491, 1175]. Cases [174]. Cash [72]. Castillo [326]. Catastrophe [729]. Categorical [778, 929, 162, 986, 310, 719, 529, 183, 193]. Categories [642]. Caulcutt [678]. Causal [1038]. Causality [515]. Cavanagh [833]. CBDEA [610]. cdv [964]. Cells [704]. Censored [1078, 209, 981, 308, 809]. Censoring [268, 719, 7]. Census [175]. Central [122, 226, 493]. Centre [37]. Century [251, 1031, 583]. Ch [903]. Chain [781, 879, 721, 1021, 503]. Chains [343, 946]. Challenge [353, 32]. Challenges [257, 1007, 354, 811]. Chambers [1139]. Change [404, 52, 1057, 128, 503]. Change-Point [1057]. Changing [573, 364]. Chaos [521, 729, 1102]. Characteristics [365]. Characterization [227, 488]. Chart [361]. Charts [208, 820, 768, 507, 210, 362, 148]. Chase [437, 892]. Chatfield [679, 860]. Chatterjee [710, 474]. Chebyshev [671]. Chemical [428]. Chen [438]. Cheney [511]. Cheng [1166]. Chernoff [173]. Chernoff-Type [173]. Chess [296]. Chi [935]. Chi-Squared [935]. Child [397]. Childhood [67]. China [332, 438]. Chinchilli [953]. Chinese [684]. Choi [439]. Choice [581, 798, 797, 1085, 1051, 335, 796, 260, 667, 794, 799, 468, 376]. Choosing [800]. Chow [933]. Christensen [1013, 1014]. Chromosomes [597]. Chronobiologic [60]. Chronobiological [56]. Chui [511]. Church [40]. Circular [606]. Cirrhosis [366]. Cities [36, 32, 39, 31, 42, 55, 68, 33, 54]. Citro [475]. City [38, 52, 53, 40]. Clark [436]. Clarke [861, 22, 476, 243]. Clarotti [549]. Class [866, 628]. Classe [1035]. Classes [1036]. Classification [839, 1028, 1103, 1036, 510]. Classifier [672]. Clifford [88]. Clinical [394, 393, 721, 56, 1070, 530, 838, 806, 505, 59, 746, 736, 780]. Clinimetric [537]. Clubs [649]. Cluster [173, 579]. Clustering [1010, 212, 299, 368]. Clusters [173]. Co [87]. Co-Operative [87]. Coakes [1055]. Cobb [1056, 1097]. Coefficient [76, 309, 338, 1128]. Coefficients [154]. Cognitive [581]. Cohen [88]. Cointegration [1058]. Cole [787]. Colell [1143]. Coleman [893, 862]. Colin [108]. Collected [1106]. Collection [5]. Collett [313, 582, 602]. Column [660]. Combining [414, 923]. Commentary [591, 1106, 857]. Comments [280, 590, 798, 699, 105, 592]. Committee [811]. Common [171, 4, 114]. Common-Factor [171]. Commonsense [158]. Communication [168]. Community [62]. Companion [1056]. Comparative [340]. Comparing [966, 884, 170, 307, 578]. Comparison [708, 650, 26, 472, 593, 723]. Comparisons [868, 385]. Competition [673]. Competitions [472]. Competitive [426]. Complete [1078, 724, 308]. Complex [877, 432, 1014, 478]. Complicated [1045]. Component [675, 429]. Components [538, 540, 292, 925, 320, 491]. Compound [717]. Computation [1157, 317]. Computational [815, 952, 993, 653]. Computer [924, 117, 783, 16, 455, 609, 383]. Computer-Based [924]. Computer-Interactive [455]. Computerized [683]. Computers [122]. Computing [732, 243, 382]. Comrey [477]. Concepts [1113, 712, 680, 33, 357, 978, 181, 553]. Conceptual [891]. Concerning [579, 927]. Conditional [230, 622, 694]. Conditions [171, 849]. Conference [575, 668, 745, 572]. Confidence [502, 641, 71, 724, 626, 297, 629, 1128, 829]. Confirmation [604]. Conflict [1002]. Conflicting [674]. Conjecture [1038]. Considerations [828, 596]. Consistency [571]. Consonance [1129]. Constant [667]. Constants [260]. Constituent [367]. Constrained [400, 647, 1159]. Constraints [1124]. Constructing [926, 877, 1036]. Construction [850, 58, 581, 1028].

Consumer [672, 361]. Consumption [717]. Content [101]. Context [52]. Contexts [372]. Contingency [1076, 442, 464, 305, 642]. Contingent [1051]. Continua [1180]. Continuous [823, 83, 384, 657, 215]. Contributions [438, 380, 113]. Control [396, 492, 647, 1095, 208, 867, 769, 820, 768, 83, 403, 507, 361, 366, 153, 113, 923, 210, 596, 362, 934, 931]. Controlled [706]. Controlling [806]. Controversies [357]. Convergence [910]. Convolution [769]. Cook [1167]. Cookbook [901]. Cooke [243, 476]. Cooked [1122]. Coombes [1140]. Cooperation [1143]. Coping [66]. Corbelled [404]. Cornell [979]. Coronary [61, 576]. Correction [461, 564, 660, 763, 694]. Correlated [6]. Correlation [76, 283, 309, 518, 154, 4, 1003]. Correspondence [726, 630, 78, 176, 127, 319, 831, 777, 1132, 709, 548, 213, 676]. Corrigendum [1078, 300, 28]. Corter [894]. Cost [722]. Costs [51, 163]. Could [921]. Council [1008, 1160, 34]. Count [419, 995, 339]. Counterexamples [908]. Counting [727, 117]. Countries [45, 211, 619]. Counts [704, 625]. Couper [1173]. Coupling [483]. Course [897, 456, 477, 710, 473, 1026, 890, 66, 1113, 1142, 476, 1034, 586]. Courts [25]. Covariance [839, 1155, 232, 1123, 597]. Cow [258]. Cows [503]. Cox [238, 478, 810, 600, 15, 631, 1046, 826, 229, 285]. Craven [243]. Credit [672]. Cressie [316]. Cricket [708, 853]. Crime [762]. Criminal [82]. Criteria [571, 796, 867]. Criterion [501]. Critical [604]. Critique [264]. Crop [22]. Cross [232, 530]. Cross-Over [530]. Cross-Section [232]. Crossing [545]. Crossover [423]. Crowder [512, 249, 835]. Cryer [1056]. Csörgo [1057]. Cuculescu [889]. Cultivation [543]. Cumulative [820, 1084]. Curnow [524]. Current [139]. Curriculum [744]. Curve [665, 881, 258, 271, 233]. CUSUM [208]. Cutting [51]. Cyclic [783]. Czitrom [1015].

D [1096, 1141, 444, 1064, 891, 1040, 487, 242, 615, 711, 781, 971, 325, 733, 516, 118, 689, 218, 1056, 727, 813, 939, 600, 490, 15, 557, 785, 1011, 1021, 610, 612, 840, 1168, 149, 1111, 457, 584, 238, 1028, 342, 713, 870, 888, 479, 243, 631, 683, 107, 455, 313, 634, 912, 450, 215, 1015, 1149, 1036, 478, 249, 944, 934, 949, 810, 1054, 116, 989, 1174, 835, 755, 13, 161, 637, 344, 318, 890, 18, 582, 833, 1101, 586, 384, 654, 601, 954, 611, 653, 857, 859, 930, 482, 779, 603, 602, 754, 863, 991, 476]. D [1167, 1027, 972, 814]. Dacre [216]. d’Adolphe [1035]. Dagnelie [513]. Dagpunar [86]. Dairy [258]. Dale [351]. Darwin [812]. Data [293, 1096, 717, 895, 822, 274, 711, 963, 209, 5, 138, 823, 724, 436, 381, 186, 490, 771, 3, 969, 1070, 964, 778, 610, 773, 423, 837, 191, 518, 133, 962, 257, 851, 455, 399, 763, 529, 650, 1152, 125, 162, 183, 1050, 354, 666, 710, 232, 692, 835, 305, 510, 1072, 1110, 375, 410, 559, 8, 503, 703, 285, 582, 814, 570, 331, 334, 310, 362, 1094, 1139, 958, 126, 719, 746, 602, 736, 754, 959, 275, 1027, 168, 339, 429, 1112, 607, 512, 939, 1145, 974, 788, 842, 836, 981, 1018, 782, 929, 313, 150]. Data [1062, 995, 606, 162, 316, 441, 322, 1174, 161, 193, 753, 833, 809, 977, 242, 659, 1103, 1171]. Databases [445]. David [238, 447, 185, 158, 968]. Davidian [754]. Davidson [971, 895]. Davis [14, 314, 858]. Dawid [932]. Daykin [603]. Dean [1168]. Debriefing [138]. Decision [797, 615, 499, 939, 259, 343, 1071, 796, 216, 111, 795, 734, 549, 583]. Decision-Making [259]. Decisions [475, 481, 1110]. Deck [375]. Deconvolution [421]. Decoupling [1177]. Definitions [206]. Degrees [1156]. Degrees-of-Freedom [1156]. DeGroot [13]. Deistler [11]. Delaney [161]. Delineate [286]. Demand [599, 824].

Demaris [440]. Dembo [514]. Demo [298]. Demo-Educational [298]. Demonstrating [1006]. Demonstrations [954]. Denise [350]. Density [457, 589, 378, 226]. Dental [749]. Dentistry [852]. Departure [642]. Dependence [965, 1177, 978]. Dependencies [810]. Dependent [282, 1068, 666, 986, 1124, 701]. Derivation [500]. Descarte [14]. Description [235]. Descriptive [735]. Design [121, 137, 1122, 1004, 142, 965, 644, 560, 190, 861, 461, 306, 20, 245, 139, 1097, 1080, 113, 44, 141, 420, 454, 838, 1168, 346, 554, 972]. Designed [317]. Designing [627, 161]. Designs [94, 979, 504, 800, 660, 290, 168, 783]. Detectable [231]. Detecting [404, 828]. Detection [365, 1158]. Determination [280, 590, 808, 592, 803, 279, 591, 77]. Determining [461, 306, 425]. Developing [45, 211, 619, 569]. Development [438, 204, 1080, 469, 1037]. Developments [261, 72]. Deviation [514]. Deviations [828]. Dhrymes [1058]. Diagnostic [538]. Diagnostics [274, 538, 1048, 205, 355]. Diagrams [401]. Dialect [961]. Diane [150]. Diaries [5]. Dichotomous [1051]. Dickey [1149]. Dictionary [1099, 389, 534]. Differences [803, 1127, 1082, 804, 866]. Different [97]. Differential [155, 517]. Difficulties [374]. Diffusions [584, 689]. Dimension [172]. Dimensionality [173]. Direct [98]. Direction [433, 147]. Directional [162, 147]. Directory [532]. Disaster [269]. Discordancy [1158]. Discovering [896]. Discovery [788, 427]. Discrete [866, 647, 742, 150, 505, 987, 867, 521, 870]. Discrete-Time [867]. Discrete-Valued [987]. Discrimination [293, 597]. Discussion [574, 878, 920, 799]. Disease [70, 1084, 882, 61, 576, 1010]. Diseases [56, 836, 731]. Dispersion [1158, 543, 409]. Dissolution [1052]. Distinguishing [1180].

Distribution [400, 393, 167, 379, 769, 774, 825, 332, 645, 125, 1084, 1120, 234, 725, 819, 334, 311, 686, 85]. Distribution-Free [1084, 725, 686, 85]. Distributions [717, 823, 104, 742, 1081, 215, 621, 802, 1157, 657, 488, 384, 154]. Diverse [1019, 197]. Diversity [66]. Dobson [159, 731]. Domain [671]. Dominance [725]. Dorling [1141]. Dosage [420]. Double [1005, 251, 376, 802]. Double-Choice [376]. Double-Sampling [802]. Drain [934]. Draper [1059]. Draw [650]. Drawing [350]. Draws [650]. Dream [14]. Driessen [87]. Drift [208]. Drug [395, 1052, 60, 1037]. Dryden [1098]. Dudewicz [253]. Dudley [184]. Duffy [150]. Duflo [1016]. Duncan [344, 583]. Dunn [780, 381]. Dunstan [349]. Dupac [1106]. Durbin [1006]. Dyke [22]. Dykstra [17]. Dynamic [521, 143, 336, 994, 426].

E. [18]. Earman [604]. East [69]. Eastern [824]. Easy [1114]. Ecological [43, 1032]. Ecology [36, 1101, 731]. Econometric [261, 1181, 995]. Econometrics [356, 380, 653, 155, 1148, 1093, 613, 89]. Economic [699, 697, 671, 507, 179, 698, 653, 61, 325]. Economically [244]. Economics [105, 217, 952, 993, 970, 380, 1147]. Economy [784, 353]. Ecuador [245]. Edgeworth [756]. Edible [469]. Edinburgh [68, 61]. Edition [684]. Editor [65, 166, 106, 289]. Editorial [121, 137, 751, 874, 50, 31]. Education [574, 887, 572, 1113, 852]. Educational [298, 445, 786]. Edward [320, 239, 82]. Edwards [755]. Eells [515]. Effect [853, 700, 543, 1086, 103, 173]. Effectiveness [854]. Effects [775, 189, 60, 286]. Efficiency [722, 628, 613, 89]. Efficient [276, 1137]. Efficiently [504]. Efron [551]. Ehrenberg [156]. Ehrlich [562]. Elderly [41]. Election [140]. Electoral [593]. Electricity [722, 364]. Electron [428]. Electronic [1056]. Element [235]. Elementary [713, 214, 449, 115]. Elements [752, 1073, 913, 936, 523, 907]. Elicitation [878, 393, 875]. Eliciting [876]. Elisabeth [673]. Elliptical [772]. Elworthy [584, 611]. Elzbieta [387]. Embleton [607]. Embrechts [1017]. EMLE [95]. Emphasising [215]. Empirical [75, 665, 910, 110, 27]. Employment [70, 244]. Encephalitis [394]. Encyclopedia [1111]. Energy [683, 258]. Engelhart [195]. Engineering [1119, 687, 384, 182, 550, 691, 326]. Engineers [1150, 188]. England [599, 69]. English [684, 649]. English/Chinese [684]. Enhancing [294]. Enrique [1090]. Entitlement [266]. Entropy [865, 1005, 98, 182, 182]. Entry [472]. Environment [1164, 43, 962, 931, 677, 718, 732, 1164]. Environmental [53, 264, 424, 734, 1101]. Envisioning [196]. Epidemic [703]. Epidemiological [851, 624, 787]. Epidemiology [396, 53, 424, 714]. Equal [1128]. Equality [6]. Equation [623, 850, 263, 947, 712, 1069]. Equations [292]. Equity [1123]. Erdos [1089]. Ergodicity [1138]. Eric [16]. Erickson [441, 182]. Erratum [90]. Error [333, 9, 142, 1002, 138, 114, 281, 292, 886, 202, 232, 295, 468, 482, 779, 1166, 944]. Errors [853, 163, 612, 1047, 415, 2, 327]. Escobar [1145]. Essays [1096]. Essential [107]. Essentials [684, 448]. Establishments [124]. Estes [323]. Estimable [644]. Estimate [721, 230]. Estimated [364]. Estimates [263, 378, 702]. Estimating [1051, 99, 704, 826, 882, 807, 312]. Estimation [333, 228, 75, 261, 1125, 88, 848, 771, 1137, 501, 457, 951, 143, 774, 506, 403, 722, 464, 507, 271, 202, 98, 4, 74, 234, 304, 2, 291, 770, 1023, 1144, 1175]. Estimator [96].

Estimators [623, 1002, 260, 667, 849]. Ethics [838]. Ethylene [704]. Europe [486, 34]. European [1170, 544, 620]. Evaluate [104]. Evaluating [292]. Evaluation [1122, 411, 564, 537, 550, 849, 940, 1061]. Event [369, 1174, 495]. Events [1017]. Everitt [516, 381, 1018, 780, 442, 632, 1099, 220]. Evidence [1001, 780, 1075, 25, 544]. Evidence-Based [780]. Evolution [412]. Evolving [25]. Eward [196]. EWMA [362]. Exact [71, 297, 97, 645, 805, 2, 1157, 692]. Examination [540, 604]. Example [718, 436, 474]. Examples [580, 494, 975]. Excellence [812]. Exchange [621]. Exercise [178]. Exercises [119]. Expansion [756]. Expectations [153]. Expected [798, 499, 1005, 420]. Experience [62]. Experiences [211, 875]. Experiment [100]. Experimental [190, 346, 461, 306, 800, 524, 454, 944, 1080, 113]. Experimentation [22]. Experiments [1122, 1045, 431, 290, 161, 702, 542, 310, 627, 861, 1168, 20, 1097, 317]. Expert [876]. Explained [926, 648, 1046, 1176]. Explanatory [926, 570]. Exploration [965, 569, 275]. Exploratory [963, 964, 962, 958, 126]. Explores [570]. Exploring [700, 959, 1112, 453]. Exponential [805, 1065]. Exposure [704, 923]. Extending [9]. Extension [1085, 296]. Extensions [988]. Extremal [1017]. Extreme [989, 334, 326].

F [1010, 895, 1164, 953, 152, 839, 937, 437, 733, 892, 509, 435, 493, 1172, 899, 17, 683, 1152, 79, 246, 349, 757, 1026, 13, 241, 389, 446, 931, 1101, 549, 677, 82, 932, 475, 1094, 553]. Faces [173]. Factor [904, 477, 171, 561, 290, 62]. Factorial [94, 1004, 644, 702, 310, 627]. Factorials [886]. Factors [883, 43, 461, 306, 70, 413, 61, 576]. Facts [156]. Failure [268]. Failures [701]. Fair [673]. Fairer [921]. Faires [218]. Falin [1100]. Falk [605]. Fallible [720]. Families [1065]. Family [365, 805, 666]. Fang [438]. Farebrother [1169]. Farhad [244]. Farnum [10]. Fatness [230]. Features [961]. Feedback [1095]. Feinberg [25]. Fellers [83]. Fentem [896]. Ferguson [489]. Fernholz [1096]. Fertility [503]. Fiction [251]. Fields [109, 989]. Figure [750]. Figure-Skating [750]. File [593]. Filter [697]. Filters [951, 671]. Final [673]. Finance [952, 989, 993, 733, 1017]. Finance-Modelling [993]. Financial [1025, 427]. Fingleton [162]. Finite [228, 1022, 554]. Fionn [327]. First [1034, 477, 1087, 710, 473, 1026, 725, 1142]. First-Order [725]. Fisher [607, 319, 643, 309, 97, 190, 670, 606, 595, 319]. Fisherian [1115]. fit [467]. Fitted [230]. Fitting [823, 881, 336, 271, 233, 1169]. Fixed [1130]. Fletcher [1101]. Flow [72]. Flury [1142]. Fly [646]. Follow [700]. Follow-Up [700]. Food [520]. Football [776, 967, 599]. Force [1125]. Forecast [426]. Forecasting [623, 325, 846, 298, 174, 688, 994, 10, 858]. Forests [294]. Foreward [360, 568]. Foreword [618, 999, 741, 845, 664, 392, 408, 256, 793, 957]. Formulae [9, 1045, 806, 544, 231]. Forward [281]. Fossati [1170]. Foster [1171]. Foundational [562]. Foundations [689, 182, 481, 1063, 1116]. Four [1087, 1086, 364]. Four-Monthly [364]. Fox [186, 940, 355, 1060]. Fractals [584]. Fractional [1004, 702, 827]. Fractionals [611]. Frailty [470]. Frames [593]. Frameworks [141]. Franfort [1019]. Franfort-Nachmias [1019]. Frank [680]. Franken [237]. Franklin [972]. Franz [32]. Fredericksen [552]. Free [1084, 725, 686, 85]. Freedman [318].

Freedom [1156]. French [1020]. Frequency [1029]. Freund [443]. Friendly [901, 1108, 680]. Front [1, 30, 49, 64, 93, 135, 136, 165, 255, 200, 224, 278, 288, 359, 391, 407, 463, 303, 329, 497, 536, 567, 588, 617, 640, 663, 696, 716, 740, 766, 792, 817, 844, 873, 915, 956, 998, 1043, 1053, 1092, 1133, 1162]. Frontiers [479]. FRS [238]. FT [621]. FT-SE [621]. Fuller [834]. Function [226]. Functional [95, 1072]. Functionals [611]. Functioning [517]. Functions [538, 104, 284, 827, 285, 312]. Fundamentals [840]. Further [645]. Future [1119].

G [752, 157, 1034, 1073, 1095, 689, 381, 1056, 780, 813, 865, 1022, 117, 785, 904, 1180, 556, 861, 693, 836, 837, 1100, 681, 315, 243, 947, 107, 321, 634, 561, 356, 1062, 1149, 162, 79, 217, 22, 192, 944, 757, 1108, 938, 1097, 1175, 948, 451, 903, 508, 344, 1110, 902, 1140, 833, 1113, 654, 1038, 525, 1055, 1114, 491, 857, 1032, 1069, 988, 977, 736, 991, 787, 476]. Galambos [973]. Gambling [870]. Game [1143, 1147]. Game-Theoretic [1143]. Gamerman [1021]. Games [87, 489, 870]. Garthwaite [730]. Gary [83, 182]. Gaspard [1102]. Gaussian [820, 424]. GCSE [21]. Gelman [711]. General [747, 225, 1144, 703, 226, 350, 719, 437, 892]. Generalisation [96]. Generalised [984]. Generalized [262, 159, 743, 666, 467]. Generated [783]. Generation [552, 86]. Genetic [1066]. Genetics [1039]. Gennaro [82]. Genomes [815]. Geochemistry [74]. Geodemographics [129]. Geoffrey [1009]. Geographic [130, 883, 962]. Geographical [569]. Geographically [960]. Geometric [991]. Geometry [1094]. George [324]. Geostatistical [274]. Geostatistics [789, 1179, 1134, 1061]. Gerber [974]. Geriatric [369]. German [62]. Germany [43]. Geschieder [324]. Getting [690].

Ghosh [1023, 1022]. Gia [280]. Gilks [781]. Gill [727]. Giltinan [754]. Giné [1177]. Given [464]. Glasgow [61]. Glen [345]. Glickman [553]. Global [35]. Glover [1172]. Glynn [1024]. Gnanadesikan [841]. Godambe [312]. Godbole [681]. Goel [558]. Goldberg [236]. Goldsmith [767]. Golf [855]. Goodness [467]. Goodness-of-fit [467]. Goovaerts [1061]. Gorman [972]. Gosset [315]. Gould [812]. Gourieroux [1025]. Governess [918]. Government [353, 1154]. Gower [863]. Gradshteyn [608]. Graduate [66]. Graham [682]. Grandeur [812]. Graphical [393, 6, 755, 701, 275, 348]. Graphics [382, 1062, 336, 977, 273, 959, 1167, 1167]. Graphing [842]. Gray [557, 1024]. Graybill [1026]. Great [350]. Green [1054]. Greenfield [893, 862, 864]. Greenwood [935]. Grenander [936]. Grenfell [731]. Griego [1114]. Griffiths [1027]. Groenen [1012]. Groenenwald [452]. Gross [583]. Ground [649]. Group [505]. Groups [307, 578, 386]. Groves [1173, 163]. Growth [697, 665, 944, 270]. Gruet [975]. Guidance [864]. Guide [1064, 974, 947, 901, 320, 893, 784, 1108, 252, 935, 1024, 384, 690, 1139, 862, 975, 679, 678]. Guidebook [787]. Gustav [252]. Gut [897]. Guttorp [782]. Guy [869].

H [353, 900, 839, 711, 614, 730, 968, 1177, 493, 518, 951, 477, 517, 555, 1104, 243, 1181, 789, 319, 1163, 1103, 454, 441, 217, 153, 949, 1179, 952, 978, 473, 113, 1093, 1026, 13, 161, 510, 389, 85, 993, 1110, 890, 712, 1059, 494, 911, 481, 680, 115, 58, 59]. Haaland [842]. Habitat [37]. Hagenaars [193]. Hajek [1106]. Hajek-with [1106]. Halberg [32]. Halbert [1181]. Hald [248]. Half [36, 542]. Half-Normal [542]. Hall [756]. Hallett [153]. Halls [55]. Hand [835, 1028, 479, 249, 863]. Handbook [632, 900, 954]. Handcock [710]. Handheld [122]. Handling [895]. Handmaiden [918]. Hannan [11]. Hanushek [475]. Härdle [732, 347]. Harff [683]. Harrison [688, 994]. Hart [1143]. Hartley [71]. Hasted [524]. Haughton [586]. Having [334]. Hawkins [553]. Hay [516]. Hayashi [1103]. Hayes [112]. Hazard [285]. Hazards [209, 545]. Hazards-Type [209]. Healey [246]. Health [38, 1164, 850, 45, 883, 32, 785, 39, 35, 700, 31, 321, 42, 40, 46, 44, 66, 47, 33, 54, 51]. Health-Related [850]. Healthy [55, 57]. Heart [70, 506, 61, 576]. Hedayat [554]. Heiberger [317]. Heidbreder [865]. Heinen [866]. Heinz [155]. Helen [21]. Henry [380]. Henze [1104]. Herbert [386]. Herdman [251]. Heresies [1044]. Hernandez [867]. Hernandez-Lerma [867]. Hersh [14]. Hershberger [1034]. Hesketh [1018]. Heterogeneity [212, 299, 424]. Heterogeneous [1156]. Heteroscedastic [333, 622, 694]. Heteroscedasticity [1047, 722, 723]. Hettmansperger [1105]. Heyde [1144]. Heyer [386]. Hidden [987]. Hierarchical [818, 427, 398]. Hierarchies [343]. High [274]. High-Interaction [274]. Higher [574, 572]. Highest [589]. Hildebrand [444]. Hilton [445]. Hinkley [238, 971]. Historical [697, 486, 899]. History [697, 370, 495, 369, 698, 248, 351, 1169, 954]. HIV [263, 398]. HIV-Infected [398]. Hjorth [609]. Hoben [154]. Hochberg [26]. Hoffmann [655]. Hoffmann-Jörgensen [655]. Holden [325]. Hölder [830]. Holland [517]. Holly [153]. Home [649, 855]. Homes [41]. Homogeneity [642]. Homogeneous [331]. Honor [1096, 324]. Honour [238]. Hoppensteadt [446]. Horváth [1057]. Hosking [1029]. Hosmer [185, 1174]. Hot [375]. Household [125, 1173]. Housing [45]. Howell [447]. Howison [733]. Howson [108]. Høyland [656]. Hoyle [712]. Hsu [868]. Huet [975]. Huisman [148, 148]. Human [1039, 836, 44, 37, 597]. Hunt [1140]. Huskova [1106]. Hussmanns [244]. Hutchinson [215, 684, 384, 448]. Hutchison [610]. Hutton [152, 1170]. Hydrology [989]. Hylleberg [480]. Hyperendemic [1084]. Hypertension [56]. Hypotheses [16, 1033]. Hypothesis [671, 266, 67, 801, 465]. Hypothesis-Based [465]. Hypothesized [333].

Ideas [1167]. Identifiability [488].

Identification [322, 701, 439]. Identified [1001]. Identifying [644, 292, 1084]. Ignatov [178]. Ihms [510]. II [996, 475]. III [982, 479, 1178]. Ikeda [611, 584]. Illustrations [969]. Image [1152]. Images [230, 44]. Implementation [104, 403]. Importance [304]. Important [1127, 627]. Imprecise [431, 499, 746]. Improved [330]. Improvement [678, 492, 893, 1015, 862]. Improvements [263]. Improving [848, 811, 475, 583, 364, 475]. Imputation [375]. Incidence [1084]. Income [332, 125, 671, 98]. Incomplete [644, 1050, 290, 339]. Incorporate [258]. Increasing [36, 1104]. Independence [230, 1177]. Independent [269, 307, 578]. Index [747, 850, 1130, 1131, 371, 471, 77, 621]. Indicators [963, 964]. Indices [626, 619]. Individual [649]. Individualized [32]. Individuals [593]. Induced [1002]. Industrial [1015, 619, 934]. Industries [520]. Industry [270, 363, 654]. Inequalities [169, 973]. Inequality [830, 1159]. Inequality-Constrained [1159]. Infant [883]. Infected [207, 398]. Infection [263]. Infectious [836, 731]. Inference [797, 880, 387, 272, 342, 554, 319, 1115, 1032, 856, 1074, 985, 730, 190, 1021, 17, 1026, 911, 600, 631]. Inference-Based [856]. Inferences [1120, 405, 422]. Infinite [702]. Inflation [544, 620]. Influence [538, 772, 401]. Information [92, 199, 302, 417, 565, 661, 764, 871, 1041, 1182, 130, 962, 143, 1155, 101, 569, 923, 613, 930, 475, 275, 888, 196]. Informative [719]. Infrared [293]. Inner [32, 39, 31, 53, 42, 40, 33, 54]. Innovative [127]. Input [507]. Institute [811, 811]. Instructor [680]. Insurance [989, 1017]. Intake [258]. Integral [647, 263]. Integrals [608]. Integrated [560, 930, 343]. Integration [752]. Intelligence [479]. Intensive [16, 609]. Interaction [274, 189]. Interactions [570, 508]. Interactive [732, 785, 273, 959, 961, 455]. Intercropping [290]. Interested [32]. Interests [66]. Interface [903, 916]. Intermediate [897, 222]. International [968, 811, 855, 13, 486]. Interpretation [540, 669, 1114, 810]. Interpreting [508, 833, 762, 869]. Interrelations [898]. Interval [796]. Interval-Based [796]. Intervals [502, 641, 629, 235, 589, 1129, 1128]. Intervention [62]. Interview [1173]. Interviewer [138]. Intraclass [4]. Introducing [8, 490]. Introduction [891, 880, 937, 780, 614, 533, 778, 159, 572, 912, 714, 789, 551, 942, 250, 949, 1179, 563, 857, 859, 481, 325, 860, 659, 899, 12, 383, 355, 815, 487, 526, 528, 1150, 861, 90, 929, 834, 757, 1097, 755, 858, 195, 525]. Introductory [66, 586, 737, 954, 484, 758]. Inverse [820, 351, 509]. Investigating [1010, 67]. Investigation [648, 345]. Investigations [220]. Ireland [575, 574, 668, 745, 744, 373, 263, 572, 270, 366]. Ischaemic [70]. Isham [836]. Issues [965, 522, 562, 712, 1037, 54, 677]. Item [517]. Iterative [1016]. Ivanov [109, 976]. Iyer [1026].

J [1064, 492, 1020, 985, 1105, 992, 839, 458, 615, 686, 711, 781, 325, 655, 489, 607, 435, 11, 758, 218, 512, 186, 1056, 738, 785, 114, 560, 658, 27, 979, 1011, 788, 841, 842, 612, 149, 867, 943, 1100, 1136, 1137, 981, 456, 735, 894, 159, 238, 1028, 555, 946, 479, 520, 683, 150, 320, 356, 86, 459, 319, 910, 452, 551, 1149, 973, 1152, 162, 898, 253, 761, 812, 838, 868, 1170, 519, 79, 249, 942, 485, 192, 783, 349, 688, 1054, 710, 950, 1151, 14, 604, 1029, 314, 585, 13, 994, 241, 637, 193, 558, 1072]. J [188, 1024, 158, 986, 189, 684, 1176, 1140, 609, 889, 908, 858, 984, 1101, 1166, 737, 178, 613, 195, 182, 324, 1055, 82, 857, 932, 583, 482, 988, 1139, 779, 1060, 341, 863, 991, 443, 1171, 905]. J. [1161]. J.-L [1152]. Jacard [189]. Jackson [320]. Jacoby [1062, 977]. Jaeger [194]. James [110, 222, 250]. Jan [155]. Jane [81]. Janssen [734]. Japan [54]. Jaroslav [1106]. Jascheck [327]. Jati [89]. Jaynes [18]. Jeffrey [608]. Jens [147]. Jensen [937]. Jewell [837]. Jiangsu [332]. Joan [354]. Joe [348, 978]. Johan [380]. John [1096, 186, 783, 111, 251, 355]. Johnson [657, 939, 1107, 784, 938, 451, 1030, 1031, 982, 449, 954]. Johnston [555]. Joint [848]. Jolivet [975]. Jolliffe [730, 553]. Jones [88, 893, 862, 790, 730, 518]. Jöreskog [904, 561]. Jörgensen [655]. Jorner [842]. Jose [1090]. Joseph [246, 590]. Judgement [414]. Judgment [697, 571]. Judgments [750]. Julius [322]. Jureckova [898]. Justice [82].

K.-W [1107]. Kadane [838]. Kadekodi [89]. Kaiser [1108]. Kalashnikov [633]. Kallenberg [1063]. Kalman [697]. Kalsbeek [482]. Kalton [344]. Kaluzny [1109]. Kanji [556]. Kaplan [96]. Karl [351]. Karlin [614]. Kasovia [785]. Kasovia-Schmitt [785]. Kaspryyk [344]. Kavalieris [1101]. Keiding [727]. Keller [217]. Kelley [450]. Kelly [733]. Kelsbeck [612]. Kempson [861, 22]. Kempton [940]. Kendall [519, 192, 9, 1085, 459]. Kenneth [236]. Kernel [969, 284, 790]. Kessler [612]. Khuri [979]. Kijima [980, 941]. Kim [481]. Kimber [512, 837]. King [1032]. Kingdom [520, 532]. Kinnear [557]. Kinney [942]. Kirkwood [250]. Klaassen [1137]. Kleijnen [452]. Klein [981, 558]. Kleinbaum [813, 634]. Klinke [732]. Klugman [1110]. Klüppelberg [1017]. Kmietowicz [179]. Knowledge [510, 944]. Knowles [857]. Koerts [380]. Koosis [1064]. Koroljuk [635]. Koronacki [492]. Kotz [982, 657, 1111, 1031]. Krause [983]. Kriging [276]. Krishnamurty [785]. Krishnan [988]. Krzanowski [839, 735]. Kuchler [1065]. Kurtosis [884, 621].

L [1096, 444, 242, 1112, 1034, 982, 325, 436, 657, 689, 218, 512, 1145, 933, 240, 560, 841, 236, 837, 1111, 981, 1057, 315, 477, 1098, 151, 17, 160, 511, 1152, 214, 445, 1175, 453, 508, 85, 1115, 987, 188, 488, 1140, 1030, 1101, 1166, 970, 311, 10, 1055, 1031, 1033, 1094, 779, 553, 1027]. Laboratories [450]. Laboratory [504]. Labour [1125]. Lactation [258]. Lad [899]. Laguna [1172]. Lai [610, 215, 384]. Lamperti [943]. Lamps [271]. Land [343]. Landmark [1094]. Lane [775]. Lange [1066]. Language [432, 1139]. Laplace [234]. Large [283, 623, 1002, 401, 650, 123, 1046, 514]. Large-Sample [283]. Larry [385]. Lasserre [867]. Latent [866, 866]. Latessa [82]. Laurel [197]. Lawler [757]. Layout [334]. Layouts [75, 658]. Lead [500, 645, 55]. Leading [1031]. Leads [53]. League [599]. Learning [924, 813, 634]. Least [881, 286]. Lebart [1112]. Lebesgue [752]. Lecture [1000]. Lectures [999, 483]. Ledger [117]. Lee [837, 195, 383, 477, 12, 559]. Leg [708]. Lehmann [1175, 1033]. Leiss [151]. Leisure [55]. Lemeshow [185, 352, 1174]. Length [369]. Leonenko [109]. LePage [453]. Lerma [867]. Leslie [8, 703]. Lessi [909]. Lessler [482]. Letchford [685]. Letter [106]. Leukaemia [67]. Level [744, 263, 235, 77, 627, 81]. Levels [51, 469, 776]. Leverage [1159]. Levine [252]. Levitas [869]. Levy [151, 352]. Lewis [607, 84, 1176, 949]. Lievesley [350]. Life [88, 470, 785, 700, 446, 812]. Life-Span [88]. Life-Tables [470]. Lifestyle [38]. Lifetime [837]. Likelihood [73, 140, 1075, 1127, 4, 378, 882, 234, 2, 856, 1144]. Likelihoods [434, 922]. Limit [466, 1057, 1068]. Limited [282, 986]. Limits [71, 626, 297, 453, 829]. Lin [1068]. Lindley [13, 798]. Lindman [454]. Lindsey [985, 758, 984]. Lindvall [483]. Line [113]. Linear [487, 309, 11, 262, 225, 848, 881, 772, 208, 401, 800, 159, 743, 1155, 628, 295, 984, 1134, 1169, 126, 1060, 286, 1014, 953, 160, 80, 79, 563, 1013, 341]. Lipsey [346]. Lipsman [1140]. Lisek [237]. Lists [807]. Litton [833]. Local [965, 963, 276, 964, 772, 958, 959]. Localities [298]. Localized [1010]. Location [1158, 820, 170]. Log [502, 126, 829, 1013]. Log-Linear [126, 1013]. Log-Normal [502]. Log-Odds [829]. Logistic [185, 759, 666, 1013, 634]. Logit [1049, 440]. Logothetis [113]. Lois [19]. Lomax [947, 1113]. London [470, 621]. Long [148, 186, 114, 986]. Longford [786]. Longitudinal [242, 53, 835, 193, 925, 518]. Look [205]. Loss [499, 420, 1110]. Lotteries [921]. Lottery [1005, 1104]. Lower [707]. Lu [1068]. Luceño [1095]. Lucia [247]. Lung [264]. Lunn [455]. Lütkepohl [900]. Lwin [27]. Lyman [187]. Lymphoblastoid [704]. Lynn [383, 350].

M

[184, 815, 608, 953, 909, 790, 1023, 512, 1022, 614, 528, 638, 163, 841, 861, 1168, 1172, 837, 980, 1065, 1106, 981, 521, 1057, 555, 243, 12, 194, 177, 941, 1165, 1016, 524, 478, 761, 1077, 519, 249, 532, 983, 22, 192, 216, 688, 996, 710, 989, 1029, 835, 13, 994, 344, 753, 935, 1173, 889, 1148, 1134, 178, 656, 954, 112, 932, 1069, 1139, 603, 181, 736, 754, 787, 317, 476, 975]. M.-A [975]. M.-G [736]. M.-L [837]. MacDonald [987]. Machin [158]. Made [921, 557]. Magnetic [230]. Magnitude [324]. Magnus [155, 341]. Magnusson [242]. Maitra [870]. Making [939, 259, 216, 549, 583]. Mallows [669]. Mammography [721]. Man [651]. Management [353, 1119, 56, 105, 217, 216, 111, 435, 734, 112, 1070]. Managers [444, 152, 949]. Mandated [151]. Mandel [658]. Manfred [11]. Mangat [907]. Manly [241, 1101]. Mann [484]. Manpower [270]. Manual [178, 680, 1109, 22]. Mapper [293]. Maps [815, 822, 425]. Marcia [383]. Marcoulides [1034]. Mardia [1098]. Marginal [73, 434, 97, 642]. Marginals [464]. Maritz [27, 686]. Mark [1004, 346, 520]. Market [129, 771, 426, 579]. Markov [689, 867, 987, 879, 721, 506, 503, 615, 781, 1021, 980, 946, 941]. Markovian [531]. Marriott [839, 389]. Marsh [293]. Martelli [521]. Martingale [996]. Martingales [689]. Marubini [736]. Maruyama [1069]. Mas [1143]. Mas-Colell [1143]. Matches [967]. Maternal [883]. Mathematic [920]. Mathematica [1024, 1140, 952, 993]. Mathematical [899, 943, 253, 258, 367, 195, 1066, 733, 149, 633, 216, 1054]. Mathematicians [945]. Mathematics [919, 917, 918, 916, 896, 446]. Mathematics-Trouble [916]. Matrices [571, 1135, 1155, 900]. Matrix [265, 8, 703, 155, 1148, 905]. Matter [29, 48, 63, 91, 1, 30, 49, 64, 120, 134, 164, 198, 277, 93, 135, 136, 165, 255, 223, 254, 287, 301, 390, 406, 416, 462, 200, 224, 278, 288, 359, 391, 407, 463, 328, 358, 303, 329, 496, 535, 566, 497, 536, 567, 587, 616, 639, 662, 588, 617, 640, 663, 695, 715, 739, 765, 696, 716, 740, 766, 791, 816, 843, 872, 792, 817, 844, 873, 914, 955, 997, 1042, 915, 956, 998, 1043, 1079, 1118, 1153, 1183, 1053, 1092, 1133, 1162]. Max [195]. Maximum [798, 1005, 4, 882, 234, 2, 719, 182, 865, 182]. Maximum-Entropy [182, 182]. Maxwell [161]. May [651, 522]. Mayo [944]. McCabe [523]. McClave [188]. McCulloch [491]. McFadden [1070]. McKean [1105]. McLachlan [988]. McNeil [455]. McPherson [345]. Mead [524, 20]. Mean [502, 1051, 628, 232, 1082]. Meaning [139]. Means [212, 299, 803, 595]. Measure [411, 564, 28, 752]. Measurement [142, 850, 292, 562, 1166, 544, 579, 776, 581, 118, 1011, 754, 560, 779]. Measurements [953, 839]. Measures [580, 648, 169, 1131, 884, 386, 1046, 594, 642, 170, 249, 889]. Measuring [1121, 419, 854, 784]. Mechanics [1102]. Medhi [738, 485, 737]. Medians [768]. Medical [901, 625, 220, 582, 602, 231, 158]. Medicine [369, 446, 852, 780, 349]. Medley [836]. Meeden [1022]. Meeker [1145]. Meel [1180]. Meeting [1007, 13]. Mehran [244]. Meier [96]. Melamed [1117]. Mémoire [1035]. Memorial [999, 1000]. Memory [720, 324]. Menard [759]. Mendenhall [456, 214, 187]. Mental [47]. Merriam [683]. Merwe [157]. Meta [370, 264, 527]. Meta-Analysis [370, 264]. Meta-Analytic [527]. Metabolizable [258]. Metcalfe [687]. Meteorology [909]. Meter [364]. Method [798, 1004, 879, 721, 881, 706, 466, 1082, 483, 340]. Methodology [121, 137, 534, 388]. Methods [1066, 940, 1034, 992, 971, 509, 6, 186, 1022, 1145, 638, 190, 899, 633, 447, 16, 146, 356, 537, 524, 522, 838, 507, 352, 748, 448, 934, 204, 692, 241, 220, 510, 559, 609, 737, 970, 601, 182, 954, 498, 310, 494, 181, 420, 991, 787, 864, 275, 402, 300, 1105, 686, 865, 27, 1107, 1103, 868, 938, 451, 314, 684, 990, 10, 656, 1060, 443, 1010, 760]. Mettrick [654]. Michael [158]. Michelson [901]. Micronuclei [704]. Microsimulation [475]. Mid [641]. Mid- [641]. Midvariances [307]. Migration [225]. Mikosch [1017]. Milk [367]. Miller [81, 902]. Mineral [683]. Mini [411, 564]. Mini-Roundabouts [411, 564]. Minimum [231]. Minitab [857].

Mirazchiiski [178]. Misconceptions [917, 927]. Misconceptions-Statistics [917]. Mishra [253]. Mislevy [552]. Missing [399, 763, 305]. Misspecification [286]. Mitchell [486]. Mixed [262, 2]. Mixing [1068]. Mixture [335, 234]. Mixtures [421]. Model [439, 1121, 1002, 261, 209, 110, 298, 883, 848, 771, 945, 1049, 414, 967, 773, 171, 296, 1155, 269, 506, 424, 258, 622, 694, 1127, 268, 72, 232, 1046, 415, 426, 1003, 609, 503, 818, 2, 1124, 427, 624, 62, 576, 168, 585]. Model-Based [1121]. Modeling [891, 712, 980, 1117, 1069]. Modelling [717, 699, 309, 850, 960, 758, 225, 343, 432, 133, 238, 782, 855, 440, 650, 543, 367, 622, 694, 621, 698, 952, 1174, 993, 646, 1083, 475, 624, 614, 947, 941, 755, 24, 313, 480, 582, 1017, 602]. Models [333, 866, 953, 1025, 274, 733, 118, 262, 263, 727, 267, 877, 335, 772, 837, 230, 584, 894, 966, 743, 160, 323, 371, 348, 23, 805, 472, 625, 666, 485, 336, 216, 810, 978, 903, 826, 1159, 510, 753, 987, 488, 1110, 986, 295, 925, 1013, 656, 409, 809, 596, 376, 597, 126, 1060, 286, 770, 754, 88, 528, 1137, 159, 1016, 237, 994, 984, 779, 1014, 836, 786]. Modern [523, 222, 1063, 55, 1040, 186, 253, 1117, 990, 1012]. Modification [706]. Modified [885]. Moeschberger [981]. Mohr [240]. Molchanov [945]. Moment [309]. Moments [226, 1029]. Monitoring [394, 1095, 361, 1101, 59]. Monte [781, 1021, 241, 623, 879]. Montgomery [487]. Monthly [364]. Moore [357]. Morgan [1114]. Morgenthaler [1096]. Morphometric [1094]. Morris [893, 862]. Mortality [470, 541, 69, 646]. Most [742, 901]. Mosteller [296]. Motion [996]. Motor [818]. Moving [885]. Moyal [1090]. Mudflats [293]. Mukhopadhyay [1023]. Multidimensional [204, 1012]. Multigroup [24]. Multilevel [883, 133, 132]. Multilocation [1004]. Multinomial [280, 330, 539, 400, 880].

Multinomial-Poisson [539]. Multiobjective [734]. Multiple [675, 851, 674, 472, 375, 189, 494, 1124, 578, 868, 26, 385, 508]. Multiple-Entry [472]. Multiple-Shrinkage [675]. Multiregression [426]. Multistate [466]. Multivariate [747, 1073, 458, 381, 760, 1048, 365, 735, 348, 1062, 337, 1054, 311, 1142, 595, 422, 1034, 839, 1180, 457, 789, 810, 1179, 950, 978, 1074]. Murphy [1146]. Murray [149]. Murtagh [1152, 327]. Musical [673]. Myors [1146].

N [1040, 940, 982, 1023, 607, 657, 727, 907, 1180, 1071, 837, 584, 238, 552, 1104, 1163, 1178, 906, 524, 606, 1103, 316, 253, 109, 239, 810, 473, 113, 1030, 654, 10, 611, 550, 1031, 1059, 786, 181, 1116]. Nachmias [1019]. Nagaraja [1163, 473]. Narayana [321]. National [1005, 1007, 445, 98]. Nations [37]. Natural [294, 904, 561, 1061, 731]. Nature [157, 1154, 498]. Nearest [672]. Neave [85]. Necessarily [142]. Necessary [171]. Needs [245]. Neighbour [672, 290]. Nenadic [450]. Neo [1115]. Neo-Fisherian [1115]. Ness [1166]. Nested [660]. Net [202, 128]. Netherlands [122, 143]. Networks [395, 937]. Neudecker [155]. Neural [421]. Nguyen [890]. Nielsen [600, 15, 631]. Nikulin [935]. Nineteenth [251]. Nineteenth-Century [251]. Noether [525]. Noise [461, 306]. Nolan [659]. Nominal [594, 642]. Non [1078, 960, 881, 700, 292, 847, 146, 300, 1081, 125, 577, 284, 233, 308, 102, 285, 525, 612, 79]. Non-Linear [881, 79]. Non-Measurement [292]. Non-Normality [1078, 308]. Non-Parametric [284, 233, 102, 525]. Non-Proportional [285]. Non-Randomized [577]. Non-Response [700, 146, 300, 125]. Non-Sampling [612]. Non-Smooth [1081]. Non-Stationarity [960]. Non-Stationary [847]. Nonlinear [953, 976, 779, 975, 1136, 951, 322, 754].

Nonnegative [1135]. Nonparametric [1105, 750, 347, 1157, 356]. Nonresponse [1173]. Nonsampling [482]. Noreen [16]. Normal [502, 104, 803, 628, 1128, 234, 311, 542, 849, 802]. Normal/Laplace [234]. Normality [1078, 337, 308]. Norris [946]. North [69, 61]. North-East [69]. Northern [373, 270, 366]. Nosanchuk [441]. Note [227, 927]. Novel [94, 275]. Number [174, 173]. Numbers [650, 398]. Numerical [167, 218]. Nutrition [321]. Nuts [469].

O [514, 909, 489, 600, 490, 15, 867, 631, 1063, 1149, 1072, 1114, 932, 531]. Obituary [1009, 928]. Objectives [66]. Observational [736]. Observations [101, 1169]. Occupancy [720]. Oceanography [509]. Ochi [177]. O’Connel [160]. Odds [1130, 829, 770, 576, 1076]. Off [113]. Off-Line [113]. Official [784, 869]. Oil [469]. Oil-Seeds [469]. Old [372, 271]. Oliver [767]. Olson [983]. Olympic [750]. Onchocerciasis [207]. Onchocerciasis-‘Infected [207]. One [502, 75, 148, 338, 334, 1005]. One-Average [148]. One-Sample [338]. One-Sided [502]. One-Way [75, 334]. OPCS [53]. Open [749]. Operational [899]. Operations [112, 1147]. Operative [87]. Optimal [335, 1144, 821, 168, 7, 153]. Optimality [867]. Optimization [598, 216, 903, 903]. Optimum [830]. Ord [459, 519, 192]. Order [1085, 473, 725, 627, 17]. Ordered [405, 310]. Ordering [472]. Ordinal [1121, 594]. O’Reilly [1036]. Organisation [510]. Organisations [532]. Organizations [921]. Origin [922]. Orr [840]. Orthogonal [99, 627]. Osborn [1140]. Ostroff [785]. Ostrom [180]. Other [504, 989, 987]. Ott [187, 444, 214]. Ottsumi [1103]. Outcomes [883, 52, 674]. Outlier [1048, 365, 1158]. Outpatient [72]. Over-Round [1131]. Overcome [206]. Overdispersed [773].

Overlapping [1128]. Overview [724, 411, 564]. Oxide [704].

P [1010, 1039, 940, 1105, 1109, 534, 484, 733, 790, 1023, 382, 727, 730, 533, 557, 785, 968, 1180, 610, 236, 693, 837, 1137, 981, 681, 782, 517, 870, 215, 1015, 83, 1058, 1061, 898, 448, 237, 562, 1054, 1151, 113, 312, 1161, 510, 558, 85, 344, 935, 1173, 684, 1176, 1012, 858, 1017, 832, 384, 513, 857, 756, 932, 690, 494, 531, 731, 1102]. Pace [1115]. Package [141]. Page [787]. Pages [448]. Paired [650]. Panchapakesan [1071]. Pandurang [928]. Panel [125, 128, 344, 145]. Panjer [1110]. Pannatier [814]. Pantula [1149]. Papastavridis [681]. Paper [280, 590, 798]. Papers [575, 668, 745, 878, 920, 18, 681, 475]. Paradigm [1075]. Paradoxes [157]. Parameter [400, 771, 167, 1144, 403, 271, 88]. Parameters [333, 880, 644, 507, 74, 405]. Parametric [284, 233, 102, 525, 985]. Parasuicide [68]. Pareto [774]. Part [839, 996]. Parthasarathy [489, 1147]. Partial [877, 101, 594, 102]. Particle [103]. Particular [56]. Patch [1120]. Pattern [936, 569]. Patterns [400, 58, 681, 369]. Paul [352, 221, 1089]. Pay [1051]. PC [557]. Pearce [22]. Pearson [643, 309, 71, 315, 351]. Peck [487]. Pedhazur [560]. Peel [325]. Peña [1177]. Penalized [378]. Tk [964]. Pennypacker [555, 555]. Pensée [1035]. Pentikainen [603]. People [921]. Percentage [104]. Percentages [650]. Performance [623, 707, 1049, 1130, 1131, 855]. Performances [966]. Periodic [331]. Periods [1084]. Permanent [294, 671]. Personalities [1031]. Perspective [35, 42, 452, 1115]. Perspectives [33, 891]. Persson [842]. Peskin [446]. Pesonen [603]. Peter [152, 12, 314, 108, 350]. Peters [636]. Petroleum [74]. Pfeiffer [221]. Pflug [903]. Pham [280]. Pham-Gia [280]. Pharmaceutical [103]. Phase [721]. Philip [14]. Philosophical [899, 562]. Physical [509, 877]. Physics [18]. Pickles [349]. Pig [230]. Pigs [366]. Pine [646]. Pisani [318]. Plackett [315]. Plan [461, 306, 7]. Plane [1014]. Planning [245, 89]. Plans [802]. Plant [940]. Plato [117, 812]. Played [708]. Players [650]. Playing [854]. Pleszczyńska [387]. Plot [294]. Plots [337, 542]. PLUS [975, 969, 1040, 983, 632]. Pocket [893, 862]. Point [1057, 227, 1175, 23]. Points [104]. Poisson [539, 282, 804]. Pole [688]. Policy [1154, 245, 475, 1170]. Politics [1141]. Polls [140]. Pollutants [543]. Pollution [51, 931]. Polynomial [827]. Pons [526]. Pooled [232, 19]. Pooling [283, 504]. Poor [69]. Population [412, 1022, 212, 299, 143, 554, 235, 139, 46, 244, 207, 350, 291]. Populations [352, 24, 731]. Pork [1122]. Positive [721]. Possibility [674]. Post [697, 144]. Post-1850 [697]. Post-Stratification [144]. Postal [123, 100]. Postcode [593]. Posterior [419, 589]. Posteriori [719]. Postgraduates [864]. Powell [356]. Power [346, 501, 705, 1146, 806, 828, 804, 231, 927]. Powerful [742]. Prabhu [1116]. Practical [678, 580, 687, 127, 440, 893, 203, 876, 787, 862, 975, 835, 112, 603]. Practice [1096, 1020, 43, 457, 413, 1030, 273, 1027, 387, 781]. Practitioners [945]. Prakasa [488]. Pre [140]. Pre-Election [140]. Precedence [724]. Precision [128]. Preclinical [721]. Predicting [400, 140, 472, 270]. Prediction [430, 400, 167, 265, 401, 1129, 284, 72]. Predictive [721, 1049]. Preference [822]. Prehistoric [404]. Preliminary [1047]. Preparations [103]. Prerequisite [769]. Present [1031]. Presented [575, 668, 745]. Pressure [58]. Pretest [1002]. Prevalence [366, 882, 8, 703]. Prevalent [576]. Prevention [62]. Price [474]. Priestley [909]. Primer [1140, 991]. Principal [538, 675, 320]. Principled [832]. Principles [537, 938, 451, 1027, 895, 191, 219, 86, 1115]. Print [690]. Prior [430, 393, 431, 499, 877, 397, 427, 746]. Priors [424, 628]. Probabilistic [376, 515]. Probabilities [75, 431, 499, 472, 746]. Probability [744, 636, 528, 414, 1150, 584, 681, 177, 1063, 337, 488, 248, 18, 889, 178, 195, 351, 611, 226, 291, 115, 605, 184, 693, 897, 713, 119, 908, 655, 943, 888, 1165, 1178, 906, 942, 386, 188, 586, 221, 930]. Probit [848]. Problem [797, 924, 212, 281, 720, 203, 405, 795, 595, 691, 1032, 679]. Problem-Solving [203]. Problems [114, 584, 125, 66, 206, 611, 54, 636, 996, 605]. Procedure [330, 283, 361, 922]. Procedures [1180, 898, 527, 26]. Proceedings [13]. Process [492, 747, 148, 921, 626, 967, 769, 23, 1015, 83, 934, 1080, 596, 522]. Processes [689, 533, 867, 980, 941, 910, 239, 703, 96, 615, 727, 738, 1065, 342, 177, 83, 757, 890, 1138]. Processing [403, 1152, 123, 436]. Product [309, 822, 466, 1080, 426]. Production [619, 367]. Productiveness [637]. Products [608]. Professional [599]. Professionals [55]. Profile [468]. Prognosis [683]. Program [432]. Programming [80, 216, 563, 1139]. Progression [721]. Prokhorov [1178]. Promoting [54]. Promotion [46]. Proper [114]. Properly [419]. Properties [94, 226]. Prophylaxis [394]. Proportion [704]. Proportional [647, 209, 285, 576]. Proportional-Integral [647]. Proportions [71, 297, 546, 589]. Proposals [583]. Prospect [364]. Prospective [72, 927]. Protocol [469]. Province [332]. Proximity [1018]. Proxy [825, 819]. Psychological [324]. Psychology [954, 447]. Public [38]. Purchase [1083]. Purves [318]. Puzey [893, 862].

Q [1020, 1145]. Qualitative [891]. Quality [492, 242, 1119, 678, 893, 361, 113, 498, 862, 113]. Quan [297]. Quantiles [578]. Quantitative [891, 114, 10]. Quantum [157]. Quasi [1144]. Quasi-Likelihood [1144]. Queen [673]. Questionnaire [139]. Questions [1114, 1014]. Quetelet [1035]. Queueing [485, 859, 1116]. Queues [430, 1100, 531]. Queuing [633].

R [581, 184, 1096, 492, 752, 242, 940, 486, 781, 110, 218, 512, 727, 865, 939, 600, 117, 15, 557, 1075, 907, 190, 678, 904, 163, 841, 842, 861, 1106, 1135, 518, 456, 652, 315, 552, 946, 1107, 17, 520, 1146, 631, 947, 160, 455, 155, 561, 438, 319, 20, 551, 893, 995, 524, 478, 454, 214, 1117, 22, 783, 250, 527, 810, 938, 116, 952, 451, 989, 1029, 1026, 453, 340, 734, 85, 188, 993, 318, 196, 189, 902, 1173, 18, 990, 1140, 449, 889, 1148, 858, 1013, 1113, 832, 549, 10, 550, 712, 809, 491, 1059, 1169]. R [869, 779, 341, 991, 787, 896, 1014, 443, 317, 605, 862, 1167, 905, 972]. R.-D [989]. Rabe [1018]. Rabe-Hesketh [1018]. Rabi [239]. Race [39]. Races [775]. Radhaiah [321]. Radicke [1082]. Radioactivity [409]. Raghavan [489, 1135]. Ragin [340]. Rainfall [212]. Raj [380]. Ralf [244]. Ramsay [1072]. Random [830, 86, 1068, 625, 109, 322, 925, 1016]. Randomization [241]. Randomized [577]. Range [148]. Rank [338]. Ranked [822]. Ranking [673, 1050, 818, 310]. Rao [909, 321, 1148, 488, 1148]. Rare [882]. Rasch [139, 118]. Rate [721]. Raters [1121]. Rates [708, 886, 804]. Rather [59]. Ratio [880, 770, 324]. Rationale [933]. Ratios [1076, 829]. Ratliff [450]. Rausand [656]. Rawlings [1149]. Ray [182]. Re [117]. Re-Counting [117]. Reaching [197]. Read [1111]. Reading [364]. Readings [555]. Real [184]. Reality [33]. Reasoning [1107, 108]. Reassessment [571]. Recapture [851]. Recidivism [209].

Recommendations [475]. Records [1163]. Recovery [173]. Reducing [51]. Rees [107]. Reference [880, 56, 58, 235]. Referenced [274]. Regimens [420]. Regional [1084, 1029]. Regions [629, 1127]. Registered [143]. Registers [593]. Regression [487, 1002, 282, 960, 848, 99, 281, 772, 675, 230, 634, 759, 260, 667, 1149, 674, 625, 666, 336, 284, 116, 233, 1174, 1046, 1159, 180, 415, 508, 753, 1003, 205, 990, 827, 1166, 1059, 1060, 1167, 975, 347, 185, 976, 79, 189, 1013, 986, 809, 355, 474, 1167]. Regression-Modelling [960]. Regressor [546]. Reid [238]. Reinsel [1073]. Reiss [989]. Related [850, 489, 1103, 677, 1060]. Relating [265]. Relation [836]. Relationship [95, 1052, 917]. Relationships [1127, 1169]. Relative [304, 231]. Release [409]. Reliability [512, 88, 1145, 837, 637, 656, 638, 23, 247, 549, 550]. Remarks [887]. Rencher [760, 1074]. Repeat [156]. Repeat-Buying [156]. Repeated [953, 839, 580, 249, 625, 754]. Repetition [68]. Reply [799]. Report [1008, 1160]. Requirements [804]. Resampling [1125, 494]. Resampling-Based [1125, 494]. Research [129, 152, 56, 43, 191, 1126, 1149, 522, 532, 124, 954, 1114, 33, 555, 242, 526, 114, 346, 530, 555, 323, 246, 445, 527, 950, 582, 602, 1147, 972, 864]. Researchers [385]. Residual [99, 281]. Resonance [230]. Resources [1061, 683]. Respondent [138]. Response [1121, 138, 52, 335, 700, 146, 300, 125, 376, 168, 979]. Restricted [17, 286]. Results [623, 926, 724, 645, 595]. Retirement [400]. Retrial [1100]. Retrospective [705, 143, 927]. Return [621]. Returns [622, 694, 1123]. Reuben [14]. Review [581, 808, 184, 1096, 1141, 444, 353, 1064, 891, 1040, 1010, 439, 1039, 815, 487, 492, 732, 895, 900, 866, 1020, 1164, 985, 1066, 483, 752, 242, 940, 1019, 1112, 608, 641, 514, 148, 157, 953, 1034, 1105, 1109, 909, 1025, 1073, 152, 992, 839, 937, 486, 534, 387, 437, 458, 615, 679, 686, 711, 781, 971, 982, 325, 484, 636, 655, 526, 733, 516, 87, 489, 790, 892, 1023, 1095, 509, 607, 110, 118, 435, 436, 11, 382, 657, 689, 758, 860, 218, 512, 381, 913, 88, 186, 345, 936, 1056, 780, 727, 813]. Review [865, 939, 1022, 1145, 600, 117, 24, 347, 490, 945, 15, 523, 614, 730, 528, 533, 638, 738, 557, 785, 933, 114, 240, 560, 658, 1075, 1076, 27, 969, 1150, 907, 974, 979, 1011, 190, 678, 1070, 904, 163, 343, 968, 1177, 580, 760, 778, 788, 811, 841, 842, 1180, 1021, 493, 556, 610, 612, 659, 222, 236, 840, 861, 1168, 1172, 149, 693, 899, 1071, 836, 837, 867, 897, 943, 980, 1065, 1100, 1106, 1111, 1136, 1137, 191, 346, 981, 633, 1135, 687, 518, 521, 25, 456, 457, 584, 635, 681, 951, 1057, 652, 735]. Review [894, 1018, 159, 238, 442, 1028, 342, 713, 315, 530, 19, 782, 219, 447, 477, 517, 552, 555, 870, 946, 1104, 888, 1098, 151, 1107, 479, 84, 243, 17, 520, 1146, 12, 90, 929, 1144, 631, 1181, 947, 16, 554, 683, 107, 160, 194, 440, 455, 313, 321, 634, 912, 150, 155, 185, 323, 450, 80, 901, 348, 759, 320, 561, 714, 789, 356, 438, 86, 154, 177, 215, 459, 319, 941, 1163, 1165, 1178, 20, 906, 910, 1063, 23, 452, 551, 632, 893, 1015, 1062, 1149, 83, 973, 995, 1016, 529, 245, 511, 524, 606, 976, 1103, 1152, 1058, 1036]. Review [522, 1061, 162, 316, 478, 898, 187, 247, 253, 183, 454, 761, 812, 838, 868, 728, 784, 26, 119, 1077, 1170, 441, 519, 79, 249, 495, 1068, 214, 246, 385, 352, 532, 445, 354, 942, 983, 1099, 1117, 485, 448, 179, 217, 22, 109, 192, 783, 834, 944, 934, 216, 250, 349, 204, 147, 237, 153, 239, 322, 527, 562, 688, 757, 949, 1108, 810, 938, 1054, 1097, 1179, 116, 996, 1175, 948, 710, 1067, 952, 451, 460, 950, 1151, 978, 989, 14, 111, 386, 473, 604, 1174, 1029, 314, 903, 480, 692, 835, 755, 113]. Review [682, 1093, 81, 252, 1026, 585, 13, 161, 312, 994, 180, 241, 453, 637, 340, 156, 193, 220, 508, 510, 734, 558, 1072, 389, 251, 85, 344, 753, 935, 1115, 987, 188, 488, 993, 1024, 1110, 158, 318, 986, 196, 248, 189, 244, 890, 902, 1173, 18, 563, 684, 1176, 990, 1140, 449, 1012, 559, 229, 446, 515, 609, 889, 908, 1030, 1148, 582, 814, 833, 858, 931, 984, 1013, 1101, 1113, 1134, 1166, 1017, 586, 737, 178, 221, 832, 970, 384, 549, 654, 380, 311, 327, 613, 10, 601, 108, 195, 656, 383, 182, 351, 685]. Review [954, 1038, 525, 611, 513, 324, 550, 653, 677, 712, 809, 1055, 1114, 491, 388, 1031, 1059, 1169, 21, 82, 89, 112, 786, 857, 859, 756, 1032, 930, 932, 1033, 1035, 1138, 475, 583, 690, 350, 482, 1069, 494, 762, 869, 988, 1142, 977, 355, 1094, 1139, 856, 1074, 779, 1060, 603, 181, 197, 911, 481, 531, 680, 341, 602, 736, 754, 863, 991, 787, 729, 864, 731, 896, 1014, 115, 443, 317, 326, 474, 476, 553, 605, 691, 862, 1037, 1167, 357, 975, 1027, 1102, 1116, 1143, 1147, 1171, 905, 972, 808, 344]. Reviews [520]. Revisions [98]. Revisited [71, 95, 297, 1081, 368, 102]. Revolutionize [273]. Reyment [904, 561]. Richard [194, 314]. Richardson [781, 197]. Riedwyl [1104]. Right [268]. Ripley [1040, 342]. Risk [883, 70, 672, 825, 819, 603, 62, 231, 576, 435]. Ritov [1137]. Road [411, 564]. Robert [24]. Robertson [17]. Robinson [353]. Robust [76, 434, 461, 306, 260, 667, 205, 210, 1105, 898]. Robyn [148]. Rogers [689]. Role [1119, 130, 267, 461, 306, 25, 363]. Rolls [298]. Roots [1058]. Rose [490, 1036]. Rosenkrantz [1150, 18]. Rosenthal [527]. Ross [528]. Rossman [788]. Rotating [128]. Rotation [145]. Round [1131]. Roundabouts [411, 564]. Row [660]. Row-Column [660]. Royall [1075]. Rubenstein [1117]. Rubin [711]. Rudas [1076]. Rudolf [245]. Rules [628, 1028]. Run [100, 72, 627]. Running [775]. Runs [681]. Ruppert [779]. Ryan [990]. Ryzhik [608].

S [1096, 1141, 439, 815, 732, 1020, 608, 1034, 1109, 992, 686, 711, 781, 982, 484, 733, 489, 657, 381, 614, 528, 933, 27, 969, 907, 974, 1071, 1111, 1136, 1135, 25, 635, 681, 1018, 442, 315, 530, 555, 554, 901, 759, 1165, 632, 893, 1149, 253, 812, 784, 352, 983, 1099, 22, 153, 322, 710, 460, 1174, 480, 692, 161, 156, 220, 508, 935, 488, 1110, 986, 446, 609, 685, 324, 1055, 491, 1031, 494, 1139, 481, 680, 474, 691, 862, 1037, 357, 975, 1143, 972, 1040, 632, 983, 1109]. S-PLUS [975, 969, 1040, 983, 632]. Safety [395, 411, 564, 23]. Sailing [1081]. Sakamoto [529]. Salem [1112]. Sales [426]. Salt [293]. Salvan [1115]. Sample [280, 330, 590, 798, 283, 294, 797, 1002, 338, 501, 800, 847, 397, 884, 546, 592, 589, 803, 796, 500, 645, 794, 799, 805, 1127, 577, 806, 1077, 1046, 279, 591, 795, 753, 141, 206, 809, 828, 807, 804, 427, 801, 231, 808]. Sample-Based [546]. Samples [1078, 175, 1049, 504, 1048, 593, 308, 350]. Sampling [280, 330, 373, 830, 769, 433, 720, 593, 124, 469, 291, 427, 420, 7, 802, 1022, 907, 554, 214, 585, 612, 352, 460]. Sampling-Based [420]. Santner [150]. Särndel [585]. Saturated [94]. Savage [562]. Saville [991]. Saw [646]. Saw-Fly [646]. Sayrs [19]. Scales [537, 579]. Scaling [204, 1012, 324]. Scattering [1102]. Scheaffer [841, 214, 188]. Schemper [1046]. Schervish [761]. Schmelkin [560]. Schmitt [785]. Schoen [24]. Schofield [901]. School [744, 298, 376]. Schott [905]. Schumacker [947]. Schumaker [511]. Schwerzel [532]. Science [142, 652, 743, 1103, 354, 748, 950, 1088, 182, 151]. Sciences [105, 1111, 66, 1031, 1035, 458, 785, 904, 236, 912, 561, 446, 1113, 970]. Scientific [345, 190, 782, 108]. Scientifically [1127]. Scientifique [1035]. Scientist [901]. Scientists [1150, 1126, 690, 490]. Scores [465]. Scoring [131, 310]. Scotland [68]. Scott [186, 114, 457]. Screening [721, 1050]. SE [621]. Sealy [315]. Search [276, 1172]. Searle [491]. Seasonality [885, 480]. Seber [79]. Second [1113]. Secondary [744]. Section [232]. Secular [828]. Seeds [469]. Seen [97]. Segments [579]. Seismic [410]. Selected [105, 681, 1181, 319, 809]. Selecting [921, 818, 570]. Selection [261, 99, 773, 706, 851, 1127, 609, 597]. Self [1064, 813, 634]. Self-Learning [813]. Self-Teaching [1064]. Semiparametric [1137, 356]. Semivariance [849]. Sen [1023, 898, 266]. Sengupta [89, 613]. Senn [530, 1037]. Sensitivity [721, 346]. Sensory [1122]. Sequences [815]. Sequential [505, 1023]. Serial [518, 1003]. Series [608, 909, 1073, 860, 212, 708, 201, 19, 847, 1058, 366, 147, 688, 314, 232, 180, 858, 596, 624, 728, 519, 192, 834, 987]. Service [52, 51]. Services [52]. Serving [1087]. Session [1008, 1160]. Set [1087]. Sets [644, 610]. Settlements [37]. Seventeenth [1031]. Several [6, 1048]. Sexual [397]. Shafer [1038]. Sham [1039]. Shamoo [191]. Shape [404, 1098, 820, 948]. Sharing [354]. Sharma [531]. Sharon [236]. Sharp [169]. Sheffield [57]. Shelly [1109]. Sherwin [637]. Shifted [922]. Shiryaev [1178, 906]. Shrinkage [675]. Sided [502, 742, 545]. Sieber [354]. Siegelman [383]. Signal [403]. Significance [933, 240, 77, 59]. Silverman [1072]. Similarities [866]. Similarity [894]. Simonelli [973]. Simonoff [992, 710]. Simple [396, 500, 826, 415, 570, 557]. Simpson [1141]. Simulated [598]. Simulation [201, 1021, 1117, 203, 903, 646, 383, 595, 452]. Simulations [1170]. Simulium [207]. Simultaneously [644]. Sinai [586]. Singh [907, 344]. Single [7, 972]. Single-Case [972]. Sinha [554]. Sir [238]. Sites [818]. Size [808, 798, 228, 797, 501, 397, 592, 589, 803, 796, 794, 799, 805, 1127, 577, 806, 1120, 279, 591, 795, 207, 173, 828, 807, 804, 801, 231]. Sizes [280, 330, 590, 800, 427]. Skating [750]. Skewed [628]. Skewness [175, 169, 884, 621]. Slater [151]. Slippage [1158]. Small [948]. Smart [569]. Smith [512, 182, 1020, 1151, 13, 182, 1059, 932]. Smoke [264]. Smooth [1081]. Smoothing [969, 770, 377, 790, 992]. Snell [238]. Snooker [547]. Soccer [649, 854]. Social [458, 526, 490, 52, 122, 1011, 652, 912, 53, 1036, 246, 532, 354, 179, 527, 124, 950, 970, 475, 1019, 522]. Society [1141, 1019, 968]. Socio [61]. Socio-Economic [61]. Software [965, 123, 270, 814]. Soil [543]. Solution [1032]. Solutions [87, 114, 670, 595]. Solving [924, 679, 203, 691]. Some [280, 590, 623, 9, 699, 1001, 374, 261, 724, 887, 3, 105, 371, 1081, 53, 592, 1044, 806, 996, 66, 206, 996]. Sorensen [1065]. Sources [414, 851, 520, 923]. Southern [824]. Space [400, 518, 368]. Space-State [518]. Space-Time [368]. Spagon [1015]. Spain [68]. Span [88]. Spatial [963, 960, 964, 342, 962, 257, 424, 316, 1120, 569, 295, 814, 958, 959, 275, 162]. Spatially [274]. SpatialStats [1109]. Spatiotemporal [541]. SPC [601, 402]. Spearman [643]. Species [1120, 807]. Specifications [877]. Spectator [194]. Spectra [428]. Spectral [378]. Spectroscopy [428]. Speed [147, 646]. Spherical [607]. Spiegelhalter [781]. Spizzichino [549]. Spline [770]. Splines [377]. Split [100]. Split-Run [100]. Spooner [949]. Sport [194]. Sports [811, 371, 471]. Sposito [80]. Spread [812]. Spreadsheet [104, 652, 172]. Sprent [690]. SPSS [974, 252, 1114, 1171, 557, 1055]. SPSS/PC [557]. Squared [935]. Squares [881, 286]. Stability [338, 771, 1138]. Stage [291]. Standard [9, 292]. Standardised [96]. Standards [58]. Standings [776]. Stanley [185, 352, 324]. Stanton [10]. Starck [1152]. STAT [963]. State [518, 558]. Stationarity [960, 1123]. Stationary [847, 237]. Statistic [1006, 669, 1156, 304]. Statistical [293, 1001, 732, 895, 1164, 985, 1066, 797, 926, 1034, 1105, 686, 11, 267, 887, 523, 974, 1011, 811, 556, 899, 1071, 1111, 346, 769, 25, 315, 243, 17, 520, 929, 160, 150, 80, 452, 632, 332, 898, 728, 1044, 942, 448, 834, 948, 775, 203, 1088, 692, 74, 389, 1115, 18, 428, 684, 827, 609, 970, 656, 852, 388, 1031, 498, 1033, 596, 376, 1074, 273, 911, 961, 553, 1102, 444, 492, 940, 387, 1095, 607, 512, 727, 1145, 730, 933, 1075, 190, 238, 342, 447, 1098, 1107, 1146, 323, 319, 1015, 1062, 83, 524, 606]. Statistical [179, 109, 934, 989, 220, 559, 1113, 1166, 737, 601, 82, 857, 977, 856, 181, 991, 443, 1037, 975]. Statistician [679, 32, 84, 893, 363, 384, 862]. Statistics [575, 574, 668, 745, 920, 1040, 919, 1019, 1119, 744, 924, 486, 534, 458, 982, 636, 758, 939, 1150, 788, 811, 842, 372, 222, 840, 105, 652, 917, 572, 479, 12, 90, 1007, 1126, 912, 155, 450, 438, 949, 710, 1088, 1026, 13, 591, 188, 158, 248, 18, 1148, 654, 954, 525, 749, 916, 932, 573, 583, 958, 718, 680, 959, 605, 824, 1157, 122, 1064, 992, 437, 484, 655, 516, 892, 382, 1056, 15, 811, 841, 456, 635, 107, 348, 356, 459, 910, 187, 253, 761, 784, 1099, 1151, 473, 682, 902, 449, 1030]. Statistics [195, 1114, 762, 869, 1142, 896, 476, 1027, 905, 1141, 1039, 1164, 913, 345, 945, 785, 236, 687, 194, 321, 316, 246, 217, 349, 938, 451, 81, 318, 1176, 931, 1101, 832, 685, 677, 21, 583, 481, 680, 691, 357]. Statistique [513]. Statman [148, 148]. Status [70, 503]. Stay [369]. Steed [1055]. Stefanski [779]. Stein [1164]. Stepwise [281]. Stern [711]. Stevens [458, 222, 324]. Stirling [1027]. Stirzaker [713]. Stochastic [614, 533, 1021, 980, 1065, 584, 272, 870, 177, 941, 478, 237, 757, 903, 547, 488, 890, 725, 172, 1138, 489, 738, 782, 485, 239]. Stock [622, 694, 621]. Stoyanov [178, 908]. Strange [926, 520]. Strategic [111, 141]. Strategies [555, 1104, 854, 197, 555]. Stratification [144]. Stratified [830, 291]. Structural [623, 850, 638, 947, 1155, 247, 1069, 712]. Structure [1045, 836, 415, 295, 173]. Structures [333, 839, 232, 341]. Stuart [459, 1009, 1091]. Stuck [1140]. Student [66, 643, 1002, 104, 315]. Students [66, 749]. Studies [396, 1051, 138, 1158, 851, 1015, 577, 231, 179, 736]. Study [1122, 138, 70, 53, 245, 139, 923, 469, 824, 62]. Studying [1167]. Style [700, 117]. Subba [909]. Subject [268]. Subjective [899, 399, 763, 693]. Subjects [398]. Subpopulation [228]. Substantial [876]. Substantive [427]. Sudderth [870]. Suddhendu [388]. Suen [219]. Sufficient [171]. Sukhatme [928]. Sullivan [490]. Sultan [563]. Sum [820, 1084]. Supply [824]. Support [734]. Surfaces [979]. Surveillance [828]. Survey [121, 137, 142, 130, 138, 436, 907, 163, 943, 133, 700, 123, 100, 214, 139, 124, 128, 585, 753, 141, 570, 126, 700, 152, 163]. Surveys [129, 1125, 122, 705, 399, 763, 1120, 145, 828, 807, 612, 840, 1077, 344, 1173, 482, 244]. Survival [837, 648, 666, 1067, 1174, 559, 285, 582, 602, 736, 813, 981, 558]. Sutherland [111]. Sweeney [654]. Sweeting [512]. Swensson [585]. Sykes [532]. Symbolic [1157]. Symmetry [227]. Symposium [574, 572]. Synaptic [421]. System [623, 294, 1136, 962, 322, 656]. Systematic [217]. Systems [130, 877, 521, 569, 11, 478, 637, 550].

T [1096, 866, 483, 1105, 1109, 909, 489, 607, 512, 730, 1076, 27, 610, 612, 1168, 837, 1135, 84, 17, 160, 450, 901, 714, 438, 215, 893, 522, 728, 119, 441, 445, 448, 216, 1067, 188, 1024, 890, 18, 684, 990, 559, 1017, 384, 786, 482, 988, 603, 787, 864, 731, 862, 1147]. Tables [643, 470, 464, 179, 305, 642, 1076, 442, 608]. Tabu [1172]. Tacq [950]. Tactics [555]. Tadeusz [387]. Taguchi [113]. Takeover [1001]. Tale [68]. Talking [516]. Tamhane [26]. Tanaka [1103]. Tanizaki [951].

Tanushev [178]. Targets [1001]. Tarozzi [157]. Tauchen [356]. Taxometric [1180]. Taylor [614]. Tcl [964]. Tcl/Tk [964]. Teach [682]. Teaching [574, 1064, 105, 572, 1126, 66, 206, 954, 749, 172, 573, 553]. Teams [708]. Technical [475]. Technique [265, 229]. Techniques [1001, 514, 15, 969, 1136, 981, 123, 950, 180, 8, 653]. Technologists [690]. Technology [111]. Teekens [245]. Templeton [1100]. Tendency [493]. Tennis [855, 1087, 1086]. Terms [361, 468, 389]. Test [434, 97, 708, 1047, 219, 671, 266, 467, 545, 4, 725, 885, 427, 1157, 552]. Testing [1078, 1004, 6, 16, 1123, 508, 308, 725, 469, 1082, 331, 494, 801, 240, 935, 786, 753, 1033]. Tests [643, 1002, 338, 750, 724, 771, 742, 1158, 674, 723, 821, 857, 376, 334, 804, 556, 552, 85]. Text [785, 813, 634, 737]. Textual [1112]. Theil [380]. Their [971, 708, 533, 836, 248, 376]. Thematic [293, 425]. Theo [87]. Theodorescu [889]. Theorems [1057]. Theoretic [796, 1143]. Theoretical [1147]. Theories [219]. Théorique [513]. Théoriques [513]. Theory [387, 11, 499, 523, 1071, 379, 769, 457, 584, 238, 552, 1181, 459, 1178, 511, 976, 868, 1068, 948, 72, 314, 1046, 795, 156, 1012, 1030, 586, 656, 611, 603, 1014, 326, 936, 943, 633, 485, 562, 604, 178, 859, 729, 115, 1116, 635, 245, 761, 1077, 1175]. Theory-Models [656]. There [855]. Thinking [444]. Third [13]. Thomas [150, 351, 154, 989, 343]. Thompson [492, 325, 110, 23, 1077, 460]. Thorne [970]. Three [1002]. Threshold [622, 694, 925, 624]. Thurstone [296]. Tibshirani [551]. Tichy [638]. Tiers [1045]. Time [909, 1073, 860, 771, 212, 201, 19, 847, 671, 728, 368, 666, 834, 366, 147, 688, 1174, 232, 987, 858, 596, 624, 275, 867, 1058, 519, 192, 314, 180]. Time-Dependent [666]. Times [268, 775]. Timmreck [714, 787]. Titterington [478]. Tobacco [264].

Todorovic [533]. Tomb [404]. Tomes [112]. Tong [311]. Tool [246, 1149]. Toolkit [1078, 308]. Tools [538, 1054, 569, 1094, 975]. Toothaker [385]. Topics [489, 388]. Totals [97]. Tournaments [855]. Town [55]. Toxoplasmic [394]. TQM [601]. Trace [235]. Traill [1176]. Training [211, 824]. Trait [866]. Transformation [539, 309, 229]. Transformations [336]. Transmission [421]. Transplant [506]. Transport [343]. Treatment [410]. Treatments [265]. Tree [894]. Trees [291]. Tremayne [523]. Trend [697]. Trends [699, 698, 828]. Trial [393, 838]. Trials [394, 530, 269, 806, 505, 736, 1070]. Triangular [825, 1081]. Tribute [32]. Trimmed [595]. Trouble [916]. True [881, 173]. Truncated [981, 809]. Tsui [1107]. Tuckwell [115]. Tufte [196]. Tukey [382, 1096]. Tumour [721]. Tuning [260, 667]. Turkkan [280]. Turkman [1164, 931, 677]. Turlach [732]. Turrisi [189]. Tutorial [680, 377]. Two [75, 1121, 148, 97, 658, 167, 742, 1158, 125, 472, 290, 545, 69, 68, 804, 291, 364, 627, 307, 578]. Two-Factor [290]. Two-Level [627]. Two-Monthly [364]. Two-Parameter [167]. Two-Process [148]. Two-Sided [742, 545]. Two-Stage [291]. Two-Way [75, 658]. Two-Year [125]. Tyne [67]. Type [209, 173, 885, 1047, 886, 973]. Types [1180].

U [936, 842, 1065, 609, 1116]. UK [1001, 744, 771, 1005, 1123, 1036]. Ultimate [207]. Umpiring [853]. Unbalanced [1049]. Unbiased [742]. Unbiasedness [750]. Uncertainties [327]. Uncertainty [410, 786, 153]. Unconfounded [370]. Underemployment [244]. Underestimating [281]. Understanding [784, 636, 240, 652, 912, 187, 441, 605, 1027]. Unemployment [143, 244]. UNESCO [36]. Unfathomable [1154]. Uniformly [742]. Unimodality [889]. Union [1170]. Unit [146, 300, 1058]. United [520, 532, 37]. Univariate [657, 977]. Universalité [1035]. Unknown [336]. Unobserved [1120]. Unreplicated [886]. Untimely [544]. Update [1111]. Updated [429]. Updating [409]. Upton [162]. Urbach [108]. Urban [38, 32, 43, 31, 42, 44, 41, 33]. Urbanization [35]. USA [744]. Usable [94]. Use [919, 921, 5, 138, 15, 122, 343, 205, 503, 703, 409, 1114, 310]. User [1109, 901, 320, 680]. User-Friendly-Workbook [680]. Uses [127, 475]. Using [293, 1122, 129, 434, 598, 209, 1005, 964, 230, 143, 632, 424, 139, 336, 426, 375, 827, 818, 542, 597, 362, 126, 286, 746, 627, 959, 576, 1171, 445, 654]. Utilities [722]. Utility [798, 933].

V [1164, 953, 937, 971, 526, 436, 1177, 836, 1106, 633, 687, 635, 238, 1098, 84, 321, 1178, 1015, 976, 22, 109, 13, 312, 931, 677, 1114]. Vaart [910]. Valencia [13, 68]. Validation [609]. Validity [933]. Valsecchi [736]. Valuation [1051]. Value [721, 334, 494, 326]. Valued [987]. Values [374, 742, 546, 989, 77]. Vanishing [171]. Vardeman [691]. Variability [650, 620, 304, 493]. Variable [99, 7]. Variables [926, 415, 570, 1124, 1068, 986]. Varian [952, 993]. Variance [73, 434, 1125, 99, 212, 299, 281, 1155, 500, 454, 1156, 593, 1128, 206, 819, 304, 2, 770, 252, 491]. Variance-Covariance [1155]. Variances [6]. Variate [86]. Variation [883, 648, 1046]. Variations [69, 331]. Variety [940]. Variogram [847]. Various [333, 923]. VARIOWIN [814]. Varying [771]. Vasudeo [928]. Vector [207]. Vega [1109]. Vegetation [807]. Vehicle [818]. Venables [1040]. Verma [244]. Version [1024]. Versus [581, 629, 796, 271]. Via [420, 578, 595, 830, 589, 1129]. View [655]. Vignettes [139]. VII [511]. Vijay [244]. Vincent [80]. Viral [67]. Visible [293]. Visiting [708]. Visits [72]. Visualization [964, 457]. Visualizing [1062]. Visweswara [321]. Vito [82]. vitro [1052]. vivo [1052]. Vogel [974]. Vogt [534]. Vol [689, 520, 182, 513, 475]. Volatility [1083]. Volume [92, 199, 302, 417, 565, 661, 764, 871, 1041, 1182, 982, 1111, 292, 162, 954, 291]. Volumes [380]. Vonesh [953]. Voss [1168]. Vote [140]. Vrieze [489].

W [581, 1096, 353, 891, 1040, 732, 1105, 839, 534, 437, 781, 892, 1056, 939, 1145, 969, 1150, 612, 943, 346, 456, 457, 735, 19, 517, 870, 151, 1107, 16, 185, 450, 356, 910, 23, 452, 1062, 511, 728, 214, 532, 179, 834, 562, 1097, 111, 1174, 180, 1072, 987, 1024, 833, 1166, 970, 10, 1169, 583, 482, 869, 977, 443, 1027]. Wackernagel [789, 1179]. Wainer [517]. Wales [599]. Walker [762]. Waller [1180]. Wallgren [842]. Wallis [1029]. Wan [189]. Wand [790]. Wang [684]. Ward [581]. Ware [954]. Warrack [217]. Water [677]. Waterman [815]. Waterton [706]. Watkins [841]. Watson [1006, 1009]. Wavelets [1108]. Way [75, 138, 658, 334, 525]. Waymire [239]. Weak [910]. Wear [67]. Weerahandi [692]. Weibull [167]. Weighted [960]. Weighting [146, 300]. Weights [540, 544]. Weinberg [236]. Weisberg [493]. Weldon [1027]. Well [35, 44]. Well-Being [35, 44]. Wellner [1137, 910]. Welsh [911]. Wermuth [810]. West [688, 994, 43, 508]. Westfall [494]. Wharton [435]. Which [644]. Whicker [383]. White [615, 1181, 1181]. Whitmore [837]. Whittaker [348]. Whitten [88]. Whittle [378]. Whittle-Likelihood [378]. Wichern [939]. Wicket [708]. Wide [749]. Wiener [611]. Wilcoxon [643]. Wild [79]. Will [55, 273]. William [315, 323, 187]. Williams [689, 783]. Willingness [1051]. Willmot [1110]. Wilmott [733]. Wilson [654, 443]. Wimbledon [1087, 1086]. Win [651, 1104, 1104]. Wind [147, 646].

Windows [1114, 1171]. Winkelmann [995]. Winnings [1005]. Wisniewski [216]. Within [919, 366, 46, 69]. Without [1055]. Witmer [841]. Wittink [116]. Wolfgang [347]. Wolfson [590]. Wood [991]. Work [139, 84]. Workbook [680]. Works [1106, 1181]. Workshop [788]. World [36]. Worthington [85]. Wretman [585]. Wright [693, 17, 912, 21]. Writing [197]. Writings [315]. Wynn [113].

X [438]. XLISP [963]. XploRe [732].

XXVIII [520].

Yajima [1103]. Yamaguchi [495]. Yang [438]. Yannoulis [179]. Year [125, 57]. Years [968, 1081, 1087, 1086]. Yor [996]. Yosef [26]. Young [494]. Yourself [682]. Yu [635, 1178].

Z [1068, 179, 684, 178]. Zeitouni [514]. Zucchini [987].

References

Anonymous:1990:FMa

[1] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 39(1):1–2, ???? 1990. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                              tronic). URL http://www.jstor.org/
                                              stable/2348188.

Sutradhar:1990:EML

[2] Brajendra C. Sutradhar. Exact maximum likelihood estimation for the mixed analysis of variance model with autocorrelated errors. Journal of the Royal Statistical Society. Series D (The Statistician), 39(1):3–9, ???? 1990. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348189.

Davies:1990:BAS

[3] J. K. W. Davies. A Bayesian analysis of some accident data. Journal of the Royal Statistical Society. Series D (The Statistician), 39(1):11–17, ???? 1990. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348190.

Paul:1990:MLE

[4] S. R. Paul and R. K. Barnwal. Maximum likelihood estimation and a C(α) test for a common intraclass correlation. Journal of the Royal Statistical Society. Series D (The Statistician), 39(1):19–24, ???? 1990. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348191.

Butcher:1990:UDD

[5] R. Butcher and J. Eldridge. The use of diaries in data collection. Journal of the Royal Statistical Society. Series D (The Statistician), 39(1):25–41, ???? 1990. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348192.

Cohen:1990:GMT

[6] Ayala Cohen. Graphical methods for testing the equality of several correlated variances. Journal of the Royal Statistical Society. Series D (The Statistician), 39(1):43–52, ???? 1990. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348193.

Yeh:1990:OSV

[7] Lam Yeh. An optimal single variable sampling plan with censoring. Journal of the Royal Statistical Society. Series D (The Statistician), 39 (1):53–66, ???? 1990. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2348194.

Saporu:1990:ILM

[8] F. W. O. Saporu. Introducing Leslie matrix techniques into the analysis of age-prevalence data. Journal of the Royal Statistical Society. Series D (The Statistician), 39(1):67–77, ???? 1990. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2348195.

Anderson:1990:ESS

[9] Oliver D. Anderson. Extending some standard error formulae of Kendall. Journal of the Royal Statistical Society. Series D (The Statistician), 39 (1):79–83, ???? 1990. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2348196.

Shearer:1990:BRB

[10] P. R. Shearer. Book review: Quantitative Forecasting Methods, by N. R. Farnum; L. W. Stanton. Journal of the Royal Statistical Society. Series D (The Statistician), 39(1):85, ???? 1990. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2348197.

Chatfield:1990:BRB

[11] Chris Chatfield. Book review: The Statistical Theory of Linear Systems, by E. J. Hannan; Manfred Deistler. Journal of the Royal Statistical Society. Series D (The Statistician), 39 (1):85–86, ???? 1990. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348198.

Higgins:1990:BRBa

[12] J. Higgins. Book review: Bayesian Statistics. An Introduction, by Peter M. Lee. Journal of the Royal Statistical Society. Series D (The Statistician), 39(1):86, ???? 1990. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348199.

Percy:1990:BRB

[13] David F. Percy. Book review: Bayesian Statistics 3: Proceedings of the Third Valencia International Meeting, by J. M. Bernardo; M. H. DeGroot; D. V. Lindley; A. F. M. Smith. Journal of the Royal Statistical Society. Series D (The Statistician), 39(1):86–87, ???? 1990. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348200.

Nason:1990:BRB

[14] Guy Nason. Book review: Descartes’ Dream, by Philip J. Davis; Reuben Hersh. Journal of the Royal Statistical Society. Series D (The Statistician), 39(1):87–88, ???? 1990. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348201.

Crowder:1990:BRB

[15] Martin Crowder. Book review: Asymptotic Techniques for Use in Statistics, by O. E. Barndorff-Nielsen; D. R. Cox. Journal of the Royal Statistical Society. Series D (The Statistician), 39(1):88, ???? 1990. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2348202.

Hollas:1990:BRB

[16] E. A. Hollas. Book review: Computer Intensive Methods for Testing Hypotheses, by Eric W. Noreen. Journal of the Royal Statistical Society. Series D (The Statistician), 39(1):89, ???? 1990. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2348203.

Hayter:1990:BRB

[17] Tony Hayter. Book review: Order Restricted Statistical Inference, by T. Robertson; F. T. Wright; R. L. Dykstra. Journal of the Royal Statistical Society. Series D (The Statistician), 39(1):90, ???? 1990. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2348204.

Ripley:1990:BRB

[18] B. D. Ripley. Book review: E. T. Jaynes: Papers on Probability, Statistics and Statistical Physics, by R. D. Rosenkrantz. Journal of the Royal Statistical Society. Series D (The Statistician), 39(1):90–91, ???? 1990. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2348205.

Grubb:1990:BRB

[19] Howard Grubb. Book review: Pooled Time Series Analysis, by Lois W.

Sayrs. Journal of the Royal Statistical Society. Series D (The Statistician), 39(1):91, ???? 1990. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348206.

Kempson:1990:BRB

[20] R. E. Kempson. Book review: The Design of Experiments, by R. Mead. Journal of the Royal Statistical Society. Series D (The Statistician), 39 (1):91–92, ???? 1990. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348207.

Stander:1990:BRBa

[21] Julian Stander. Book review: Statistics for GCSE, by Helen Wright. Journal of the Royal Statistical Society. Series D (The Statistician), 39(1):92–93, ???? 1990. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348208.

Martin:1990:BRB

[22] R. J. Martin. Book review: Manual of Crop Experimentation, by S. C. Pearce; G. M. Clarke; G. V. Dyke; R. E. Kempson. Journal of the Royal Statistical Society. Series D (The Statistician), 39(1):93, ???? 1990. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348209.

Kimber:1990:BRB

[23] Alan Kimber. Book review: Point Process Models with Applications to Safety and Reliability, by W. A. Thompson. Journal of the Royal Statistical Society. Series D (The Statistician), 39(1):94, ???? 1990. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2348210.

Cortina-Borja:1990:BRB

[24] Mario Cortina-Borja. Book review: Modelling Multigroup Populations, by Robert Schoen. Journal of the Royal Statistical Society. Series D (The Statistician), 39(1):94–95, ???? 1990. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2348211.

Gillespie:1990:BRB

[25] E. S. Gillespie. Book review: The Evolving Role of Statistical Assessments as Evidence in the Courts, by S. E. Feinberg. Journal of the Royal Statistical Society. Series D (The Statistician), 39(1):95–96, ???? 1990. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2348212.

Liu:1990:BRB

[26] Wei Liu. Book review: Multiple Comparison Procedures, by Yosef Hochberg; Ajit C. Tamhane. Journal of the Royal Statistical Society. Series D (The Statistician), 39(1):96, ???? 1990. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2348213.

Denham:1990:BRB

[27] Mike Denham. Book review: Empirical Bayes Methods, by J. S. Maritz; T. Lwin. Journal of the Royal Statistical Society. Series D (The Statistician), 39(1):97, ???? 1990. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348214.

Kinsella:1990:CBB

[28] A. Kinsella. Corrigendum: Bootstrapping a bioequivalence measure. Journal of the Royal Statistical Society. Series D (The Statistician), 39(1):98, ???? 1990. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348215.

Anonymous:1990:BMa

[29] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 39(1):99–100, ???? 1990. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348216.

Anonymous:1990:FMb

[30] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 39(2):101– 102, ???? 1990. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348528.

Holland:1990:EHI

[31] D. A. Holland. Editorial: Health of inner cities and urban areas. Journal of the Royal Statistical Society. Series D (The Statistician), 39(2):103, ???? 1990. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348529.

Cornelissen:1990:CSI

[32] Germaine Cornélissen. Challenge to the statistician interested in the individualized assessment of health in the inner cities and urban areas — a tribute to Franz Halberg. Journal of the Royal Statistical Society. Series D (The Statistician), 39(2):104– 109, ???? 1990. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348530.

Wilson:1990:HIC

[33] D. W. Wilson. Health of inner cities and urban areas: Concepts, perspectives, reality and research. Journal of the Royal Statistical Society. Series D (The Statistician), 39(2):111– 113, ???? 1990. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348531.

Hartley:1990:ABC

[34] Raymond Hartley. An address on behalf of the Council of Europe. Journal of the Royal Statistical Society. Series D (The Statistician), 39(2):115– 120, ???? 1990. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348532.

Goldstein:1990:UHW

[35] G. Goldstein. Urbanization, health and well-being: A global perspective. Journal of the Royal Statistical Society. Series D (The Statistician), 39 (2):121–133, ???? 1990. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348533.

Celecia:1990:HWI

[36] John Celecia. Half of the world and increasing: Cities, ecology and UNESCO’s action. Journal of the Royal Statistical Society. Series D (The Statistician), 39(2):135–141, ???? 1990. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348534.

Ramachandran:1990:ABU

[37] Arcot Ramachandran. An address on behalf of the United Nations Centre for Human Settlements (Habitat). Journal of the Royal Statistical Society. Series D (The Statistician), 39 (2):143–146, ???? 1990. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348535.

Ashton:1990:ULP

[38] John Ashton. Urban lifestyle and public health in the city. Journal of the Royal Statistical Society. Series D (The Statistician), 39(2):147– 156, ???? 1990. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348536.

Day:1990:RHI

[39] Michael Day. Race and health of inner cities. Journal of the Royal Statistical Society. Series D (The Statistician), 39(2):157–161, ???? 1990. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348537.

Nairn-Briggs:1990:CHI

[40] George Nairn-Briggs. The Church and health in the inner city. Journal of the Royal Statistical Society. Series D (The Statistician), 39(2):163– 167, ???? 1990. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348538.

Stros:1990:HEU

[41] O. Stros and J. Lener. Homes for the elderly in urban areas. Journal of the Royal Statistical Society. Series D (The Statistician), 39(2):169– 172, ???? 1990. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348539.

McGivern:1990:HIC

[42] William McGivern. Health of inner cities and urban areas — a Belfast perspective. Journal of the Royal Statistical Society. Series D (The Statistician), 39(2):173–184, ???? 1990. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348540.

Dietrich:1990:BEF

[43] Richard J. Dietrich. Biological and ecological factors in urban built environment: Recent research and practice in West Germany. Journal of the Royal Statistical Society. Series D (The Statistician), 39(2):185–190, ???? 1990. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348541.

Parker:1990:IHU

[44] John Parker. Images of health, urban design and human well-being. Journal of the Royal Statistical Society. Series D (The Statistician), 39 (2):191–197, ???? 1990. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348542.

Britten:1990:HHD

[45] John R. Britten. Housing and health in developing countries. Journal of the Royal Statistical Society. Series D (The Statistician), 39(2):199–203, ???? 1990. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348543.

Oliver:1990:HPW

[46] Caroline Oliver. Health promotion within an ageing population. Journal of the Royal Statistical Society. Series D (The Statistician), 39(2):205– 206, ???? 1990. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348544.

Whitehead:1990:MHA

[47] Tony Whitehead. Mental health and ageing. Journal of the Royal Statistical Society. Series D (The Statistician), 39(2):207–211, ???? 1990. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348545.

Anonymous:1990:BMb

[48] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 39(2):212, ???? 1990. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348546.

Anonymous:1990:FMc

[49] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 39(3):213– 214, ???? 1990. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2349037.

Holland:1990:E

[50] D. A. Holland. Editorial. Journal of the Royal Statistical Society. Series D (The Statistician), 39(3):215, ???? 1990. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2349038.

Heinz:1990:CCH

[51] Ingo Heinz. Cutting the costs in the Health Service by reducing air pollution levels. Journal of the Royal Statistical Society. Series D (The Statistician), 39(3):217–227, ???? 1990. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2349039.

Davies:1990:SSC

[52] Bleddyn Davies. Social services in the city: Context change, service response and service outcomes. Journal of the Royal Statistical Society. Series D (The Statistician), 39(3):229– 245, ???? 1990. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2349040.

Jones:1990:SEE

[53] D. R. Jones and P. O. Goldblatt. Social and environmental epidemiology in the inner city: Some leads from the OPCS longitudinal study. Journal of the Royal Statistical Society. Series D (The Statistician), 39 (3):247–257, ???? 1990. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2349041.

Yamada:1990:PHP

[54] Yoshimasa Yamada. Problems and health promoting issues in inner cities in Japan. Journal of the Royal Statistical Society. Series D (The Statistician), 39(3):259–262, ???? 1990. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2349042.

Morgan:1990:THT

[55] Les Morgan. Town halls, through modern leisure professionals, will lead to healthy cities. Journal of the Royal Statistical Society. Series D (The Statistician), 39(3):263–265, ???? 1990. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2349043.

Cugini:1990:CRC

[56] P. Cugini, G. Leone, L. Di Palma, P. Battisti, S. Coppola, S. Di Simone, C. Di Massimo Simonetti, and M. Comin. Chronobiological research into the clinical management of cardiovascular diseases with particular reference to arterial hypertension. Journal of the Royal Statistical Society. Series D (The Statistician), 39(3):267– 274, ???? 1990. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2349044.

Rice:1990:HSY

[57] John Noel Rice. A healthy Sheffield by the year 2000. Journal of the Royal Statistical Society. Series D (The Statistician), 39(3):275–283, ???? 1990. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2349045.

Cugini:1990:CRS

[58] P. Cugini, G. Leone, L. Di Palma, P. Battisti, D. W. Wilson, T. Kawasaki, K. Tamura, F. Halberg, and G. Cornelissen. Construction of reference standards for blood pressure 24-h patterns. Journal of the Royal Statistical Society. Series D (The Statistician), 39(3):285–299, ???? 1990. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2349046.

Tamura:1990:CSA

[59] Kohji Tamura, Hiroyuki Ishii, Shigeo Mukaiyama, and Franz Halberg. Clinical significance of ABPM monitoring for 48 h rather than 24 h. Journal of the Royal Statistical Society. Series D (The Statistician), 39 (3):301–306, ???? 1990. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2349047.

Tamura:1990:CAA

[60] Kohji Tamura, Hiroyuki Ishii, Shigeo Mukaiyama, Hiroshi Ijiri, Ikuo Tsuruta, and Takao Komaya. Chronobiologic assessment of antihypertensive drug effects. Journal of the Royal Statistical Society. Series D (The Statistician), 39(3):307–317, ???? 1990. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2349048.

Woodward:1990:CHD

[61] M. Woodward, M. C. Shewry, W. C. S. Smith, and Hugh Tunstall-Pedoe. Coronary heart disease and socioeconomic factors in Edinburgh and North Glasgow. Journal of the Royal Statistical Society. Series D (The Statistician), 39(3):319–329, ???? 1990. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2349049.

Wendt:1990:RFM

[62] Guenter Wendt. The risk factor model as a basis for community intervention — experience from the German Cardiovascular Prevention Study. Journal of the Royal Statistical Society. Series D (The Statistician), 39 (3):331–335, ???? 1990. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2349050.

Anonymous:1990:BMc

[63] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 39(3):??, ???? 1990. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2349051.

Anonymous:1990:FMd

[64] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 39(4):337– 340, ???? 1990. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2349076.

Anonymous:1990:E

[65] Anonymous. From the Editor. Journal of the Royal Statistical Society. Series D (The Statistician), 39(4):339, ???? 1990. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2349077.

Sahai:1990:SPT

[66] Hardeo Sahai. Some problems of teaching an introductory biostatistics course to graduate students in health sciences: Coping with the diversity of student aptitudes, interests and objectives. Journal of the Royal Statistical Society. Series D (The Statistician), 39(4):341–347, ???? 1990. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2349078.

Raybould:1990:CLT

[67] S. Raybould. Childhood leukaemia in the Tyne and Wear: Investigating a viral hypothesis. Journal of the Royal Statistical Society. Series D (The Statistician), 39(4):349–363, ???? 1990. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2349079.

Smeeton:1990:TTC

[68] N. C. Smeeton, D. G. Wilkinson, M. Gomez Beneyto, and V. Llopis. A tale of two cities: Repetition of parasuicide in Edinburgh (Scotland) and Valencia (Spain). Journal of the Royal Statistical Society. Series D (The Statistician), 39(4):365–371, ???? 1990. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2349080.

Phillimore:1990:MVW

[69] Peter Phillimore. Mortality variations within two poor areas in North-East England. Journal of the Royal Statistical Society. Series D (The Statistician), 39(4):373–383, ???? 1990. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2349081.

Gilbert:1990:ESR

[70] J. F. Gilbert, J. E. J. Gallacher, P. M. Sweetnam, J. W. G. Yarnell, and P. C. Elwood. Employment status and risk factors for ischaemic heart disease: The Caerphilly study. Journal of the Royal Statistical Society. Series D (The Statistician), 39 (4):385–390, ???? 1990. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2349082.

Brenner:1990:ECL

[71] D. J. Brenner and H. Quan. Exact confidence limits for binomial proportions — Pearson and Hartley revisited. Journal of the Royal Statistical Society. Series D (The Statistician), 39(4):391–397, ???? 1990. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2349083. See

Moye:1990:PMO

[72] Lemuel A. Moyé, Asha S. Kapadia, and Irina M. Cech. A prediction model for outpatient visits and cash flow: Developments in prospective run theory. Journal of the Royal Statistical Society. Series D (The Statistician), 39 (4):399–414, ???? 1990. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2349084.

Bertolino:1990:MLA

[73] Francesco Bertolino, Ludovico Piccinato, and Walter Racugno. A marginal likelihood approach to analysis of variance. Journal of the Royal Statistical Society. Series D (The Statistician), 39(4):415–424, ???? 1990. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2349085.

Perriman:1990:SEP

[74] William S. Perriman. Statistical estimation of petroleum geochemistry parameters. Journal of the Royal Statistical Society. Series D (The Statistician), 39(4):425–436, ???? 1990. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2349086.

Barry:1990:EBE

[75] Daniel Barry. Empirical Bayes estimation of binomial probabilities in one-way and two-way layouts. Journal of the Royal Statistical Society. Series D (The Statistician), 39(4):437– 453, ???? 1990. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2349087.

Abdullah:1990:RCC

[76] Mokhtar Bin Abdullah. On a robust correlation coefficient. Journal of the Royal Statistical Society. Series D (The Statistician), 39(4):455– 460, ???? 1990. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2349088.

Sutton:1990:VID

[77] J. B. Sutton. Values of the index of determination at the 5% significance level. Journal of the Royal Statistical Society. Series D (The Statistician), 39(4):461–463, ???? 1990. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2349089.

Chanter:1990:C

[78] D. O. Chanter, K. F. Yee, and Nathan Mantel. Correspondence. Journal of the Royal Statistical Society. Series D (The Statistician), 39(4):465– 466, ???? 1990. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2349090.

Loynes:1990:BRB

[79] R. M. Loynes. Book review: Non-Linear Regression, by G. A. F. Seber; C. J. Wild. Journal of the Royal Statistical Society. Series D (The Statistician), 39(4):467, ???? 1990. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2349091.

Jajuga:1990:BRB

[80] Krzysztof Jajuga. Book review: Linear Programming with Statistical Applications, by Vincent A. Sposito. Journal of the Royal Statistical Society. Series D (The Statistician), 39(4):467– 468, ???? 1990. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2349092.

Peace:1990:BRB

[81] Ralph Peace. Book review: Statistics for Advanced Level, by Jane Miller. Journal of the Royal Statistical Society. Series D (The Statistician), 39 (4):468–469, ???? 1990. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2349093.

Stander:1990:BRBb

[82] Julian Stander. Book review: Statistical Applications in Criminal Justice, by Gennaro F. Vito; Edward J. Latessa. Journal of the Royal Statistical Society. Series D (The Statistician), 39(4):469, ???? 1990. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2349094.

Kinsella:1990:BRB

[83] A. Kinsella. Book review: Statistical Process Control for Continuous Processes, by Gary P. Fellers. Journal of the Royal Statistical Society. Series D (The Statistician), 39(4):470, ???? 1990. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2349095.

Haugh:1990:BRB

[84] Larry D. Haugh. Book review: The Statistician at Work, by T. Lewis; V.

Barnett. Journal of the Royal Statistical Society. Series D (The Statistician), 39(4):470–471, ???? 1990. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2349096.

Prys-Williams:1990:BRB

[85] Allan G. Prys-Williams. Book review: Distribution-Free Tests, by H. R. Neave; P. L. Worthington. Journal of the Royal Statistical Society. Series D (The Statistician), 39(4):471– 472, ???? 1990. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2349097.

Kemp:1990:BRB

[86] C. David Kemp. Book review: Principles of Random Variate Generation, by J. Dagpunar. Journal of the Royal Statistical Society. Series D (The Statistician), 39(4):472–473, ???? 1990. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2349098.

Cannings:1990:BRB

[87] C. Cannings. Book review: Co-Operative Games, Solutions and Applications, by Theo Driessen. Journal of the Royal Statistical Society. Series D (The Statistician), 39(4):473, ???? 1990. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2349099.

Coles:1990:BRB

[88] Stuart Coles. Book review: Parameter Estimation in Reliability and Life-Span Models, by A. Clifford Cohen; Betty Jones Whitten. Journal of the Royal Statistical Society. Series D (The Statistician), 39(4):473– 474, ???? 1990. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2349100.

Stapenhurst:1990:BRB

[89] Tim Stapenhurst. Book review: Econometrics of Planning and Efficiency, by Jati K. Sengupta; K. Kadekodi. Journal of the Royal Statistical Society. Series D (The Statistician), 39(4):474, ???? 1990. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2349101.

Higgins:1990:BRBb

[90] J. Higgins. Book review: Erratum: Bayesian Statistics. An Introduction. Journal of the Royal Statistical Society. Series D (The Statistician), 39(4):475, ???? 1990. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2349102.

Anonymous:1990:BMd

[91] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 39(4):477– 479, ???? 1990. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2349103.

Anonymous:1990:VI

[92] Anonymous. Volume information. Journal of the Royal Statistical Society. Series D (The Statistician), 39 (4):481–485, ???? 1990. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2349075.

Anonymous:1991:FMa

[93] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 40(1):1, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348217.

Baker:1991:NUP

[94] A. G. Baker. Novel and usable properties of saturated factorial designs. Journal of the Royal Statistical Society. Series D (The Statistician), 40(1):3–9, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348218.

Conniffe:1991:FRR

[95] Denis Conniffe. Functional relationship revisited: The EMLE approach. Journal of the Royal Statistical Society. Series D (The Statistician), 40 (1):11–15, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348219.

Schemper:1991:GKM

[96] Michael Schemper. Generalisation of the Kaplan–Meier estimator for standardised processes. Journal of the Royal Statistical Society. Series D (The Statistician), 40(1):17–25, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348220.

Cormack:1991:FET

[97] R. S. Cormack and Nathan Mantel. Fisher’s exact test: The marginal totals as seen from two different angles. Journal of the Royal Statistical Society. Series D (The Statistician), 40(1):27–34, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348221.

Patterson:1991:DEE

[98] K. D. Patterson and S. M. Heravi. Direct estimation of entropy and revisions to the national income accounts. Journal of the Royal Statistical Society. Series D (The Statistician), 40(1):35–50, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348222.

Copas:1991:ERV

[99] J. B. Copas and Tianyong Long. Estimating the residual variance in orthogonal regression with variable selection. Journal of the Royal Statistical Society. Series D (The Statistician), 40(1):51–59, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348223.

Lynn:1991:SRE

[100] Peter Lynn. A split-run experiment in a postal survey. Journal of the Royal Statistical Society. Series D (The Statistician), 40(1):61–66, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348224.

Lien:1991:ICP

[101] Donald Lien and David Rearden. On the information content of partial observations: A Bayesian approach. Journal of the Royal Statistical Society. Series D (The Statistician), 40(1):67–72, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348225.

Schemper:1991:NPP

[102] Michael Schemper. Non-parametric partial association revisited. Journal of the Royal Statistical Society. Series D (The Statistician), 40(1):73–76, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348226.

Saunders:1991:EPA

[103] R. Saunders. The effect of particle agglomeration in pharmaceutical preparations. Journal of the Royal Statistical Society. Series D (The Statistician), 40(1):77–86, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348227.

Cox:1991:IFE

[104] M. A. A. Cox. The implementation of functions to evaluate percentage points of the normal and Student’s t distributions on a spreadsheet. Journal of the Royal Statistical Society. Series D (The Statistician), 40(1):87–94, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348228.

Gandhi:1991:TSB

[105] B. V. R. Gandhi, Hardeo Sahai, and Nancy Acevedo. Teaching of statistics in business, economics and management sciences: Some comments and a selected bibliography. Journal of the Royal Statistical Society. Series D (The Statistician), 40(1):95–106, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2348229.

Michael:1991:LE

[106] Yee Lok Shun Michael. Letter to the Editor. Journal of the Royal Statistical Society. Series D (The Statistician), 40(1):107, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2348230.

Howel:1991:BRBa

[107] Denise Howel. Book review: Essential Statistics, by D. G. Rees. Journal of the Royal Statistical Society. Series D (The Statistician), 40(1):109, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2348231.

Sheldon:1991:BRBa

[108] Neil Sheldon. Book review: Scientific Reasoning: The Bayesian Approach, by Colin Howson; Peter Urbach. Journal of the Royal Statistical Society. Series D (The Statistician), 40(1):109–110, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2348232.

Martin:1991:BRBa

[109] R. J. Martin. Book review: Statistical Analysis of Random Fields, by A. V. Ivanov; N. N. Leonenko. Journal of the Royal Statistical Society. Series D (The Statistician), 40 (1):110–111, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348233.

Chalmers:1991:BRBa

[110] Colin P. Chalmers. Book review: Empirical Model Building, by James R. Thompson. Journal of the Royal Statistical Society. Series D (The Statistician), 40(1):111–112, ???? 1991. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348234.

Nason:1991:BRB

[111] Guy Nason. Book review: Towards a Strategic Management and Decision Technology, by John W. Sutherland. Journal of the Royal Statistical Society. Series D (The Statistician), 40 (1):112–113, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348235.

Stapenhurst:1991:BRB

[112] Tim Stapenhurst. Book review: Practical Operations Management, by M. Hayes; A. Tomes. Journal of the Royal Statistical Society. Series D (The Statistician), 40(1):113, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348236.

OConnor:1991:BRB

[113] P. D. T. O’Connor. Book review: Quality Through Design: Experimental Design, Off-Line Quality Control and Taguchi’s Contributions, by N. Logothetis; H. P. Wynn. Journal of the Royal Statistical Society. Series D (The Statistician), 40(1):113– 114, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348237.

Day:1991:BRB

[114] Simon Day. Book review: Common Problems Proper Solutions: Avoiding Error in Quantitative Research, by J. Scott Long. Journal of the Royal Statistical Society. Series D (The Statistician), 40(1):114–115, ???? 1991. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348238.

Wilson:1991:BRB

[115] Jacky Wilson. Book review: Elementary Applications of Probability Theory, by H. C. Tuckwell. Journal of the Royal Statistical Society. Series D (The Statistician), 40(1):115, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348239.

Milo:1991:BRB

[116] Wladyslaw Milo. Book review: The Application of Regression Analysis, by D. R. Wittink. Journal of the Royal Statistical Society. Series D (The Statistician), 40(1):116–117, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348240.

Corlett:1991:BRB

[117] Tom Corlett. Book review: Re-Counting Plato: A Computer Analysis of Plato’s Style, by G. R. Ledger. Journal of the Royal Statistical Society. Series D (The Statistician), 40 (1):117–119, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348241.

Chalmers:1991:BRBb

[118] Colin P. Chalmers. Book review: Rasch Models for Measurement, by D. Andrich. Journal of the Royal Statistical Society. Series D (The Statistician), 40(1):119–120, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348242.

Liu:1991:BRB

[119] Wei Liu. Book review: Exercises in Probability, by T. Cacoullos. Journal of the Royal Statistical Society. Series D (The Statistician), 40(1):120– 121, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348243.

Anonymous:1991:BMa

[120] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 40(1):??, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348244.

Anonymous:1991:ESDa

[121] Anonymous. Editorial: Survey design, methodology and analysis. Journal of the Royal Statistical Society. Series D (The Statistician), 40(2):123, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348484.

DeHeer:1991:UHC

[122] W. F. De Heer. The use of handheld computers in social surveys of the Netherlands Central Bureau of Statistics. Journal of the Royal Statistical Society. Series D (The Statistician), 40(2):125–138, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348485.

Lamb:1991:TSP

[123] Joanne Lamb and Lyn Middleton. Techniques and software for the processing of a large postal survey. Journal of the Royal Statistical Society. Series D (The Statistician), 40(2):139– 143, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348486.

Millward:1991:SES

[124] Neil Millward. Sampling establishments for social survey research. Journal of the Royal Statistical Society. Series D (The Statistician), 40(2):145– 152, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348487.

Laaksonen:1991:ANR

[125] S. S. Laaksonen. Adjustment for non-response in two-year panel data: Applications to problems of household income distribution. Journal of the Royal Statistical Society. Series D (The Statistician), 40(2):153–168, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348488.

Upton:1991:EAS

[126] Graham J. G. Upton. The exploratory analysis of survey data using log-linear models. Journal of the Royal Statistical Society. Series D (The Statistician), 40(2):169–182, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348489.

Higgs:1991:PIU

[127] N. T. Higgs. Practical and innovative uses of correspondence analysis. Journal of the Royal Statistical Society. Series D (The Statistician), 40 (2):183–194, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348490.

Nieuwenbroek:1991:PNC

[128] N. J. Nieuwenbroek. Precision of net change in a rotating panel survey. Journal of the Royal Statistical Society. Series D (The Statistician), 40(2):195–201, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348491.

Baker:1991:UGM

[129] Ken Baker. Using geodemographics in market research surveys. Journal of the Royal Statistical Society. Series D (The Statistician), 40(2):203– 207, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348492.

Bond:1991:RGI

[130] Derek Bond and Paula Devine. The role of geographic information systems in survey analysis. Journal of the Royal Statistical Society. Series D (The Statistician), 40(2):209–216, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348493.

Knott:1991:SAA

[131] Martin Knott, Maria Teresa Albanese, and Jane Galbraith. Scoring attitudes to abortion. Journal of the Royal Statistical Society. Series D (The Statistician), 40(2):217–223, ???? 1991. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2348494.

Wiggins:1991:MAA

[132] R. D. Wiggins, K. Ashworth, C. A. O’Muircheartaigh, and J. I. Galbraith. Multilevel analysis of attitudes to abortion. Journal of the Royal Statistical Society. Series D (The Statistician), 40(2):225–234, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2348495.

Goldstein:1991:MMS

[133] Harvey Goldstein. Multilevel modelling of survey data. Journal of the Royal Statistical Society. Series D (The Statistician), 40(2):235–244, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348496.

Anonymous:1991:BMb

[134] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 40(2):245– 248, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348497.

Anonymous:1991:FMb

[135] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 40(2):??, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348483.

Anonymous:1991:FMc

[136] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 40(3):249, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348276.

Anonymous:1991:ESDb

[137] Anonymous. Editorial: Survey design, methodology and analysis. Journal of the Royal Statistical Society. Series D (The Statistician), 40(3):251, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348277.

Campanelli:1991:URI

[138] Pamela C. Campanelli, Elizabeth A. Martin, and Jennifer M. Rothgeb. The use of respondent and interviewer debriefing studies as a way to study response error in survey data. Journal of the Royal Statistical Society. Series D (The Statistician), 40(3):253– 264, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348278.

Martin:1991:ARA

[139] Elizabeth A. Martin, Pamela C. Campanelli, and Robert E. Fay. An application of Rasch analysis to questionnaire design: Using vignettes to study the meaning of ‘work’ in the current population survey. Journal of the Royal Statistical Society. Series D (The Statistician), 40(3):265–276, ???? 1991. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348279.

Bolstein:1991:PLV

[140] Richard Bolstein. Predicting the likelihood to vote in pre-election polls. Journal of the Royal Statistical Society. Series D (The Statistician), 40 (3):277–283, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348280.

Rennolls:1991:SFS

[141] Keith Rennolls. Strategic frameworks for a sample survey design package.

Journal of the Royal Statistical Society. Series D (The Statistician), 40 (3):285–290, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348281.

Bell:1991:BNB

[142] John F. Bell. Big is not necessarily beautiful in survey design: Measurement error and the APU science survey. Journal of the Royal Statistical Society. Series D (The Statistician), 40(3):291–300, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348282.

Heerschop:1991:RUN

[143] M. J. Heerschop and C. A. J. Lieftinck-Koeijers. Registered unemployment in the Netherlands: Estimation of a dynamic population using retrospective information. Journal of the Royal Statistical Society. Series D (The Statistician), 40(3):301–314, ???? 1991. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348283.

Smith:1991:PS

[144] T. M. F. Smith. Post-stratification. Journal of the Royal Statistical Society. Series D (The Statistician), 40 (3):315–323, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348284.

Sharot:1991:ARP

[145] Trevor Sharot. Attrition and rotation in panel surveys. Journal of the Royal Statistical Society. Series D (The Statistician), 40(3):325–331, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348285.

Holt:1991:MWU

[146] D. Holt and D. Elliot. Methods of weighting for unit non-response. Journal of the Royal Statistical Society. Series D (The Statistician), 40 (3):333–342, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2348286. See cor-

rigendum [300].

Mellor:1991:BRB

[147] Dale Mellor. Book review: The Analysis of Directional Time Series: Applications to Wind Speed and Direction, by Jens Breckling. Journal of the Royal Statistical Society. Series D (The Statistician), 40(3):343, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2348287.

Bissell:1991:BRB

[148] Derek Bissell. Book review: Statman One-Average and Range Charts, by Alan Long; Robyn Huisman; Statman Two-Process Capability, by Alan Long; Robyn Huisman. Journal of the Royal Statistical Society. Series D (The Statistician), 40(3):344, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2348288.

France:1991:BRB

[149] J. France. Book review: Mathematical Biology, by J. D. Murray.

Journal of the Royal Statistical Society. Series D (The Statistician), 40 (3):344–345, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348289.

Iyer:1991:BRBa

[150] R. Iyer. Book review: The Statistical Analysis of Discrete Data, by Thomas J. Santner; Diane E. Duffy. Journal of the Royal Statistical Society. Series D (The Statistician), 40(3):345, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348290.

Harris:1991:BRB

[151] Ray Harris. Book review: Mandated Science, by L. Slater; E. Levy; W. Leiss. Journal of the Royal Statistical Society. Series D (The Statistician), 40(3):345–346, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348291.

Bound:1991:BRB

[152] J. A. Bound. Book review: Survey Research for Managers, by Peter F. Hutton. Journal of the Royal Statistical Society. Series D (The Statistician), 40(3):346, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348292.

Metcalfe:1991:BRB

[153] A. V. Metcalfe. Book review: Optimal Control, Expectations and Uncertainty, by S. Holly; A. H. Hallett. Journal of the Royal Statistical Society. Series D (The Statistician), 40(3):347, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2348293.

Kemp:1991:BRBa

[154] A. W. Kemp. Book review: Distributions of Correlation Coefficients, by Hoben Thomas. Journal of the Royal Statistical Society. Series D (The Statistician), 40(3):347–348, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348294.

Iyer:1991:BRBb

[155] R. Iyer. Book review: Matrix Differential Calculus with Applications in Statistics and Econometrics, by Jan R. Magnus; Heinz Neudecker. Journal of the Royal Statistical Society. Series D (The Statistician), 40(3):349, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348295.

Pigden:1991:BRB

[156] Norman Pigden. Book review: Repeat-Buying: Facts, Theory and Applications, by A. C. S. Ehrenberg. Journal of the Royal Statistical Society. Series D (The Statistician), 40(3):349– 350, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348296.

Blodwell:1991:BRB

[157] J. F. Blodwell. Book review: The Nature of Quantum Paradoxes, by G.

Tarozzi; A. van der Merwe. Journal of the Royal Statistical Society. Series D (The Statistician), 40(3):350– 351, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348297.

Rees:1991:BRB

[158] D. G. Rees. Book review: Medical Statistics — A Commonsense Approach, by Michael J. Campbell; David Machin. Journal of the Royal Statistical Society. Series D (The Statistician), 40(3):351, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348298.

Gough:1991:BRB

[159] Kevin Gough. Book review: An Introduction to Generalized Linear Models, by J. Annette Dobson. Journal of the Royal Statistical Society. Series D (The Statistician), 40(3):352, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348299.

Howel:1991:BRBb

[160] Denise Howel. Book review: Linear Statistical Models: An Applied Approach, by B. L. Bowerman; R. T. O’Connel. Journal of the Royal Statistical Society. Series D (The Statistician), 40(3):352–353, ???? 1991. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348300.

Percy:1991:BRB

[161] David F. Percy. Book review: Designing Experiments and Analyzing Data, by S. E. Maxwell; H. D. Delaney. Journal of the Royal Statistical Society. Series D (The Statistician), 40 (3):353–354, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2348301.

Lawson:1991:BRB

[162] Andrew Lawson. Book review: Spatial Data Analysis: Volume 2 Categorical and Directional Data, by G. J. G. Upton; B. Fingleton. Journal of the Royal Statistical Society. Series D (The Statistician), 40(3):354– 355, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348302.

Edwards:1991:BRB

[163] Phil Edwards. Book review: Survey Errors and Survey Costs, by R. M. Groves. Journal of the Royal Statistical Society. Series D (The Statistician), 40(3):355–356, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2348303.

Anonymous:1991:BMc

[164] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 40(3):357– 360, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348304.

Anonymous:1991:FMd

[165] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 40(4):361, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348723.

Haws:1991:E

[166] Tony Haws. From the Editor. Journal of the Royal Statistical Society. Series D (The Statistician), 40(4):363, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348724.

Dellaportas:1991:NPT

[167] Petros Dellaportas and David E. Wright. Numerical prediction for the two-parameter Weibull distribution. Journal of the Royal Statistical Society. Series D (The Statistician), 40 (4):365–372, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348725.

Yanagisawa:1991:ODB

[168] Yukio Yanagisawa. Optimal designs for a beta binary response model and the analysis of binary communication data. Journal of the Royal Statistical Society. Series D (The Statistician), 40(4):373–385, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348726.

Groeneveld:1991:SIS

[169] Richard A. Groeneveld. Sharp inequalities for skewness measures. Journal of the Royal Statistical Society. Series D (The Statistician), 40(4):387– 392, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348727.

Wilcox:1991:CBM

[170] Rand R. Wilcox. Comparing biweight measures of location. Journal of the Royal Statistical Society. Series D (The Statistician), 40(4):393– 399, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348728.

Haagen:1991:NSC

[171] Klaus Haagen. Necessary and sufficient conditions for vanishing the arbitrariness in a common-factor model. Journal of the Royal Statistical Society. Series D (The Statistician), 40 (4):401–408, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2348729.

Stent:1991:SSN

[172] A. F. Stent and L. G. McAlevey. The stochastic spreadsheet: A new teaching dimension. Journal of the Royal Statistical Society. Series D (The Statistician), 40(4):409–413, ???? 1991. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2348730.

Saxena:1991:ECS

[173] P. C. Saxena and K. Navaneetham. The effect of cluster size, dimensionality, and number of clusters on recovery of true cluster structure through Chernoff-type faces. Journal of the Royal Statistical Society. Series D (The Statistician), 40(4):415–425, ????

1991.  CODEN ????  ISSN

0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348731.

Gamerman:1991:FNA

[174] Dani Gamerman and Hélio S. Migon. Forecasting the number of AIDS cases in Brazil. Journal of the Royal Statistical Society. Series D (The Statistician), 40(4):427–442, ???? 1991. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348732.

Butcher:1991:SSC

[175] R. Butcher. Skewness in samples from the census. Journal of the Royal Statistical Society. Series D (The Statistician), 40(4):443–448, ???? 1991. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348733.

Greene:1991:C

[176] Brook A. Greene and Samuel Cameron. Correspondence. Journal of the Royal Statistical Society. Series D (The Statistician), 40(4):449–450, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348734.

Kemp:1991:BRBb

[177] A. W. Kemp. Book review: Applied Probability and Stochastic Processes, by M. K. Ochi. Journal of the Royal Statistical Society. Series D (The Statistician), 40(4):451–452, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348735.

Selby:1991:BRB

[178] Michael J. P. Selby. Book review: Exercise Manual in Probability Theory, by J. Stoyanov; I. Mirazchiiski; Z. Ignatov; M. Tanushev. Journal of the Royal Statistical Society. Series D (The Statistician), 40(4):452, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2348736.

Marriott:1991:BRB

[179] Alan Marriott. Book review: Statistical Tables for Economic, Business and Social Studies, by Z. W. Kmietowicz; Y. Yannoulis. Journal of the Royal Statistical Society. Series D (The Statistician), 40(4):452–453, ???? 1991. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2348737.

Pewsey:1991:BRB

[180] Arthur Pewsey. Book review: Time Series Analysis (Regression Techniques), by C. W. Ostrom. Journal of the Royal Statistical Society. Series D (The Statistician), 40(4):453–455, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348738.

Vaughan:1991:BRBa

[181] T. B. Vaughan. Book review: Statistical Methods and Concepts, by M. N. Das. Journal of the Royal Statistical Society. Series D (The Statistician), 40(4):455–456, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2348739.

Shoesmith:1991:BRB

[182] Eddie Shoesmith. Book review: Maximum-Entropy and Bayesian Methods in Science and Engineering. Vol. 1, Foundations, by Gary J. Erickson; C. Ray Smith; Maximum-Entropy and Bayesian Methods in Science and Engineering. Vol. 2, Applications, by Gary J. Erickson; C. Ray Smith. Journal of the Royal Statistical Society. Series D (The Statistician), 40(4):456– 457, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348740.

Lendrem:1991:BRB

[183] Dennis Lendrem. Book review: Categorical Data Analysis, by A. Agresti. Journal of the Royal Statistical Society. Series D (The Statistician), 40(4): 457, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348741.

Ahmad:1991:BRB

[184] R. Ahmad. Book review: Real Analysis and Probability, by R. M. Dudley. Journal of the Royal Statistical Society. Series D (The Statistician), 40(4):457–458, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348742.

Iyer:1991:BRBc

[185] R. Iyer. Book review: Applied Logistic Regression, by David W. Hosmer; Stanley Lemeshow. Journal of the Royal Statistical Society. Series D (The Statistician), 40(4):458, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2348743.

Collins:1991:BRB

[186] Alan J. Collins. Book review: Modern Methods of Data Analysis, by John Fox; J. Scott Long. Journal of the Royal Statistical Society. Series D (The Statistician), 40(4):458–459, ???? 1991. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2348744.

Legg:1991:BRB

[187] L. Legg. Book review: Understanding Statistics, by Lyman Ott; William Mendenhall. Journal of the Royal Statistical Society. Series D (The Statistician), 40(4):459–460, ???? 1991. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2348745.

Ramachandran:1991:BRB

[188] G. Ramachandran. Book review: Probability and Statistics for Engineers, by R. L. Scheaffer; J. T. McClave. Journal of the Royal Statistical Society. Series D (The Statistician), 40 (4):460–461, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2348746.

Richardson:1991:BRB

[189] S. C. Richardson. Book review: Interaction Effects in Multiple Regression, by J. Jacard; R. Turrisi; C. K. Wan. Journal of the Royal Statistical Society. Series D (The Statistician), 40(4):461–462, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348747.

Disney:1991:BRB

[190] John Disney. Book review: Statistical Methods, Experimental Design and Scientific Inference, by R. A. Fisher. Journal of the Royal Statistical Society. Series D (The Statistician), 40 (4):462–463, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348748.

Garrett:1991:BRB

[191] A. D. Garrett. Book review: Principles of Research Data Audit, by Adil E. Shamoo. Journal of the Royal Statistical Society. Series D (The Statistician), 40(4):463, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348749.

Martin:1991:BRBb

[192] R. J. Martin. Book review: Time Series, by M. G. Kendall; J. K. Ord. Journal of the Royal Statistical Society. Series D (The Statistician), 40 (4):463–464, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348750.

Plewis:1991:BRB

[193] Ian Plewis. Book review: Categorical Longitudinal Data, by J. A. Hagenaars. Journal of the Royal Statistical Society. Series D (The Statistician), 40(4):464, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348751.

Howel:1991:BRBc

[194] Denise Howel. Book review: Statistics: A Spectator Sport, by Richard M. Jaeger. Journal of the Royal Statistical Society. Series D (The Statistician), 40(4):465, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2348753.

Sheldon:1991:BRBb

[195] Neil Sheldon. Book review: Introduction to Probability and Mathematical Statistics, by Lee J. Bain; Max Engelhart. Journal of the Royal Statistical Society. Series D (The Statistician), 40(4):465, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2348752.

Reese:1991:BRB

[196] R. A. Reese. Book review: Envisioning Information, by Eward R. Tufte. Journal of the Royal Statistical Society. Series D (The Statistician), 40(4):466, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2348754.

Vaughan:1991:BRBb

[197] T. B. Vaughan. Book review: Writing Strategies: Reaching Diverse Audiences, by Laurel Richardson. Journal of the Royal Statistical Society. Series D (The Statistician), 40(4):466– 467, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348755.

Anonymous:1991:BMd

[198] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 40(4):469, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348756.

Anonymous:1991:VI

[199] Anonymous. Volume information. Journal of the Royal Statistical Society. Series D (The Statistician), 40 (4):471–475, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348722.

Anonymous:1992:FMa

[200] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 41(1):1, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348630.

Dunne:1992:TSS

[201] Adrian Dunne. Time series simulation. Journal of the Royal Statistical Society. Series D (The Statistician), 41(1):3–8, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348631.

Meyer:1992:EAB

[202] Denny Meyer and Paul Fatti. Estimation of audit bounds for net error. Journal of the Royal Statistical Society. Series D (The Statistician), 41(1):9–16, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348632.

Mullins:1992:SAP

[203] Eamonn Mullins and Michael Stuart. Simulation as an aid in practical statistical problem-solving. Journal of the Royal Statistical Society. Series D (The Statistician), 41(1):17–26, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348633.

Mead:1992:RDM

[204] A. Mead. Review of the development of multidimensional scaling methods. Journal of the Royal Statistical Society. Series D (The Statistician), 41(1):27–39, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348634.

Ritschard:1992:RLU

[205] Gilbert Ritschard and Gérard Antille. A robust look at the use of regression diagnostics. Journal of the Royal Statistical Society. Series D (The Statistician), 41(1):41–53, ???? 1992. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348635.

Sahai:1992:DSV

[206] Hardeo Sahai and Satish Misra. Definitions of sample variance: Some teaching problems to be overcome. Journal of the Royal Statistical Society. Series D (The Statistician), 41(1):55–64, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348636.

Saporu:1992:USO

[207] F. W. O. Saporu. The ultimate size of onchocerciasis-‘infected’ simulium vector population. Journal of the Royal Statistical Society. Series D (The Statistician), 41(1):65–70, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348637.

Gan:1992:CCC

[208] F. F. Gan. CUSUM control charts under linear drift. Journal of the Royal Statistical Society. Series D (The Statistician), 41(1):71–84, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348638.

Bunday:1992:ACR

[209] Brian D. Bunday and Victor A. Kiri. Analysis of censored recidivism data using a proportional hazards-type model. Journal of the Royal Statistical Society. Series D (The Statistician), 41(1):85–96, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348639.

Rocke:1992:CRC

[210] David M. Rocke. X̄_{Q} and R_{Q} charts: Robust control charts. Journal of the Royal Statistical Society. Series D (The Statistician), 41(1):97–104, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348640.

Fielding:1992:BTD

[211] W. J. Fielding and F. B. Lauckner. Biometry training in developing countries: Caribbean experiences. Journal of the Royal Statistical Society. Series D (The Statistician), 41 (1):105–111, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348641.

Cowpertwait:1992:CPM

[212] Paul S. P. Cowpertwait and Trevor F. Cox. Clustering population means under heterogeneity of variance with an application to a rainfall time series problem. Journal of the Royal Statistical Society. Series D (The Statistician), 41(1):113–121, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348642.

Walters:1992:C

[213] D. E. Walters. Correspondence. Journal of the Royal Statistical Society. Series D (The Statistician), 41(1):123– 124, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348643.

Lynn:1992:BRBa

[214] Peter Lynn. Book review: Elementary Survey Sampling, by R. L. Scheaffer; W. Mendenhall; L. Ott. Journal of the Royal Statistical Society. Series D (The Statistician), 41(1):125, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348644.

Kemp:1992:BRBa

[215] A. W. Kemp. Book review: Continuous Bivariate Distributions, Emphasising Applications, by T. P. Hutchinson; C. D. Lai. Journal of the Royal Statistical Society. Series D (The Statistician), 41(1):125–127, ???? 1992. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348645.

McGovern:1992:BRB

[216] Bernadette McGovern. Book review: Mathematical Programming: Optimization Models for Business and Management Decision Making, by M. Wisniewski; T. Dacre. Journal of the Royal Statistical Society. Series D (The Statistician), 41(1):127, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348646.

Marriott:1992:BRB

[217] Alan Marriott. Book review: Statistics for Management and Economics: A Systematic Approach, by G. Keller; B. Warrack; H. Bartell. Journal of the Royal Statistical Society. Series D (The Statistician), 41(1):128, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348647.

Clarke:1992:BRB

[218] G. M. Clarke. Book review: Numerical Analysis, by R. L. Burden; J. D. Faires. Journal of the Royal Statistical Society. Series D (The Statistician), 41(1):128–129, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348648.

Gudgeon:1992:BRBa

[219] Andrew C. Gudgeon. Book review: Principles of Test Theories, by Hoi K. Suen. Journal of the Royal Statistical Society. Series D (The Statistician), 41(1):129–130, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2348649.

Porter:1992:BRB

[220] Max Porter. Book review: Statistical Methods for Medical Investigations, by B. S. Everitt. Journal of the Royal Statistical Society. Series D (The Statistician), 41(1):130, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2348650.

Selby:1992:BRB

[221] Michael J. P. Selby. Book review: Probability for Applications, by Paul E. Pfeiffer. Journal of the Royal Statistical Society. Series D (The Statistician), 41(1):130–131, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2348651.

Firbank:1992:BRBa

[222] Les Firbank. Book review: Intermediate Statistics: A Modern Approach, by James Stevens. Journal of the Royal Statistical Society. Series D (The Statistician), 41(1):131– 132, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348652.

Anonymous:1992:BMa

[223] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 41(1):??, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348653.

Anonymous:1992:FMb

[224] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 41(2):131, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348245.

Congdon:1992:AGL

[225] Peter Congdon. Aspects of general linear modelling of migration. Journal of the Royal Statistical Society. Series D (The Statistician), 41(2):133– 153, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348246.

Sproule:1992:GBP

[226] R. A. Sproule. The general Bernoulli probability density function and the properties of its central moments. Journal of the Royal Statistical Society. Series D (The Statistician), 41(2):155– 159, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348247.

Kochar:1992:NCS

[227] Subhash C. Kochar. A note on characterization of symmetry about a point. Journal of the Royal Statistical Society. Series D (The Statistician), 41 (2):161–163, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348248.

Bacon-Shone:1992:BEF

[228] John Bacon-Shone. Bayesian estimation of finite subpopulation size. Journal of the Royal Statistical Society. Series D (The Statistician), 41 (2):165–167, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348249.

Sakia:1992:BCT

[229] R. M. Sakia. The Box–Cox transformation technique: A review. Journal of the Royal Statistical Society. Series D (The Statistician), 41(2):169– 178, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348250.

Glasbey:1992:RMF

[230] C. A. Glasbey and P. A. Fowler. Regression models fitted using conditional independence to estimate pig fatness from magnetic resonance images. Journal of the Royal Statistical Society. Series D (The Statistician), 41(2):179–184, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348251.

Woodward:1992:FSS

[231] Mark Woodward. Formulae for sample size, power and minimum detectable relative risk in medical studies. Journal of the Royal Statistical Society. Series D (The Statistician), 41(2):185–196, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348252.

Nuamah:1992:ECS

[232] Nicholas N. N. N. Nuamah. The error and covariance structures of the mean approach model of pooled crosssection and time series data. Journal of the Royal Statistical Society. Series D (The Statistician), 41(2):197– 207, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348253.

Moussa:1992:NPR

[233] Mohamed A. A. Moussa and Mohamed Y. Cheema. Non-parametric regression in curve fitting. Journal of the Royal Statistical Society. Series D (The Statistician), 41(2):209– 225, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348254.

Scallan:1992:MLE

[234] A. J. Scallan. Maximum likelihood estimation for a normal/laplace mixture distribution. Journal of the Royal Statistical Society. Series D (The Statistician), 41(2):227–231, ???? 1992. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348255.

Holst:1992:IDB

[235] Erik Holst and Jytte Molin Christensen. Intervals for the description of the biological level of a trace element in a reference population. Journal of the Royal Statistical Society. Series D (The Statistician), 41(2):233– 242, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348256.

Firbank:1992:BRBb

[236] Les Firbank. Book review: Statistics for the Behavioral Sciences, by Sharon L. Weinberg; Kenneth P. Goldberg. Journal of the Royal Statistical Society. Series D (The Statistician), 41(2):243, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348257.

Mellor:1992:BRB

[237] Dale Mellor. Book review: Stationary Stochastic Models, by A. Brandt; P. Franken; B. Lisek. Journal of the Royal Statistical Society. Series D (The Statistician), 41(2):243–244, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348258.

Gough:1992:BRBa

[238] Kevin Gough. Book review: Statistical Theory and Modelling. In Honour of Sir David Cox, FRS, by D. V. Hinkley; N. Reid; E. J. Snell. Journal of the Royal Statistical Society. Series D (The Statistician), 41 (2):244–245, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348259.

Metcalfe:1992:BRB

[239] A. V. Metcalfe. Book review: Stochastic Processes with Applications, by Rabi N. Bhattacharya; Edward C. Waymire. Journal of the Royal Statistical Society. Series D (The Statistician), 41(2):246, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348260.

Day:1992:BRB

[240] Simon Day. Book review: Understanding Significance Testing, by L. B. Mohr. Journal of the Royal Statistical Society. Series D (The Statistician), 41(2):246–247, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348261.

Pewsey:1992:BRBa

[241] Arthur Pewsey. Book review: Randomization and Monte Carlo Methods in Biology, by B. F. J. Manly. Journal of the Royal Statistical Society. Series D (The Statistician), 41 (2):247–248, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348262.

Bell:1992:BRB

[242] J. F. Bell. Book review: Data Quality on Longitudinal Research, by D. Magnusson; L. R. Bergman. Journal of the Royal Statistical Society. Series D (The Statistician), 41(2):248– 249, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348263.

Hayes:1992:BRB

[243] Steven Hayes. Book review: Basic Statistical Computing, by D. Cooke; A. H.

Craven; G. M. Clarke. Journal of the Royal Statistical Society. Series D (The Statistician), 41(2):249, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348264.

Richardson:1992:BRB

[244] S. C. Richardson. Book review: Surveys of Economically Active Population, Employment, Unemployment and Underemployment, by Ralf Hussmanns; Farhad Mehran; Vijay Verma. Journal of the Royal Statistical Society. Series D (The Statistician), 41 (2):249–250, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348265.

Krotki:1992:BRB

[245] Karol J. Krotki. Book review: Theory and Policy Design for Basic Needs Planning: A Case Study of Ecuador, by Rudolf Teekens. Journal of the Royal Statistical Society. Series D (The Statistician), 41(2):250–251, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348266.

Lynn:1992:BRBb

[246] Peter Lynn. Book review: Statistics: A Tool for Social Research, by Joseph F. Healey. Journal of the Royal Statistical Society. Series D (The Statistician), 41(2):251–252, ???? 1992. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348267.

Leitch:1992:BRBa

[247] R. D. Leitch. Book review: Advances in Structural Reliability, by Alfredo C. Lucia. Journal of the Royal Statistical Society. Series D (The Statistician), 41(2):252–254, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348268.

Reilly:1992:BRB

[248] James Reilly. Book review: A History of Probability and Statistics and Their Applications Before 1750, by A. Hald. Journal of the Royal Statistical Society. Series D (The Statistician), 41(2):254, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348269.

Loynes:1992:BRB

[249] R. M. Loynes. Book review: Analysis of Repeated Measures, by M. J. Crowder; D. J. Hand. Journal of the Royal Statistical Society. Series D (The Statistician), 41(2):254–255, ???? 1992. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348270.

McKellar:1992:BRB

[250] John McKellar. Book review: An Introduction to Analysis, by James R. Kirkwood. Journal of the Royal Statistical Society. Series D (The Statistician), 41(2):255–256, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348271.

Prod:1992:BRB

[251] S. J. Prod. Book review: The Double in Nineteenth-Century Fiction, by John Herdman. Journal of the Royal Statistical Society. Series D (The Statistician), 41(2):256, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2348272.

Peace:1992:BRB

[252] Ralph Peace. Book review: A Guide to SPSS for Analysis of Variance, by Gustav Levine. Journal of the Royal Statistical Society. Series D (The Statistician), 41(2):257, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2348273.

Leitch:1992:BRBb

[253] R. D. Leitch. Book review: Modern Mathematical Statistics, by E. J. Dudewicz; S. N. Mishra. Journal of the Royal Statistical Society. Series D (The Statistician), 41(2):257– 258, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348274.

Anonymous:1992:BMb

[254] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 41(2):259– 266, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348275.

Anonymous:1991:FMe

[255] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 41(3):267– 268, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348547.

MacKenzie:1991:F

[256] Gilbert MacKenzie. Foreword. Journal of the Royal Statistical Society. Series D (The Statistician), 41(3):269, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348548.

Haslett:1991:SDA

[257] John Haslett. Spatial data analysischallenges. Journal of the Royal Statistical Society. Series D (The Statistician), 41(3):271–284, ???? 1991. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348549.

Lennox:1991:MML

[258] S. D. Lennox, E. A. Goddall, and C. S. Mayne. A mathematical model of the lactation curve of the dairy cow to incorporate metabolizable energy intake. Journal of the Royal Statistical Society. Series D (The Statistician), 41(3):285–293, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348550.

Donegan:1991:NAA

[259] H. A. Donegan, F. J. Dodd, and T. B. M. McMaster. A new approach to AHP decision-making. Journal of the Royal Statistical Society. Series D (The Statistician), 41(3):295– 302, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348551.

Kelly:1991:RRE

[260] Gabrielle E. Kelly. Robust regression estimators — the choice of tuning constants. Journal of the Royal Statistical Society. Series D (The Statistician), 41(3):303–314, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2348552.

Bond:1991:SRD

[261] Derek Bond and Michael J. Harrison. Some recent developments in econometric model selection and estimation. Journal of the Royal Statistical Society. Series D (The Statistician), 41 (3):315–325, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2348553.

Clayton:1991:GLM

[262] David Clayton. Generalized linear mixed models in biostatistics. Journal of the Royal Statistical Society. Series D (The Statistician), 41(3):327– 328, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348554.

Comiskey:1991:IIE

[263] C. M. Comiskey. Improvements in integral equation models for estimates of the level of HIV infection in Ireland. Journal of the Royal Statistical Society. Series D (The Statistician), 41(3):329, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348555.

Kilpatrick:1991:ETS

[264] S. J. Kilpatrick. Environmental tobacco smoke and lung cancer: A critique of a meta-analysis. Journal of the Royal Statistical Society. Series D (The Statistician), 41(3):331, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348556.

Gillespie:1991:APM

[265] E. S. Gillespie, L. Stewart, and R. N. Wilson. The application of the prediction matrix technique to analyse treatments relating to cancer of the bladder. Journal of the Royal Statistical Society. Series D (The Statistician), 41(3):333, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348557.

McGregor:1991:TSE

[266] Pat McGregor and Ian Cantley. A test of Sen’s entitlement hypothesis. Journal of the Royal Statistical Society. Series D (The Statistician), 41 (3):335–341, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348558.

Cox:1991:RSM

[267] D. R. Cox. The role of statistical models. Journal of the Royal Statistical Society. Series D (The Statistician), 41(3):343, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348559.

MacKenzie:1991:MBF

[268] G. MacKenzie. A model for bivariate failure times subject to right censoring. Journal of the Royal Statistical Society. Series D (The Statistician), 41(3):345, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2348560.

Jackson:1991:ITM

[269] David A. Jackson. Independent trials are a model for disaster. Journal of the Royal Statistical Society. Series D (The Statistician), 41(3):347– 348, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348561.

McClean:1991:PGM

[270] Sally McClean, Pat Lundy, and Wallace Ewart. Predicting the growth of manpower for the northern Ireland software industry. Journal of the Royal Statistical Society. Series D (The Statistician), 41(3):349–356, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348562.

McKeown:1991:PEV

[271] J. J. McKeown and D. Sprevak. Parameter estimation versus curve fitting: ‘new lamps for old’. Journal of the Royal Statistical Society. Series D (The Statistician), 41(3):357– 361, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348563.

Green:1991:SAB

[272] Peter J. Green. Stochastic algorithms and Bayesian inference. Journal of the Royal Statistical Society. Series D (The Statistician), 41(3):363, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348564.

Unwin:1991:HIG

[273] Antony Unwin. How interactive graphics will revolutionize statistical practice. Journal of the Royal Statistical Society. Series D (The Statistician), 41(3):365–369, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348565.

Bradley:1991:HID

[274] Ronan Bradley and John Haslett. High-interaction diagnostics for geostatistical models of spatially referenced data. Journal of the Royal Statistical Society. Series D (The Statistician), 41(3):371–380, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348566.

Wills:1991:NGM

[275] Graham Wills. Novel graphical methods for exploration of spatial data with time information. Journal of the Royal Statistical Society. Series D (The Statistician), 41(3):381, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348567.

Dillon:1991:SEL

[276] M. Dillon. Search for efficient local kriging algorithm. Journal of the Royal Statistical Society. Series D (The Statistician), 41(3):383, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348568.

Anonymous:1991:BMe

[277] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 41(3):385– 386, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348569.

Anonymous:1992:FMc

[278] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 41(4):387, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2349002.

Pham-Gia:1992:SSD

[279] T. Pham-Gia and N. Turkkan. Sample size determination in Bayesian analysis. Journal of the Royal Statistical Society. Series D (The Statistician), 41(4):389–397, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2349003. See com-

ments [280].

Adcock:1992:BAD

[280] C. J. Adcock. Bayesian approaches to the determination of sample sizes for binomial and multinomial sampling — some comments on the paper by Pham-Gia and Turkkan. Journal of the Royal Statistical Society. Series D (The Statistician), 41(4):399– 404, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2349004. See [279].

Freedman:1992:PUR

[281] L. S. Freedman, D. Pee, and D. N. Midthune. The problem of underestimating the residual error variance in forward stepwise regression. Journal of the Royal Statistical Society. Series D (The Statistician), 41(4):405– 412, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2349005.

Brannas:1992:LDP

[282] Kurt Brännäs. Limited dependent Poisson regression. Journal of the Royal Statistical Society. Series D (The Statistician), 41(4):413–423, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2349006.

Ahmed:1992:LSP

[283] S. E. Ahmed. Large-sample pooling procedure for correlation. Journal of the Royal Statistical Society. Series D (The Statistician), 41(4):425– 438, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2349007.

Michels:1992:AKF

[284] Paul Michels. Asymmetric kernel functions in non-parametric regression analysis and prediction. Journal of the Royal Statistical Society. Series D (The Statistician), 41(4):439–454, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2349008.

Schemper:1992:CAS

[285] Michael Schemper. Cox analysis of survival data with non-proportional hazard functions. Journal of the Royal Statistical Society. Series D (The Statistician), 41(4):455–465, ???? 1992. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2349009.

Walker:1992:URL

[286] Esteban Walker and Ralph G. O’Brien. Using restricted least squares to delineate the effects of misspecification in linear models. Journal of the Royal Statistical Society. Series D (The Statistician), 41(4):467–476, ???? 1992. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2349010.

Anonymous:1992:BMc

[287] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 41(4):477– 482, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2349011.

Anonymous:1992:FMd

[288] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 41(5):483, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348913.

Smeeton:1992:E

[289] Nigel Smeeton. From the Editor. Journal of the Royal Statistical Society. Series D (The Statistician), 41(5):485, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348914.

Monod:1992:TFN

[290] Hervé Monod. Two-factor neighbour designs in incomplete blocks for intercropping experiments. Journal of the Royal Statistical Society. Series D (The Statistician), 41(5):487–497, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348915.

Valentine:1992:EAB

[291] Harry T. Valentine, Timothy G. Gregoire, and George M. Furnival. Estimation of the aggregate bole volume of a population of trees by stratified, two-stage, probability sampling. Journal of the Royal Statistical Society. Series D (The Statistician), 41 (5):499–507, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348916.

Gregoire:1992:IEC

[292] Timothy G. Gregoire and Mike Williams. Identifying and evaluating the components of non-measurement error in the application of standard volume equations. Journal of the Royal Statistical Society. Series D (The Statistician), 41(5):509–518, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348917.

Agha:1992:SAD

[293] M. Agha and M. A. Bussell. Statistical analysis of the discrimination between salt Marsh and mudflats using visible and infrared thematic mapper data. Journal of the Royal Statistical Society. Series D (The Statistician), 41(5):519–523, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2348918.

Beetson:1992:EPS

[294] Trevor Beetson, Marks Nester, and Jerry Vanclay. Enhancing a permanent sample plot system in natural forests. Journal of the Royal Statistical Society. Series D (The Statistician), 41(5):525–538, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2348919.

Richardson:1992:SLM

[295] S. Richardson, C. Guihenneuc, and V. Lasserre. Spatial linear models with autocorrelated error structure. Journal of the Royal Statistical Society. Series D (The Statistician), 41 (5):539–557, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2348920.

Henery:1992:ETM

[296] Robert J. Henery. An extension to the Thurstone–Mosteller model for chess. Journal of the Royal Statistical Society. Series D (The Statistician), 41 (5):559–567, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348921.

Copas:1992:ECL

[297] J. B. Copas. Exact confidence limits for binomial proportions — Brenner & Quan revisited. Journal of the Royal Statistical Society. Series D (The Statistician), 41(5):569–572, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348922. See [71].

Congdon:1992:DEM

[298] Peter Congdon and Ian McCallum. A demo-educational model for forecasting school rolls for localities. Journal of the Royal Statistical Society. Series D (The Statistician), 41(5):573– 590, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348923.

Cox:1992:CPM

[299] Trevor F. Cox and Paul S. P. Cowpertwait. Clustering population means under heterogeneity of variance. Journal of the Royal Statistical Society. Series D (The Statistician), 41(5):591– 598, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348924.

Holt:1992:CMW

[300] D. Holt and D. Elliot. Corrigendum: Methods of weighting for unit nonresponse. Journal of the Royal Statistical Society. Series D (The Statistician), 41(5):599, ???? 1992. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2348925. See

[146].

Anonymous:1992:BMd

[301] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 41(5):601, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348926.

Anonymous:1992:VIa

[302] Anonymous. Volume information. Journal of the Royal Statistical Society. Series D (The Statistician), 41 (5):603–608, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2348912.

Anonymous:1993:FMa

[303] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 42(1):1, ???? 1993. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2348104.

Simonoff:1993:RIB

[304] Jeffrey S. Simonoff. The relative importance of bias and variability in the estimation of the variance of a statistic. Journal of the Royal Statistical Society. Series D (The Statistician), 42(1):3–7, ???? 1993. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2348105.

Phillips:1993:CTM

[305] Michael J. Phillips. Contingency tables with missing data. Journal of the Royal Statistical Society. Series D (The Statistician), 42(1):9–18, ???? 1993. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348106.

Ghosh:1993:REP

[306] Subir Ghosh and Edna Derderian. Robust experimental plan and its role in determining robust design against noise factors. Journal of the Royal Statistical Society. Series D (The Statistician), 42(1):19–28, ???? 1993. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348107.

Wilcox:1993:CBM

[307] Rand R. Wilcox. Comparing the biweight midvariances of two independent groups. Journal of the Royal Statistical Society. Series D (The Statistician), 42(1):29–35, ???? 1993. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348108.

Royston:1993:TTN

[308] Patrick Royston. A toolkit for testing for non-normality in complete and censored samples. Journal of the Royal Statistical Society. Series D (The Statistician), 42(1):37–43, ???? 1993. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348109. See

corrigendum [1078].

Bartlett:1993:LMP

[309] R. F. Bartlett. Linear modelling of Pearson’s product moment correlation coefficient: An application of Fisher’s z-transformation. Journal of the Royal Statistical Society. Series D (The Statistician), 42(1):45–53, ???? 1993. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348110.

Thomas:1993:URS

[310] G. E. Thomas and S. S. Kiwanga. Use of ranking and scoring methods in the analysis of ordered categorical data from factorial experiments. Journal of the Royal Statistical Society. Series D (The Statistician), 42 (1):55–67, ???? 1993. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348111.

Shanmugam:1993:BRBa

[311] R. Shanmugam. Book review: The Multivariate Normal Distribution, by Y. L. Tong. Journal of the Royal Statistical Society. Series D (The Statistician), 42(1):69, ???? 1993. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348112.

Percy:1993:BRB

[312] David F. Percy. Book review: Estimating Functions, by V. P. Godambe. Journal of the Royal Statistical Society. Series D (The Statistician), 42(1):69–70, ???? 1993. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348113.

Howel:1993:BRBa

[313] D. Howel. Book review: Modelling Binary Data, by D. Collett. Journal of the Royal Statistical Society. Series D (The Statistician), 42(1):70–71, ???? 1993. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348114.

North:1993:BRB

[314] P. M. North. Book review: Time Series: Theory and Methods, by Peter J. Brockwell; Richard A. Davis. Journal of the Royal Statistical Society. Series D (The Statistician), 42 (1):71–72, ???? 1993. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348115.

Grieve:1993:BRB

[315] A. P. Grieve. Book review: ‘Student’ — A Statistical Biography of William Sealy Gosset: Based on the Writings of E. S. Pearson, by R. L. Plackett; G. A. Barnard. Journal of the Royal Statistical Society. Series D (The Statistician), 42(1):72, ???? 1993. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348116.

Lawson:1993:BRB

[316] A. Lawson. Book review: Statistics for Spatial Data, by N. Cressie. Journal of the Royal Statistical Society. Series D (The Statistician), 42(1):73, ???? 1993. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348117.

Woodward:1993:BRBa

[317] P. W. Woodward. Book review: Computation for the Analysis of Designed Experiments, by R. M. Heiberger. Journal of the Royal Statistical Society. Series D (The Statistician), 42(1):73–74, ???? 1993. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348118.

Rees:1993:BRB

[318] D. G. Rees. Book review: Statistics, by D. Freedman; R. Pisani; R. Purves; A. Adhikari. Journal of the Royal Statistical Society. Series D (The Statistician), 42(1):74–75, ???? 1993. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348119.

Kemp:1993:BRB

[319] A. W. Kemp. Book review: Statistical Inference and Analysis: Selected Correspondence of R. A. Fisher, by R. A. Fisher; J. H. Bennett. Journal of the Royal Statistical Society. Series D (The Statistician), 42(1):75–76, ???? 1993. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348120.

Jolliffe:1993:BRB

[320] I. T. Jolliffe. Book review: A User’s Guide to Principal Components, by J. Edward Jackson. Journal of the Royal Statistical Society. Series D (The Statistician), 42(1):76–77, ???? 1993. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348121.

Howel:1993:BRBb

[321] D. Howel. Book review: Statistics in Health and Nutrition, by K. Visweswara Rao; G. Radhaiah; V. Narayana. Journal of the Royal Statistical Society. Series D (The Statistician), 42(1):77, ???? 1993. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348122.

Metcalfe:1993:BRB

[322] A. V. Metcalfe. Book review: Nonlinear System Analysis and Identification from Random Data, by Julius S. Bendat. Journal of the Royal Statistical Society. Series D (The Statistician), 42(1):77–78, ???? 1993. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348123.

Iyer:1993:BRB

[323] R. Iyer. Book review: Statistical Models in Behavioural Research, by William K. Estes. Journal of the Royal Statistical Society. Series D (The Statistician), 42(1):78, ???? 1993. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348124.

Smith:1993:BRB

[324] Gordon L. Smith. Book review: Ratio Scaling of Psychological Magnitude: In Honor of the Memory of S. S. Stevens, by Stanley J. Bolanowski,; George A. Geschieder. Journal of the Royal Statistical Society. Series D (The Statistician), 42(1):78–79, ???? 1993. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348125.

Bullard:1993:BRB

[325] Peter Bullard. Book review: Economic Forecasting: An Introduction, by K. Holden; D. A. Peel; J. L. Thompson. Journal of the Royal Statistical Society. Series D (The Statistician), 42(1):79, ???? 1993. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2348126.

Woodward:1993:BRBb

[326] P. W. Woodward. Book review: Extreme Value Theory in Engineering, by E. Castillo. Journal of the Royal Statistical Society. Series D (The Statistician), 42(1):79–80, ???? 1993. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2348127.

Shanmugam:1993:BRBb

[327] R. Shanmugam. Book review: Errors, Bias and Uncertainties in Astronomy, by Carlos Jascheck; Fionn Murtagh. Journal of the Royal Statistical Society. Series D (The Statistician), 42(1):80–82, ???? 1993. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2348128.

Anonymous:1993:BMa

[328] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 42(1):83–88, ???? 1993. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348129.

Anonymous:1993:FMb

[329] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 42(2):89, ???? 1993. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348972.

Adcock:1993:IBP

[330] C. J. Adcock. An improved Bayesian procedure for calculating sample sizes in multinomial sampling. Journal of the Royal Statistical Society. Series D (The Statistician), 42(2):91–95, ???? 1993. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348973.

Sutradhar:1993:THV

[331] Brajendra C. Sutradhar and Estela Bee Dagum. Testing for homogeneous variations in periodic data. Journal of the Royal Statistical Society. Series D (The Statistician), 42(2):97–106, ???? 1993. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348974.

Kmietowicz:1993:SAI

[332] Z. W. Kmietowicz and H. Ding. Statistical analysis of income distribution in the Jiangsu Province of China. Journal of the Royal Statistical Society. Series D (The Statistician), 42 (2):107–121, ???? 1993. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348975.

Adjibolosoo:1993:EPH

[333] Senyo B-S. K. Adjibolosoo. Estimation of parameters of heteroscedastic error models under various hypothesized error structures. Journal of the Royal Statistical Society. Series D (The Statistician), 42(2):123–133, ???? 1993. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348976.

Thiagarajah:1993:TAO

[334] K. Thiagarajah and S. R. Paul. C(α) tests for the analysis of one-way layout of data having extreme value distribution. Journal of the Royal Statistical Society. Series D (The Statistician), 42(2):135–143, ???? 1993. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348977.

Dunsmore:1993:OMC

[335] Ian R. Dunsmore and Nick J. Ward. Optimal mixture choice for binary response models. Journal of the Royal Statistical Society. Series D (The Statistician), 42(2):145–152, ???? 1993. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348978.

McCulloch:1993:FRM

[336] R. E. McCulloch. Fitting regression models with unknown transformations using dynamic graphics. Journal of the Royal Statistical Society. Series D (The Statistician), 42(2):153– 160, ???? 1993. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348979.

Koziol:1993:PPA

[337] James A. Koziol. Probability plots for assessing multivariate normality. Journal of the Royal Statistical Society. Series D (The Statistician), 42 (2):161–173, ???? 1993. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348980.

Buning:1993:CSR

[338] H. Büning. A coefficient of stability of rank tests in the one-sample case. Journal of the Royal Statistical Society. Series D (The Statistician), 42(2):175–180, ???? 1993. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348981.

Yannaros:1993:AIC

[339] Nikos Yannaros. Analyzing incomplete count data. Journal of the Royal Statistical Society. Series D (The Statistician), 42(2):181–187, ???? 1993. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348982.

Pidgen:1993:BRB

[340] N. D. Pidgen. Book review: The Comparative Method, by C. R. Ragin. Journal of the Royal Statistical Society. Series D (The Statistician), 42(2):189, ???? 1993. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348983.

Watkins:1993:BRB

[341] A. J. Watkins. Book review: Linear Structures, by J. R. Magnus. Journal of the Royal Statistical Society. Series D (The Statistician), 42(2):190, ???? 1993. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348984.

Green:1993:BRB

[342] David Green. Book review: Statistical Inference for Spatial Processes, by B. D. Ripley. Journal of the Royal Statistical Society. Series D (The Statistician), 42(2):190–191, ???? 1993. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348985.

Edwards:1993:BRB

[343] Phil Edwards. Book review: Integrated Land Use and Transport Modelling: Decision Chains and Hierarchies, by Thomás de la Barra. Journal of the Royal Statistical Society. Series D (The Statistician), 42(2):191– 193, ???? 1993. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348986.

Prys-Williams:1993:BRB

[344] Allan G. Prys-Williams. Book review: Review of Panel Surveys, by D. Kaspryyk; G. Duncan; G. Kalton; M. P. Singh. Journal of the Royal Statistical Society. Series D (The Statistician), 42(2):193–194, ???? 1993. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348987.

Collins:1993:BRB

[345] A. J. Collins. Book review: Statistics in Scientific Investigation, by Glen McPherson. Journal of the Royal Statistical Society. Series D (The Statistician), 42(2):194, ???? 1993. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348988.

Garrett:1993:BRB

[346] A. D. Garrett. Book review: Design Sensitivity: Statistical Power for Experimental Research, by Mark W. Lipsey. Journal of the Royal Statistical Society. Series D (The Statistician), 42(2):195, ???? 1993. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348989.

Coshall:1993:BRB

[347] John Coshall. Book review: Applied Nonparametric Regression, by Wolfgang Hardle. Journal of the Royal Statistical Society. Series D (The Statistician), 42(2):195–196, ???? 1993. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348990.

Jeffers:1993:BRB

[348] J. N. R. Jeffers. Book review: Graphical Models in Applied Multivariate Statistics, by Joe Whittaker. Journal of the Royal Statistical Society. Series D (The Statistician), 42(2):196– 197, ???? 1993. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348991.

McKellar:1993:BRB

[349] John McKellar. Book review: Statistics in Medicine, by F. Dunstan; J.

Pickles. Journal of the Royal Statistical Society. Series D (The Statistician), 42(2):197–198, ???? 1993. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348992.

Szuster:1993:BRB

[350] Fearnley Szuster. Book review: Drawing General Population Samples in Great Britain, by Peter Lynn; Denise Lievesley. Journal of the Royal Statistical Society. Series D (The Statistician), 42(2):198, ???? 1993. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348993.

Shoesmith:1993:BRB

[351] Eddie Shoesmith. Book review: A History of Inverse Probability from Thomas Bayes to Karl Pearson, by Andrew I. Dale. Journal of the Royal Statistical Society. Series D (The Statistician), 42(2):199, ???? 1993. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348994.

Lynn:1993:BRB

[352] Peter Lynn. Book review: Sampling of Populations: Methods and Applications, by Paul S. Levy; Stanley Lemeshow. Journal of the Royal Statistical Society. Series D (The Statistician), 42(2):199–200, ???? 1993. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348995.

Apostolides:1993:BRB

[353] M. G. Apostolides. Book review: The Challenge to Government: Management of a Capitalist Economy, by H. W. Robinson. Journal of the Royal Statistical Society. Series D (The Statistician), 42(2):200, ???? 1993. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348996.

Mabbett:1993:BRB

[354] Alan Mabbett. Book review: Sharing Social Science Data — Advantages and Challenges, by Joan E. Sieber. Journal of the Royal Statistical Society. Series D (The Statistician), 42 (2):200–201, ???? 1993. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348997.

Trustrum:1993:BRB

[355] Kathleen Trustrum. Book review: Regression Diagnostics: An Introduction, by John Fox. Journal of the Royal Statistical Society. Series D (The Statistician), 42(2):201, ???? 1993. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348998.

Jones:1993:BRB

[356] M. C. Jones. Book review: Nonparametric and Semiparametric Methods in Econometrics and Statistics, by W. A. Barnett; J. Powell; G. E. Tauchen. Journal of the Royal Statistical Society. Series D (The Statistician), 42 (2):201–202, ???? 1993. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348999.

Wooff:1993:BRB

[357] David Wooff. Book review: Statistics: Concepts and Controversies, by David S. Moore. Journal of the Royal Statistical Society. Series D (The Statistician), 42(2):202–203, ???? 1993. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2349000.

Anonymous:1993:BMb

[358] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 42(2):204– 207, ???? 1993. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2349001.

Anonymous:1992:FMe

[359] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 42(3):209– 210, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348794.

Anonymous:1992:F

[360] Anonymous. Foreward. Journal of the Royal Statistical Society. Series D (The Statistician), 42(3):211, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2348795.

McEntee:1992:CCP

[361] E. F. McEntee. A control chart procedure for monitoring quality in terms of consumer appraisal. Journal of the Royal Statistical Society. Series D (The Statistician), 42(3):213, ???? 1992.

CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348796.

Trevanich:1992:ECC

[362] Anothai Trevanich and Patrick Bourke. EWMA control charts using attributes data. Journal of the Royal Statistical Society. Series D (The Statistician), 42(3):215, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348797.

Porter:1992:RSI

[363] M. A. Porter. The role of the statistician in industry. Journal of the Royal Statistical Society. Series D (The Statistician), 42(3):217–227, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348798.

Walsh:1992:IAE

[364] Brendan Walsh. Improving the accuracy of estimated electricity bills with the prospect of changing from two-monthly meter reading to fourmonthly meter reading. Journal of the Royal Statistical Society. Series D (The Statistician), 42(3):229, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348799.

Gillespie:1992:AMO

[365] E. S. Gillespie. An application of multivariate outlier detection in assessing family characteristics for bank advertisements. Journal of the Royal Statistical Society. Series D (The Statistician), 42(3):231–235, ???? 1992. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2348800.

McLoughlin:1992:ATS

[366] E. M. McLoughlin, E. A. Goodall, F. D. Menzies, and S. G. McIlroy. The application of time series analysis to assess and control the prevalence of cirrhosis of pigs within northern Ireland. Journal of the Royal Statistical Society. Series D (The Statistician), 42 (3):237–247, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2348801.

Lennox:1992:MMM

[367] S. D. Lennox, E. A. Goodall, and C. S. Mayne. Mathematical modelling of milk constituent production. Journal of the Royal Statistical Society. Series D (The Statistician), 42 (3):249–260, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2348802.

MacKenzie:1992:STC

[368] Gilbert MacKenzie. Space-time clustering revisited. Journal of the Royal Statistical Society. Series D (The Statistician), 42(3):261, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2348803.

McClean:1992:PLS

[369] Sally McClean and Peter Millard. Patterns of length of stay after admission in geriatric medicine: An event history approach. Journal of the Royal Statistical Society. Series D (The Statistician), 42(3):263–274, ???? 1992. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348804.

Kelly:1992:MAU

[370] Alan Kelly. Meta-analysis unconfounded: A case history. Journal of the Royal Statistical Society. Series D (The Statistician), 42(3):275, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348805.

Jackson:1992:SMI

[371] David A. Jackson. Some models for index betting on sports. Journal of the Royal Statistical Society. Series D (The Statistician), 42(3):277, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348806.

Fieller:1992:AOS

[372] N. R. J. Fieller. Archaeostatistics: Old statistics in ancient contexts. Journal of the Royal Statistical Society. Series D (The Statistician), 42(3):279– 295, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348807.

Bond:1992:ABS

[373] Derek Bond and Denise Lievesley. Address-based sampling in northern Ireland. Journal of the Royal Statistical Society. Series D (The Statistician), 42(3):297–304, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348808.

Barry:1992:SDV

[374] Dan Barry. Some difficulties with pvalues. Journal of the Royal Statistical Society. Series D (The Statistician), 42(3):305, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2348809.

Reilly:1992:DAU

[375] Marie Reilly. Data analysis using hot deck multiple imputation. Journal of the Royal Statistical Society. Series D (The Statistician), 42(3):307– 313, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348810.

Tesarikova:1992:PMS

[376] Eva Tesarikova. Probabilistic models of school achievement tests with double-choice response and their statistical analysis. Journal of the Royal Statistical Society. Series D (The Statistician), 42(3):315, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2348811.

Yandell:1992:SST

[377] Brian S. Yandell. Smoothing splines — a tutorial. Journal of the Royal Statistical Society. Series D (The Statistician), 42(3):317–319, ???? 1992. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2348812.

Pawitan:1992:PWL

[378] Yudi Pawitan. Penalized Whittlelikelihood estimates of spectral density. Journal of the Royal Statistical Society. Series D (The Statistician), 42(3):321, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348813.

Garvin:1992:BDT

[379] J. Garvin and S. J. McClean. The binormal distribution: Theory and application. Journal of the Royal Statistical Society. Series D (The Statistician), 42(3):323, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348814.

Shanmugam:1992:BRB

[380] Ramalingam Shanmugam. Book review: Henry Theil’s Contributions to Economics and Econometrics Volumes, by Baldev Raj; Johan Koerts. Journal of the Royal Statistical Society. Series D (The Statistician), 42(3):325, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348815.

Coad:1992:BRB

[381] D. S. Coad. Book review: Applied Multivariate Data Analysis, by B. S. Everitt; G. Dunn. Journal of the Royal Statistical Society. Series D (The Statistician), 42(3):325– 326, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348816.

Chatfield:1992:BRB

[382] Chris Chatfield. Book review: Computing and Graphics in Statistics, by A. Buja; P. A. Tukey. Journal of the Royal Statistical Society. Series D (The Statistician), 42(3):326–327, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348817.

Shirley:1992:BRB

[383] Eryl Shirley. Book review: Computer Simulation Application: An Introduction, by Marcia Lynn Whicker; Lee Siegelman. Journal of the Royal Statistical Society. Series D (The Statistician), 42(3):327–328, ???? 1992. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2348818.

Shade:1992:BRB

[384] John Shade. Book review: The Engineering Statistician’s Guide to Continuous Bivariate Distributions, by T. P. Hutchinson; C. D. Lai. Journal of the Royal Statistical Society. Series D (The Statistician), 42(3):328, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2348819.

Lynn:1992:BRBc

[385] Peter Lynn. Book review: Multiple Comparisons for Researchers, by Larry E. Toothaker. Journal of the Royal Statistical Society. Series D (The Statistician), 42(3):329, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2348820.

Nason:1992:BRB

[386] Guy Nason. Book review: Probability Measures on Groups, by Herbert Heyer. Journal of the Royal Statistical Society. Series D (The Statistician), 42(3):329–330, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348821.

Brooks:1992:BRBa

[387] Steve Brooks. Book review: Statistical Inference: Theory and Practice, by Tadeusz Bromek; Elzbieta Pleszczyńska. Journal of the Royal Statistical Society. Series D (The Statistician), 42(3):330, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348822.

Sprent:1992:BRB

[388] P. Sprent. Book review: Topics in Statistical Methodology, by Suddhendu Biswas. Journal of the Royal Statistical Society. Series D (The Statistician), 42(3):330–331, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348823.

Pridmore:1992:BRB

[389] W. A. Pridmore. Book review: A Dictionary of Statistical Terms, by F. H. C. Marriott. Journal of the Royal Statistical Society. Series D (The Statistician), 42(3):331–332, ???? 1992. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348824.

Anonymous:1992:BMe

[390] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 42(3):333– 336, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348825.

Anonymous:1992:FMf

[391] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 42(4):337, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348467.

Fearn:1992:Fa

[392] T. Fearn and A. O’Hagan. Foreword. Journal of the Royal Statistical Society. Series D (The Statistician), 42(4):339, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348468.

Chaloner:1992:GEP

[393] Kathryn Chaloner, Timothy Church, Thomas A. Louis, and John P. Matts. Graphical elicitation of a prior distribution for a clinical trial. Journal of the Royal Statistical Society. Series D (The Statistician), 42(4):341– 353, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348469.

Carlin:1992:BAM

[394] Bradley P. Carlin, Kathryn Chaloner, Timothy Church, Thomas A. Louis, and John P. Matts. Bayesian approaches for monitoring clinical trials with an application to toxoplasmic encephalitis prophylaxis. Journal of the Royal Statistical Society. Series D (The Statistician), 42(4):355– 367, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348470.

Cowell:1992:BNA

[395] R. G. Cowell, A. P. Dawid, T. A. Hutchinson, S. Roden, and D. J. Spiegelhalter. Bayesian networks for the analysis of drug safety. Journal of the Royal Statistical Society. Series D (The Statistician), 42(4):369– 384, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348471.

Ashby:1992:SBA

[396] Deborah Ashby, Jane L. Hutton, and Magnus A. McGee. Simple Bayesian analyses for case-control studies in cancer epidemiology. Journal of the Royal Statistical Society. Series D (The Statistician), 42(4):385–397, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348472.

Hutton:1992:BSS

[397] J. L. Hutton and R. G. Owens. Bayesian sample size calculations and prior beliefs about Child sexual abuse. Journal of the Royal Statistical Society. Series D (The Statistician), 42 (4):399–404, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348473.

Wild:1992:HBA

[398] Pascal Wild, Daniel Commenges, and Brigitte Etcheverry. A hierarchical Bayesian approach to the backcalculation of numbers of HIV-infected subjects. Journal of the Royal Statistical Society. Series D (The Statistician), 42(4):405–414, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2348474.

Kadane:1992:SBA

[399] Joseph B. Kadane. Subjective Bayesian analysis for surveys with missing data. Journal of the Royal Statistical Society. Series D (The Statistician), 42 (4):415–426, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2348475.

Berger:1992:PRP

[400] James O. Berger and Ming-Hui Chen. Predicting retirement patterns: Prediction for a multinomial distribution with constrained parameter space. Journal of the Royal Statistical Society. Series D (The Statistician), 42 (4):427–443, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2348476.

Goldstein:1992:PUI

[401] Michael Goldstein, Malcolm Farrow, and Takis Spiropoulos. Prediction under the influence: Bayes linear influence diagrams for prediction in a large brewery. Journal of the Royal Statistical Society. Series D (The Statistician), 42(4):445–459, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2348477.

Woodward:1992:ABM

[402] Philip W. Woodward and John C. Naylor. An application to Bayesian methods in SPC. Journal of the Royal Statistical Society. Series D (The Statistician), 42(4):461–469, ???? 1992. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348478.

Kulhavy:1992:IBP

[403] Rudolf Kulhavý. Implementation of Bayesian parameter estimation in adaptive control and signal processing. Journal of the Royal Statistical Society. Series D (The Statistician), 42(4):471–482, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348479.

Buck:1992:DCS

[404] C. E. Buck, C. D. Litton, and D. A. Stephens. Detecting a change in the shape of a prehistoric corbelled tomb. Journal of the Royal Statistical Society. Series D (The Statistician), 42 (4):483–490, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348480.

Pettit:1992:IAO

[405] L. I. Pettit. Inferences about ordered parameters — an astronomical problem. Journal of the Royal Statistical Society. Series D (The Statistician), 42(4):491–496, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348481.

Anonymous:1992:BMf

[406] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 42(4):??, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2348482.

Anonymous:1992:FMg

[407] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 42(5):497, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348673.

Fearn:1992:Fb

[408] T. Fearn and A. O’Hagan. Foreword. Journal of the Royal Statistical Society. Series D (The Statistician), 42(5):499, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2348674.

Smith:1992:BUA

[409] Jim Smith and Simon French. Bayesian updating of atmospheric dispersion models for use after an accidental release of radioactivity. Journal of the Royal Statistical Society. Series D (The Statistician), 42(5):501–511, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348675.

Rotondi:1992:BAT

[410] R. Rotondi, J. J. Egozcue, and G. Zonno. Bayesian approach to the treatment of uncertainty in seismic data. Journal of the Royal Statistical Society. Series D (The Statistician), 42(5):513–523, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348676.

Ibrahim:1992:BOE

[411] Kamarulzaman Bin Ibrahim and A. V. Metcalfe. Bayesian overview for evaluation of mini-roundabouts as a road safety measure. Journal of the Royal Statistical Society. Series D (The Statistician), 42(5):525–540, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/

stable/2348677. See correction [564].

Bernardo:1992:BAP

[412] José M. Bernardo and Javier Muñoz. Bayesian analysis of population evolution. Journal of the Royal Statistical Society. Series D (The Statistician), 42(5):541–550, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348678.

Kass:1992:BFP

[413] Robert E. Kass. Bayes factors in practice. Journal of the Royal Statistical Society. Series D (The Statistician), 42(5):551–560, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348679.

DelaFuente:1992:APJ

[414] E. I. De la Fuente, H. Trujillo, A. R. Ortega, I. Martín, and R. Estarelles. Assessment in probability judgement:

A model for combining sources. Journal of the Royal Statistical Society. Series D (The Statistician), 42(5):561– 570, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348680.

Polasek:1992:BRM

[415] Wolfgang Polasek and Andreas Krause. Bayesian regression model with simple errors in variables structure. Journal of the Royal Statistical Society. Series D (The Statistician), 42(5):571– 580, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348681.

Anonymous:1992:BMg

[416] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 42(5):581, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348683.

Anonymous:1992:VIb

[417] Anonymous. Volume information. Journal of the Royal Statistical Society. Series D (The Statistician), 42 (5):583–589, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348672.

Anonymous:1992:A

[418] Anonymous. Announcement. Journal of the Royal Statistical Society. Series D (The Statistician), 42(5):591, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348682.

Grieve:1992:MYP

[419] Andrew P. Grieve, Alison Meredith, and Trevor J. Sweeting. Measuring your posterior beliefs that you can count properly. Journal of the Royal Statistical Society. Series D (The Statistician), 43(1):1–12, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348928.

Wakefield:1992:ELA

[420] Jon Wakefield. An expected loss approach to the design of dosage regimens via sampling-based methods. Journal of the Royal Statistical Society. Series D (The Statistician), 43(1):13–29, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348929.

West:1992:DMA

[421] Mike West and Dennis A. Turner. Deconvolution of mixtures in analysis of neural synaptic transmission. Journal of the Royal Statistical Society. Series D (The Statistician), 43 (1):31–43, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348930.

duPlessis:1992:IMB

[422] J. L. du Plessis and A. J. Van Der Merwe. Inferences in multivariate Bayesian calibration. Journal of the Royal Statistical Society. Series D (The Statistician), 43(1):45–60, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2348931.

Forster:1992:BAA

[423] Jonathan J. Forster. A Bayesian approach to the analysis of binary crossover data. Journal of the Royal Statistical Society. Series D (The Statistician), 43(1):61–68, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2348932.

Lawson:1992:USG

[424] Andrew B. Lawson. Using spatial Gaussian priors to model heterogeneity in environmental epidemiology. Journal of the Royal Statistical Society. Series D (The Statistician), 43(1):69–76, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348933.

Green:1992:DAT

[425] Edwin J. Green and William E. Strawderman. Determining accuracy of thematic maps. Journal of the Royal Statistical Society. Series D (The Statistician), 43(1):77–85, ???? 1992. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2348934.

Queen:1992:UMD

[426] Catriona M. Queen. Using the multiregression dynamic model to forecast brand sales in a competitive product market. Journal of the Royal Statistical Society. Series D (The Statistician), 43(1):87–98, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348935.

VanBatenburg:1992:BDS

[427] Paul C. Van Batenburg, Anthony O’Hagan, and Ruud H. Veenstra. Bayesian discovery sampling in financial auditing: A hierarchical prior model for substantive test sample sizes. Journal of the Royal Statistical Society. Series D (The Statistician), 43 (1):99–110, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348936.

Ritter:1992:SAS

[428] Christian Ritter. Statistical analysis of spectra from electron spectroscopy for chemical analysis. Journal of the Royal Statistical Society. Series D (The Statistician), 43(1):111– 127, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348937.

Young:1992:BAU

[429] K. D. S. Young. A Bayesian analysis of updated component data. Journal of the Royal Statistical Society. Series D (The Statistician), 43(1):129– 137, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348938.

Armero:1992:PAP

[430] C. Armero and M. J. Bayarri. Prior assessments for prediction in queues. Journal of the Royal Statistical Society. Series D (The Statistician), 43 (1):139–153, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2348939.

Coolen:1992:BEI

[431] F. P. A. Coolen. On Bernoulli experiments with imprecise prior probabilities. Journal of the Royal Statistical Society. Series D (The Statistician), 43(1):155–167, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2348940.

Gilks:1992:LPC

[432] W. R. Gilks, A. Thomas, and D. J. Spiegelhalter. A language and program for complex Bayesian modelling. Journal of the Royal Statistical Society. Series D (The Statistician), 43 (1):169–177, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2348941.

Gilks:1992:ADS

[433] W. R. Gilks, G. O. Roberts, and E. I. George. Adaptive direction sampling. Journal of the Royal Statistical Society. Series D (The Statistician), 43 (1):179–189, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2348942.

Bertolino:1992:RBA

[434] F. Bertolino and W. Racugno. Robust Bayesian analysis in analysis of variance and the χ^{2}-test by using marginal likelihoods. Journal of the Royal Statistical Society. Series D (The Statistician), 43(1):191–201, ???? 1992. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348943.

Chalmers:1992:BRBa

[435] C. P. Chalmers. Book review: Risk: Analysis, Assessment and Management, by J. Ansell; F. Wharton. Journal of the Royal Statistical Society. Series D (The Statistician), 43(1):203, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348944.

Chalmers:1992:BRBb

[436] C. P. Chalmers. Book review: Processing Data — The Survey Example, by L. B. Bourque; V. A. Clark. Journal of the Royal Statistical Society. Series D (The Statistician), 43 (1):203–204, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348945.

Brooks:1992:BRBb

[437] Stephen Brooks. Book review: General Statistics, by W. Chase; F. Brown. Journal of the Royal Statistical Society. Series D (The Statistician), 43 (1):204–205, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348946.

Judge:1992:BRB

[438] E. J. Judge. Book review: The Development of Statistics: Recent Contributions from China, by X. R. Chen; K. T. Fang; C. C. Yang. Journal of the Royal Statistical Society. Series D (The Statistician), 43(1):205, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2348947.

Baker:1992:BRB

[439] Rose Baker. Book review: ARMA Model Identification, by B. S. Choi. Journal of the Royal Statistical Society. Series D (The Statistician), 43(1): 206, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348948.

Howel:1992:BRBa

[440] Denise Howel. Book review: Logit Modelling: Practical Applications, by A. Demaris. Journal of the Royal Statistical Society. Series D (The Statistician), 43(1):206–207, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2348949.

Lounds:1992:BRB

[441] W. S. Lounds. Book review: Understanding Data, by B. H. Erickson; T. A. Nosanchuk. Journal of the Royal Statistical Society. Series D (The Statistician), 43(1):207, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2348950.

Gough:1992:BRBb

[442] Kevin Gough. Book review: The Analysis of Contingency Tables, by B. S. Everitt. Journal of the Royal Statistical Society. Series D (The Statistician), 43(1):207–208, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2348951.

Wisniewski:1992:BRB

[443] Mik Wisniewski. Book review: Statistical Methods, by R. J. Freund; W. J. Wilson. Journal of the Royal Statistical Society. Series D (The Statistician), 43(1):209, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348952.

Apostolides:1992:BRB

[444] M. G. Apostolides. Book review: Statistical Thinking for Managers, by D. Hildebrand; L. Ott. Journal of the Royal Statistical Society. Series D (The Statistician), 43(1):209–210, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348953.

Mabbett:1992:BRB

[445] Alan Mabbett. Book review: Using National Databases in Educational Research, by T. L. Hilton. Journal of the Royal Statistical Society. Series D (The Statistician), 43(1):210– 211, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348954.

Saldanha:1992:BRB

[446] Richard A. Saldanha. Book review: Mathematics in Medicine and the Life Sciences, by F. C. Hoppensteadt; C. S. Peskin. Journal of the Royal Statistical Society. Series D (The Statistician), 43(1):211, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348955.

Gudgeon:1992:BRBb

[447] Andrew C. Gudgeon. Book review: Statistical Methods for Psychology, by David C. Howell. Journal of the Royal Statistical Society. Series D (The Statistician), 43(1):211– 212, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348956.

Mann:1992:BRB

[448] S. L. Mann. Book review: Essentials of Statistical Methods, in 41 Pages, by T. P. Hutchinson. Journal of the Royal Statistical Society. Series D (The Statistician), 43(1):212– 213, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348957.

Ruxton:1992:BRB

[449] Graeme Ruxton. Book review: Elementary Statistics, by R. Johnson. Journal of the Royal Statistical Society. Series D (The Statistician), 43(1): 213, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348958.

Jain:1992:BRB

[450] S. Jain. Book review: Basic Statistics for Laboratories, by W. D. Kelley; T. A. Ratliff,; C. Nenadic. Journal of the Royal Statistical Society. Series D (The Statistician), 43(1):214, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2348960.

Morris:1992:BRBa

[451] Pamela Morris. Book review: Statistics: Principles and Methods, by R. A. Johnson; G. K. Bhattacharyya. Journal of the Royal Statistical Society. Series D (The Statistician), 43(1):214, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348959.

Kimber:1992:BRB

[452] Alan Kimber. Book review: Simulation: A Statistical Perspective, by J. Kleijnen; W. Van Groenenwald. Journal of the Royal Statistical Society. Series D (The Statistician), 43(1):215, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348961.

Pewsey:1992:BRBb

[453] Arthur Pewsey. Book review: Exploring the Limits of Bootstrap, by R. LePage; L. Billard. Journal of the Royal Statistical Society. Series D (The Statistician), 43(1):215–216, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348962.

Lendrem:1992:BRB

[454] Dennis Lendrem. Book review: Analysis of Variance in Experimental Design, by H. R. Lindman. Journal of the Royal Statistical Society. Series D (The Statistician), 43(1):216, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348963.

Howel:1992:BRBb

[455] Denise Howel. Book review: Computer-Interactive Data Analysis, by A. D. Lunn; D. R. McNeil. Journal of the Royal Statistical Society. Series D (The Statistician), 43(1):216–217, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348964.

Gillespie:1992:BRB

[456] E. S. Gillespie. Book review: A Course in Business Statistics, by W. Mendenhall; R. J. Beaver. Journal of the Royal Statistical Society. Series D (The Statistician), 43(1):217– 218, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348965.

Glendinning:1992:BRB

[457] R. H. Glendinning. Book review: Multivariate Density Estimation, Theory, Practice and Visualization, by D. W. Scott. Journal of the Royal Statistical Society. Series D (The Statistician), 43(1):218–219, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348966.

Brooks:1992:BRBc

[458] Stephen Brooks. Book review: Applied Multivariate Statistics for the Social Sciences, by J. Stevens. Journal of the Royal Statistical Society. Series D (The Statistician), 43(1):219– 220, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348967.

Kemp:1992:BRBb

[459] A. W. Kemp. Book review: Kendall’s Advanced Theory of Statistics, by A. Stuart; J. K. Ord. Journal of the Royal Statistical Society. Series D (The Statistician), 43(1):220–221, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348968.

Morris:1992:BRBb

[460] Pamela Morris. Book review: Sampling, by S. K. Thompson. Journal of the Royal Statistical Society. Series D (The Statistician), 43(1):221, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348969.

Ghosh:1992:CRE

[461] Subir Ghosh and Edna Derderian. Correction: Robust experimental plan and its role in determining robust design against noise factors. Journal of the Royal Statistical Society. Series D (The Statistician), 43(1):222, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348970.

Anonymous:1992:BMh

[462] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 43(1):??, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348971.

Anonymous:1992:FMh

[463] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 43(1):??, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348927.

Lipsitz:1994:ECT

[464] Stuart R. Lipsitz and L. P. Zhao. Estimation in contingency tables with given marginals. Journal of the Royal Statistical Society. Series D (The Statistician), 43(2):223–230, ???? 1994. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348339.

Wiper:1994:HBC

[465] M. P. Wiper, S. French, and R. Cooke. Hypothesis-based calibration scores. Journal of the Royal Statistical Society. Series D (The Statistician), 43 (2):231–236, ???? 1994. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348340.

Islam:1994:APL

[466] M. Ataharul Islam. An alternative to the product limit method: The multistate approach. Journal of the Royal Statistical Society. Series D (The Statistician), 43(2):237–245, ???? 1994. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348341.

McLaren:1994:GGF

[467] C. E. McLaren, J. M. Legler, and G. M. Brittenham. The generalized χ^{2} goodness-of-fit test. Journal of the Royal Statistical Society. Series D (The Statistician), 43(2):247–258, ???? 1994. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348342.

Singer:1994:CAE

[468] Julio M. Singer and Dalton F. Andrade. On the choice of appropriate error terms in profile analysis. Journal of the Royal Statistical Society. Series D (The Statistician), 43 (2):259–266, ???? 1994. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348343.

Sharkey:1994:CSD

[469] A. J. Sharkey, O. G. Roch, and R. D. Coker. A case-study on the development of a sampling and testing protocol for aflatoxin levels in edible nuts and oil-seeds. Journal of the Royal Statistical Society. Series D (The Statistician), 43(2):267–275, ???? 1994. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348344.

Congdon:1994:AML

[470] Peter Congdon. Analysing mortality in London: Life-tables with frailty. Journal of the Royal Statistical Society. Series D (The Statistician), 43 (2):277–308, ???? 1994. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348345.

Jackson:1994:IBS

[471] David A. Jackson. Index betting on sports. Journal of the Royal Statistical Society. Series D (The Statistician), 43(2):309–315, ???? 1994. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2348346.

Lo:1994:CBT

[472] Victor S. Y. Lo and John Bacon-Shone. A comparison between two models for predicting ordering probabilities in multiple-entry competitions. Journal of the Royal Statistical Society. Series D (The Statistician), 43 (2):317–327, ???? 1994. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2348347.

Nason:1994:BRB

[473] Guy Nason. Book review: A First Course in Order Statistics, by B. C. Arnold; N. Balakrishnan; H. N. Nagaraja. Journal of the Royal Statistical Society. Series D (The Statistician), 43(2):329, ???? 1994. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2348348.

Woodward:1994:BRBa

[474] P. W. Woodward. Book review: Regression Analysis by Example, by S. Chatterjee; B. Price. Journal of the Royal Statistical Society. Series D (The Statistician), 43(2):329–330, ???? 1994. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348349.

Swires-Hennessy:1994:BRB

[475] E. Swires-Hennessy. Book review: Improving Information for Social Policy Decisions: The Uses of Microsimulation Modelling, Vol. I, Review and Recommendations; Improving Information for Social Policy Decisions: The Uses of Microsimulation Modelling Vol. II, Technical Papers, by C. F. Citro; E. A. Hanushek. Journal of the Royal Statistical Society. Series D (The Statistician), 43(2):330–331, ???? 1994. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348350.

Woodward:1994:BRBb

[476] P. W. Woodward. Book review: A Basic Course in Statistics, by G. M. Clarke; D. Cooke. Journal of the Royal Statistical Society. Series D (The Statistician), 43(2):331–332, ???? 1994. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348351.

Gudgeon:1994:BRBa

[477] Andrew C. Gudgeon. Book review: A First Course in Factor Analysis, by A. L. Comrey; H. B. Lee. Journal of the Royal Statistical Society. Series D (The Statistician), 43(2):332– 333, ???? 1994. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348352.

Lawson:1994:BRB

[478] Andrew Lawson. Book review: Complex Stochastic Systems, by D. R. Cox; D. M. Titterington. Journal of the Royal Statistical Society. Series D (The Statistician), 43(2):333, ???? 1994. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348353.

Hart:1994:BRB

[479] Anna Hart. Book review: Artificial Intelligence Frontiers in Statistics: AI and Statistics III, by D. J. Hand. Journal of the Royal Statistical Society. Series D (The Statistician), 43 (2):333–334, ???? 1994. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2348354.

Nuamah:1994:BRB

[480] Nicholas N. N. N. Nuamah. Book review: Modelling Seasonality, by S. Hylleberg. Journal of the Royal Statistical Society. Series D (The Statistician), 43(2):334–335, ???? 1994. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2348355.

Wates:1994:BRBa

[481] R. T. Wates. Book review: Statistics and Decisions: An Introduction to Foundations, by S. H. Kim. Journal of the Royal Statistical Society. Series D (The Statistician), 43(2):335– 336, ???? 1994. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348356.

Szuster:1994:BRB

[482] Fearnley Szuster. Book review: Nonsampling Error in Surveys, by J. T. Lessler; W. D. Kalsbeek. Journal of the Royal Statistical Society. Series D (The Statistician), 43(2):336, ???? 1994. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2348357.

Baxter:1994:BRB

[483] L. A. Baxter. Book review: Lectures on the Coupling Method, by T. Lindvall. Journal of the Royal Statistical Society. Series D (The Statistician), 43(2):336–337, ???? 1994. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348358.

Bullard:1994:BRB

[484] Peter Bullard. Book review: Introductory Statistics, by P. S. Mann. Journal of the Royal Statistical Society. Series D (The Statistician), 43 (2):337–338, ???? 1994. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348359.

Madrigal:1994:BRB

[485] J. L. Madrigal. Book review: Stochastic Models in Queueing Theory, by J. Medhi. Journal of the Royal Statistical Society. Series D (The Statistician), 43(2):338, ???? 1994. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348360.

Briscoe:1994:BRBa

[486] Simon Briscoe. Book review: International Historical Statistics, Europe 1750–1988, by B. R. Mitchell. Journal of the Royal Statistical Society. Series D (The Statistician), 43(2):339, ???? 1994. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348361.

Barbur:1994:BRBa

[487] Vicki A. Barbur. Book review: Introduction to Linear Regression Analysis, by D. C. Montgomery; E. A. Peck. Journal of the Royal Statistical Society. Series D (The Statistician), 43(2):339–341, ???? 1994. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348362.

Rao:1994:BRB

[488] T. Subba Rao. Book review: Identifiability in Stochastic Models: Characterization of Probability Distributions, by B. L. S. Prakasa Rao. Journal of the Royal Statistical Society. Series D (The Statistician), 43(2):341– 342, ???? 1994. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348363.

Cannings:1994:BRB

[489] Chris Cannings. Book review: Stochastic Games and Related Topics, by T. E. S. Raghavan; T. S. Ferguson; T. Parthasarathy; O. J. Vrieze. Journal of the Royal Statistical Society. Series D (The Statistician), 43(2):342, ???? 1994. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348364.

Coshall:1994:BRB

[490] J. Coshall. Book review: Introducing Data Analysis for Social Scientists, by D. Rose; O. Sullivan. Journal of the Royal Statistical Society. Series D (The Statistician), 43(2):342– 344, ???? 1994. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348365.

Spencer:1994:BRB

[491] Neil Spencer. Book review: Variance Components, by S. R. Searle; G. Casella; C. E. McCulloch. Journal of the Royal Statistical Society. Series D (The Statistician), 43(2):344– 345, ???? 1994. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348366.

Barbur:1994:BRBb

[492] Vicki A. Barbur. Book review: Statistical Process Control for Quality Improvement, by J. R. Thompson; J. Koronacki. Journal of the Royal Statistical Society. Series D (The Statistician), 43(2):345–346, ???? 1994. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348367.

Feinstein:1994:BRBa

[493] Alvan R. Feinstein. Book review: Central Tendency and Variability, by H. F. Weisberg. Journal of the Royal Statistical Society. Series D (The Statistician), 43(2):346–347, ???? 1994. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348368.

Thomas:1994:BRB

[494] G. E. Thomas. Book review: Resampling-Based Multiple Testing: Examples and Methods for p-Value Adjustment, by P. H. Westfall; S. S. Young. Journal of the Royal Statistical Society. Series D (The Statistician), 43(2):347–348, ???? 1994. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2348369.

Loynes:1994:BRB

[495] R. M. Loynes. Book review: Event History Analysis, by K. Yamaguchi. Journal of the Royal Statistical Society. Series D (The Statistician), 43(2): 348, ???? 1994. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348370.

Anonymous:1994:BMa

[496] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 43(2):??, ???? 1994. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2348371.

Anonymous:1994:FMa

[497] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 43(2):??, ???? 1994. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2348338.

Stickley:1994:NQA

[498] Adrian Stickley and Alan Winterbottom. The nature of quality assurance and statistical methods in BS 5750. Journal of the Royal Statistical Society. Series D (The Statistician), 43(3):349–370, ???? 1994. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2348571.

Coolen:1994:BEL

[499] F. P. A. Coolen. Bounds for expected loss in Bayesian decision theory with imprecise prior probabilities. Journal of the Royal Statistical Society. Series D (The Statistician), 43 (3):371–379, ???? 1994. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348572.

Kmietowicz:1994:VSL

[500] Z. W. Kmietowicz. The variance of the sample lead: A simple derivation. Journal of the Royal Statistical Society. Series D (The Statistician), 43 (3):381–388, ???? 1994. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348573.

Gillett:1994:APC

[501] Raphael Gillett. An average power criterion for sample size estimation. Journal of the Royal Statistical Society. Series D (The Statistician), 43 (3):389–394, ???? 1994. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348574.

Angus:1994:BOS

[502] John E. Angus. Bootstrap one-sided confidence intervals for the log-normal mean. Journal of the Royal Statistical Society. Series D (The Statistician), 43(3):395–401, ???? 1994. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348575.

Saporu:1994:UMC

[503] F. W. O. Saporu. The use of a Markov chain model in the analysis of data on the change in fertility status in cows. Journal of the Royal Statistical Society. Series D (The Statistician), 43(3):403–412, ???? 1994. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348576.

Federer:1994:POD

[504] Walter T. Federer. Pooling and other designs for analysing laboratory samples more efficiently. Journal of the Royal Statistical Society. Series D (The Statistician), 43(3):413–422, ???? 1994. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348577.

Moussa:1994:DBG

[505] Mohamed A. A. Moussa. Discrete boundaries in group sequential clinical trials. Journal of the Royal Statistical Society. Series D (The Statistician), 43(3):423–429, ???? 1994. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348578.

Klotz:1994:EMH

[506] Jerome H. Klotz and Linda D. Sharpless. Estimation for a Markov heart transplant model. Journal of the Royal Statistical Society. Series D (The Statistician), 43(3):431–438, ???? 1994. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348579.

Lo:1994:EMI

[507] Victor S. Y. Lo. On estimation methods of input parameters for economic attribute control charts. Journal of the Royal Statistical Society. Series D (The Statistician), 43(3):439–451, ???? 1994. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348580.

Porter:1994:BRBa

[508] M. A. Porter. Book review: Multiple Regression: Testing and Interpreting Interactions, by L. S. Aike; S. G. West. Journal of the Royal Statistical Society. Series D (The Statistician), 43(3):453, ???? 1994. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348581.

Challenor:1994:BRB

[509] Peter Challenor. Book review: Inverse Methods in Physical Oceanography, by A. F. Bennett. Journal of the Royal Statistical Society. Series D (The Statistician), 43(3):453– 454, ???? 1994. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348582.

Porter:1994:BRBb

[510] M. A. Porter. Book review: Classification, Data Analysis and Knowledge Organisation: Models and Methods with Applications, by H. H. Bock; P. Ihms. Journal of the Royal Statistical Society. Series D (The Statistician), 43(3):454, ???? 1994. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2348583.

Kumar:1994:BRBa

[511] Kuldeep Kumar. Book review: Approximation Theory VII, by E. W. Cheney; C. K. Chui; L. L. Schumaker. Journal of the Royal Statistical Society. Series D (The Statistician), 43(3):455, ???? 1994. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2348584.

Clarke:1994:BRB

[512] G. M. Clarke. Book review: Statistical Analysis of Reliability Data, by M. J. Crowder; A. C. Kimber; R. L. Smith; T. J. Sweeting. Journal of the Royal Statistical Society. Series D (The Statistician), 43(3):455– 456, ???? 1994. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348585.

Smeeton:1994:BRB

[513] Nigel Smeeton. Book review: Statistique Théorique et Appliquée, Vol. 1, Les Bases Théoriques, by P. Dagnelie. Journal of the Royal Statistical Society. Series D (The Statistician), 43 (3):456–457, ???? 1994. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2348586.

Biggins:1994:BRB

[514] J. D. Biggins. Book review: Large Deviation Techniques, and Applications, by A. Dembo; O. Zeitouni. Journal of the Royal Statistical Society. Series D (The Statistician), 43(3):457, ???? 1994. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348587.

Saldanha:1994:BRB

[515] Richard A. Saldanha. Book review: Probabilistic Causality, by E. Eells. Journal of the Royal Statistical Society. Series D (The Statistician), 43 (3):457–458, ???? 1994. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348588.

Calder:1994:BRB

[516] A. B. Calder. Book review: Talking About Statistics, by B. Everitt; D. Hay. Journal of the Royal Statistical Society. Series D (The Statistician), 43(3):458–459, ???? 1994. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348589.

Gudgeon:1994:BRBb

[517] Andrew C. Gudgeon. Book review: Differential Item Functioning, by P. W. Holland; H. Wainer. Journal of the Royal Statistical Society. Series D (The Statistician), 43(3):459–460, ???? 1994. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348590.

Giles:1994:BRBa

[518] Ronald Giles. Book review: Longitudinal Data with Serial Correlation: A Space-State Approach, by R. H. Jones. Journal of the Royal Statistical Society. Series D (The Statistician), 43 (3):460–461, ???? 1994. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2348591.

Lounds:1994:BRB

[519] W. S. Lounds. Book review: Time Series, by M. Kendall; J. K. Ord. Journal of the Royal Statistical Society. Series D (The Statistician), 43(3):461, ???? 1994. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348592.

Heath:1994:BRB

[520] David Heath. Book review: Reviews of United Kingdom Statistical Sources, Vol. XXVIII, The Food Industries, by J. Mark; R. Strange. Journal of the Royal Statistical Society. Series D (The Statistician), 43 (3):461–462, ???? 1994. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2348593.

Giles:1994:BRBb

[521] Ronald Giles. Book review: Discrete Dynamic Systems and Chaos, by M. Martelli. Journal of the Royal Statistical Society. Series D (The Statistician), 43(3):462, ???? 1994. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2348594.

Langford:1994:BRB

[522] I. Langford. Book review: Social Research: Issues, Methods and Process, by T. May. Journal of the Royal Statistical Society. Series D (The Statistician), 43(3):463, ???? 1994. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348595.

Crowder:1994:BRB

[523] Martin Crowder. Book review: Elements of Modern Asymptotic Theory with Statistical Applications, by B. McCabe; A. Tremayne. Journal of the Royal Statistical Society. Series D (The Statistician), 43(3):463– 464, ???? 1994. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348596.

Kumar:1994:BRBb

[524] Vijay Kumar. Book review: Statistical Methods in Agriculture and Experimental Biology, by R. Mead; R. N. Curnow; A. M. Hasted. Journal of the Royal Statistical Society. Series D (The Statistician), 43(3):464–465, ???? 1994. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348597.

Simakani:1994:BRB

[525] June E. Simakani. Book review: Introduction to Statistics: The Non-Parametric Way, by G. E. Noether. Journal of the Royal Statistical Society. Series D (The Statistician), 43(3): 465, ???? 1994. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348598.

Butcher:1994:BRB

[526] Bob Butcher. Book review: Introduction to Social Research, by V. Pons. Journal of the Royal Statistical Society. Series D (The Statistician), 43(3):466, ???? 1994. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2348599.

Metcalfe:1994:BRBa

[527] Andrew Metcalfe. Book review: Meta-Analytic Procedures for Social Research, by R. Rosenthal. Journal of the Royal Statistical Society. Series D (The Statistician), 43(3):466– 467, ???? 1994. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348600.

Csenki:1994:BRBa

[528] A. Csenki. Book review: Introduction to Probability Models, by S. M. Ross. Journal of the Royal Statistical Society. Series D (The Statistician), 43(3):467–468, ???? 1994. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2348601.

Konijn:1994:BRB

[529] H. S. Konijn. Book review: Categorical Data Analysis, by AIC, by Y. Sakamoto. Journal of the Royal Statistical Society. Series D (The Statistician), 43(3):468–469, ???? 1994. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2348602.

Grieve:1994:BRB

[530] A. P. Grieve. Book review: Cross-Over Trials in Clinical Research, by S. Senn. Journal of the Royal Statistical Society. Series D (The Statistician), 43(3):469–470, ???? 1994. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348603.

Wates:1994:BRBb

[531] R. T. Wates. Book review: Markovian Queues, by O. P. Sharma. Journal of the Royal Statistical Society. Series D (The Statistician), 43(3):470– 471, ???? 1994. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348604.

Lynn:1994:BRB

[532] Peter Lynn. Book review: Directory of Social Research Organisations in the United Kingdom, by W. Sykes; M. Bulmer; M. Schwerzel. Journal of the Royal Statistical Society. Series D (The Statistician), 43(3):471, ???? 1994. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348605.

Csenki:1994:BRBb

[533] A. Csenki. Book review: An Introduction to Stochastic Processes and Their Applications, by P. Todorovic. Journal of the Royal Statistical Society. Series D (The Statistician), 43(3):472, ???? 1994. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348606.

Briscoe:1994:BRBb

[534] Simon Briscoe. Book review: Dictionary of Statistics and Methodology, by W. P. Vogt. Journal of the Royal Statistical Society. Series D (The Statistician), 43(3):473, ???? 1994. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348607.

Anonymous:1994:BMb

[535] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 43(3):474, ???? 1994. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348608.

Anonymous:1994:FMb

[536] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 43(3):??, ???? 1994. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2348570.

Kianifard:1994:ECS

[537] Farid Kianifard. Evaluation of clinimetric scales: Basic principles and methods. Journal of the Royal Statistical Society. Series D (The Statistician), 43(4):475–482, ???? 1994. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2348132.

Brooks:1994:DPC

[538] S. P. Brooks. Diagnostics for principal components: Influence functions as diagnostic tools. Journal of the Royal Statistical Society. Series D (The Statistician), 43(4):483– 494, ???? 1994. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348133.

Baker:1994:MPT

[539] Stuart G. Baker. The multinomial-Poisson transformation. Journal of the Royal Statistical Society. Series D (The Statistician), 43(4):495–504, ???? 1994. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348134.

Delap:1994:IAW

[540] Martin R. Delap. The interpretation of achieved weights of examination components. Journal of the Royal Statistical Society. Series D (The Statistician), 43(4):505–511, ???? 1994. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348135.

Congdon:1994:SAA

[541] Peter Congdon. Spatiotemporal analysis of area mortality. Journal of the Royal Statistical Society. Series D (The Statistician), 43(4):513–528, ???? 1994. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348136.

Taylor:1994:AEU

[542] G. A. R. Taylor. Analysis of experiments by using half-normal plots. Journal of the Royal Statistical Society. Series D (The Statistician), 43 (4):529–536, ???? 1994. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348137.

Lane:1994:MEC

[543] Peter W. Lane and Steve P. McGrath. Modelling the effect of cultivation on the dispersion of pollutants in soil. Journal of the Royal Statistical Society. Series D (The Statistician), 43 (4):537–549, ???? 1994. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348138.

Silver:1994:MIU

[544] Mick Silver and Christos Ioannidis. The measurement of inflation; untimely weights and alternative formulae: European evidence. Journal of the Royal Statistical Society. Series D (The Statistician), 43(4):551–562, ???? 1994. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348139.

OQuigley:1994:TST

[545] John O’Quigley. On a two-sided test for crossing hazards. Journal of the Royal Statistical Society. Series D (The Statistician), 43(4):563– 569, ???? 1994. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348140.

Jonsson:1994:SBP

[546] Bo Jonsson. Sample-based proportions as values on a regressor. Journal of the Royal Statistical Society. Series D (The Statistician), 43(4):571– 583, ???? 1994. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348141.

Percy:1994:SS

[547] David F. Percy. Stochastic snooker. Journal of the Royal Statistical Society. Series D (The Statistician), 43 (4):585–594, ???? 1994. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348142.

Thomas:1994:C

[548] R. G. Thomas and Oliver D. Anderson. Correspondence. Journal of the Royal Statistical Society. Series D (The Statistician), 43(4):595–597, ???? 1994. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348143.

Shade:1994:BRB

[549] John Shade. Book review: Reliability and Decision Making, by R. E. Barlow; C. A. Clarotti; F. Spizzichino. Journal of the Royal Statistical Society. Series D (The Statistician), 43 (4):598–599, ???? 1994. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348144.

Smith:1994:BRB

[550] J. U. M. Smith. Book review: Reliability Evaluation of Engineering Systems, by R. Billington; R. N. Allan. Journal of the Royal Statistical Society. Series D (The Statistician), 43 (4):599–600, ???? 1994. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348145.

Kimber:1994:BRB

[551] Alan Kimber. Book review: An Introduction to the Bootstrap, by B. Efron; R. J. Tibshirani. Journal of the Royal Statistical Society. Series D (The Statistician), 43(4):600, ???? 1994. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348146.

Gudgeon:1994:BRBc

[552] Andrew C. Gudgeon. Book review: Test Theory for a New Generation of Tests, by N. Fredericksen; R. Mislevy; I. I. Bejar. Journal of the Royal Statistical Society. Series D (The Statistician), 43(4):601, ???? 1994. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348147.

Woodward:1994:BRBc

[553] P. W. Woodward. Book review: Teaching Statistical Concepts, by A. Hawkins; F. Jolliffe; L. Glickman. Journal of the Royal Statistical Society. Series D (The Statistician), 43(4): 602, ???? 1994. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348148.

Holmes:1994:BRB

[554] Andrew Holmes. Book review: Design and Inference in Finite Population Sampling, by A. S. Hedayat; B. K. Sinha. Journal of the Royal Statistical Society. Series D (The Statistician), 43(4):603, ???? 1994. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348149.

Gudgeon:1994:BRBd

[555] Andrew C. Gudgeon. Book review: Strategies and Tactics of Behavioral Research, by J. M. Johnston; H. S. Pennypacker; Readings for Strategies and Tactics of Behavioral Research, by J. M. Johnston; H. S. Pennypacker. Journal of the Royal Statistical Society. Series D (The Statistician), 43 (4):603–604, ???? 1994. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348150.

Feinstein:1994:BRBb

[556] Alvan R. Feinstein. Book review: 100 Statistical Tests, by G. K. Kanji. Journal of the Royal Statistical Society. Series D (The Statistician), 43(4):605, ???? 1994. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348151.

Currall:1994:BRB

[557] James Currall. Book review: SPSS/PC + Made Simple, by P. R. Kinnear; C. D. Gray. Journal of the Royal Statistical Society. Series D (The Statistician), 43(4):605–607, ???? 1994. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348152.

Powell:1994:BRB

[558] Suzanne J. Powell. Book review: Survival Analysis: State of the Art, by J. P. Klein; P. K. Goel. Journal of the Royal Statistical Society. Series D (The Statistician), 43(4):607, ???? 1994. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348153.

Saggu:1994:BRB

[559] P. S. Saggu. Book review: Statistical Methods for Survival Data Analysis, by E. T. Lee. Journal of the Royal Statistical Society. Series D (The Statistician), 43(4):607–608, ???? 1994. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348154.

Day:1994:BRB

[560] Simon Day. Book review: Measurement, Design and Analysis: An Integrated Approach, by E. J. Pedhazur; L. Pedhazur Schmelkin. Journal of the Royal Statistical Society. Series D (The Statistician), 43(4):608–609, ???? 1994. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348155.

Jolliffe:1994:BRB

[561] Ian Jolliffe. Book review: Applied Factor Analysis in the Natural Sciences, by R. A. Reyment; K. G. Jöreskog. Journal of the Royal Statistical Society. Series D (The Statistician), 43 (4):609–610, ???? 1994. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2348156.

Metcalfe:1994:BRBb

[562] Andrew V. Metcalfe. Book review: Philosophical and Foundational Issues in Measurement Theory, by C. W. Savage; P. Ehrlich. Journal of the Royal Statistical Society. Series D (The Statistician), 43(4):610–611, ???? 1994. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348157.

Roberts:1994:BRB

[563] Martin Roberts. Book review: Linear Programming: An Introduction with Applications, by A. Sultan. Journal of the Royal Statistical Society. Series D (The Statistician), 43(4):611– 612, ???? 1994. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348158.

Ibrahim:1994:CBO

[564] Kamarulzaman Bin Ibrahim and A. V. Metcalfe. Correction: Bayesian overview for evaluation of miniroundabouts as a road safety measure. Journal of the Royal Statistical Society. Series D (The Statistician), 43(4):612, ???? 1994. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348159. See

[411].

Anonymous:1994:VI

[565] Anonymous. Volume information. Journal of the Royal Statistical Society. Series D (The Statistician), 43 (4):613–616, ???? 1994. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348130.

Anonymous:1994:BMc

[566] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 43(4):??, ???? 1994. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348160.

Anonymous:1994:FMc

[567] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 43(4):??, ???? 1994. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348131.

Anonymous:1995:Fa

[568] Anonymous. Foreward. Journal of the Royal Statistical Society. Series D (The Statistician), 44(1):1, ???? 1995. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348610.

Openshaw:1995:DAS

[569] Stan Openshaw. Developing automated and Smart spatial pattern exploration tools for geographical information systems applications. Journal of the Royal Statistical Society. Series D (The Statistician), 44(1):3–16, ???? 1995. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348611.

Stack:1995:SEV

[570] Jim Stack and Antony R. Unwin. Selecting explanatory variables from survey data: A simple approach that explores for interactions. Journal of the Royal Statistical Society. Series D (The Statistician), 44(1):17–29, ???? 1995. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348612.

Dodd:1995:RCC

[571] F. J. Dodd, H. A. Donegan, and T. B. M. McMaster. Reassessment of consistency criteria in judgment matrices. Journal of the Royal Statistical Society. Series D (The Statistician), 44(1):31–41, ???? 1995. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348613.

Harrison:1995:CAS

[572] Michael J. Harrison. Conference on Applied Statistics in Ireland, 1993: Symposium on the Teaching of Statistics in Higher Education in Ireland: Introduction. Journal of the Royal Statistical Society. Series D (The Statistician), 44(1):43, ???? 1995. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348614.

Stuart:1995:CTS

[573] Michael Stuart. Changing the teaching of statistics. Journal of the Royal Statistical Society. Series D (The Statistician), 44(1):45–54, ???? 1995. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348615.

Anonymous:1995:DST

[574] Anonymous. Discussion on the symposium on the teaching of statistics in higher education in Ireland. Journal of the Royal Statistical Society. Series D (The Statistician), 44(1):55–58, ???? 1995. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348616.

Anonymous:1995:APP

[575] Anonymous. [abstracts of papers presented at the Conference on Applied Statistics in Ireland, 1993]. Journal of the Royal Statistical Society. Series D (The Statistician), 44(1):59–67, ???? 1995. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348617.

Woodward:1995:ARF

[576] Mark Woodward, Kirsty Laurent, and Hugh Tunstall-Pedoe. An analysis of risk factors for prevalent coronary heart disease by using the proportional odds model. Journal of the Royal Statistical Society. Series D (The Statistician), 44(1):69–80, ???? 1995. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2348618.

Lipsitz:1995:SSC

[577] Stuart R. Lipsitz and Michael Parzen. Sample size calculations for nonrandomized studies. Journal of the Royal Statistical Society. Series D (The Statistician), 44(1):81–90, ???? 1995. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2348619.

Wilcox:1995:CTI

[578] Rand R. Wilcox. Comparing two independent groups via multiple quantiles. Journal of the Royal Statistical Society. Series D (The Statistician), 44(1):91–99, ???? 1995. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2348620.

Silver:1995:SMC

[579] Mick Silver. Scales of measurement and cluster analysis: An application concerning market segments in the babyfood market. Journal of the Royal Statistical Society. Series D (The Statistician), 44(1):101–112, ???? 1995. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2348621.

Everitt:1995:ARM

[580] B. S. Everitt. The analysis of repeated measures: A practical review with examples. Journal of the Royal Statistical Society. Series D (The Statistician), 44(1):113–135, ???? 1995. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348622.

Adams:1995:BRB

[581] Robert Adams. Book review: Construction Versus Choice in Cognitive Measurement, by R. E. Bennett; W. C. Ward. Journal of the Royal Statistical Society. Series D (The Statistician), 44(1):137, ???? 1995. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348623.

Scott:1995:BRB

[582] E. M. Scott. Book review: Modelling Survival Data in Medical Research, by D. Collett. Journal of the Royal Statistical Society. Series D (The Statistician), 44(1):137–138, ???? 1995. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348624.

Swires-Hennessy:1995:BRB

[583] E. Swires-Hennessy. Book review: Statistics for the 21st Century: Proposals for Improving Statistics for Better Decision Making, by J. W. Duncan; A. C. Gross. Journal of the Royal Statistical Society. Series D (The Statistician), 44(1):138–139, ???? 1995. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348625.

Glendinning:1995:BRBa

[584] R. H. Glendinning. Book review: Asymptotic Problems in Probability Theory: Stochastic Models and Diffusions on Fractals, by K. D. Elworthy; N. Ikeda. Journal of the Royal Statistical Society. Series D (The Statistician), 44(1):139–140, ???? 1995. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2348626.

Penny:1995:BRB

[585] Richard N. Penny. Book review: Model Assisted Survey Sampling, by C.-E. Särndel; B. Swensson; J. Wretman. Journal of the Royal Statistical Society. Series D (The Statistician), 44 (1):140–142, ???? 1995. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2348627.

Sedgwick:1995:BRB

[586] Philip Sedgwick. Book review: Probability Theory: An Introductory Course, by Y. Sinai; D. Haughton. Journal of the Royal Statistical Society. Series D (The Statistician), 44(1):142, ???? 1995. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348628.

Anonymous:1995:BMa

[587] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 44(1):??, ???? 1995. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2348629.

Anonymous:1995:FMa

[588] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 44(1):??, ???? 1995. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348609.

Joseph:1995:SSC

[589] Lawrence Joseph, David B. Wolfson, and Roxane Du Berger. Sample size calculations for binomial proportions via highest posterior density intervals. Journal of the Royal Statistical Society. Series D (The Statistician), 44 (2):143–154, ???? 1995. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348439. See com-

ments [590].

Adcock:1995:BAD

[590] C. J. Adcock. The Bayesian approach to determination of sample sizes — some comments on the paper by Joseph, Wolfson and du Berger. Journal of the Royal Statistical Society. Series D (The Statistician), 44(2):155– 161, ???? 1995. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348440. See [589].

Pham-Gia:1995:SSD

[591] T. Pham-Gia. Sample size determination in Bayesian statistics — a commentary. Journal of the Royal Statistical Society. Series D (The Statistician), 44(2):163–166, ???? 1995. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348441.

Joseph:1995:SCB

[592] Lawrence Joseph, David B. Wolfson, and Roxane du Berger. Some comments on Bayesian sample size determination. Journal of the Royal Statistical Society. Series D (The Statistician), 44(2):167–171, ???? 1995. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2348442.

Lynn:1995:BVS

[593] Peter Lynn and Bridget Taylor. On the bias and variance of samples of individuals: A comparison of the electoral registers and postcode address file as sampling frames. Journal of the Royal Statistical Society. Series D (The Statistician), 44(2):173–194, ???? 1995. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2348443.

Olszak:1995:BNO

[594] Michael Olszak and Gilbert Ritschard. The behaviour of nominal and ordinal partial association measures. Journal of the Royal Statistical Society. Series D (The Statistician), 44(2):195– 212, ???? 1995. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348444.

Wilcox:1995:SRS

[595] Rand R. Wilcox. Simulation results on solutions to the multivariate Behrens– Fisher problem via trimmed means. Journal of the Royal Statistical Society. Series D (The Statistician), 44 (2):213–225, ???? 1995. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348445.

Stone:1995:TSM

[596] Richard Stone and Mark Taylor. Time series models in statistical process control: Considerations of applicability. Journal of the Royal Statistical Society. Series D (The Statistician), 44 (2):227–234, ???? 1995. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348446.

Theobald:1995:DUC

[597] C. M. Theobald and S. P. J. Kirby. Discrimination using covariance selection models in the automated allocation of human chromosomes. Journal of the Royal Statistical Society. Series D (The Statistician), 44(2):235– 240, ???? 1995. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348447.

Brooks:1995:OUS

[598] S. P. Brooks and B. J. T. Morgan. Optimization using simulated annealing. Journal of the Royal Statistical Society. Series D (The Statistician), 44(2):241–257, ???? 1995. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348448.

Dobson:1995:DPL

[599] S. M. Dobson and J. A. Goddard. The demand for Professional League Football in England and Wales, 1925–92. Journal of the Royal Statistical Society. Series D (The Statistician), 44 (2):259–277, ???? 1995. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2348449.

Corcoran:1995:BRB

[600] Stephen Corcoran. Book review: Inference and Asymptotics, by O. E. Barndorff-Nielsen; D. R. Cox. Journal of the Royal Statistical Society. Series D (The Statistician), 44(2):279, ???? 1995. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348450.

Sheil:1995:BRB

[601] J. Sheil. Book review: Statistical Methods for SPC and TQM, by D. Bissell. Journal of the Royal Statistical Society. Series D (The Statistician), 44(2):279–281, ???? 1995. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2348451.

Watkins:1995:BRB

[602] James F. Watkins. Book review: Modelling Survival Data in Medical Research, by D. Collett. Journal of the Royal Statistical Society. Series D (The Statistician), 44(2):281–282, ???? 1995. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348452.

Vajda:1995:BRB

[603] S. Vajda. Book review: Practical Risk Theory for Actuaries, by C. D. Daykin; T. Pentikainen; M. Pesonen. Journal of the Royal Statistical Society. Series D (The Statistician), 44 (2):282–283, ???? 1995. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348453.

Neuberg:1995:BRB

[604] Leland G. Neuberg. Book review: Bayes or Bust? — A Critical Examination of Bayesian Confirmation Theory, by J. Earman. Journal of the Royal Statistical Society. Series D (The Statistician), 44(2):283–284, ???? 1995. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348454.

Woodward:1995:BRB

[605] P. W. Woodward. Book review: Understanding Probability and Statistics: A Book of Problems, by R. Falk. Journal of the Royal Statistical Society. Series D (The Statistician), 44 (2):284–285, ???? 1995. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348455.

Kumar:1995:BRB

[606] Kuldeep Kumar. Book review: Statistical Analysis of Circular Data, by N. I. Fisher. Journal of the Royal Statistical Society. Series D (The Statistician), 44(2):285–286, ???? 1995. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348456.

Challenor:1995:BRB

[607] Peter Challenor. Book review: Statistical Analysis of Spherical Data, by N. I. Fisher; T. Lewis; B. J. J. Embleton. Journal of the Royal Statistical Society. Series D (The Statistician), 44(2):286–287, ???? 1995. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2348457.

Bello:1995:BRB

[608] Abdul L. Bello. Book review: Tables of Integrals, Series, and Products, by I. S. Gradshteyn; I. M. Ryzhik; A. Jeffrey. Journal of the Royal Statistical Society. Series D (The Statistician), 44(2):287–288, ???? 1995. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2348458.

Saldanha:1995:BRB

[609] Richard A. Saldanha. Book review: Computer Intensive Statistical Methods: Validation, Model Selection and Bootstrap, by J. S. U. Hjorth. Journal of the Royal Statistical Society. Series D (The Statistician), 44(2):288, ???? 1995. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348459.

Fielding:1995:BRBa

[610] A. Fielding. Book review: The Data Sets from CBDEA, by T. P. Hutchison; C. D. Lai. Journal of the Royal Statistical Society. Series D (The Statistician), 44(2):288–289, ???? 1995. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2348460.

Simakani:1995:BRB

[611] J. E. Simakani. Book review: Asymptotic Problems in Probability Theory: Wiener Functionals and Fractionals, by N. Ikeda; K. D. Elworthy. Journal of the Royal Statistical Society. Series D (The Statistician), 44(2):289, ???? 1995. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348461.

Fielding:1995:BRBb

[612] A. Fielding. Book review: Non-Sampling Errors in Surveys, by J. T. Kessler; W. D. Kelsbeck. Journal of the Royal Statistical Society. Series D (The Statistician), 44(2):289– 290, ???? 1995. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348462.

Shanmugam:1995:BRB

[613] Ramalingam Shanmugam. Book review: Econometrics of Information and Efficiency, by J. K. Sengupta. Journal of the Royal Statistical Society. Series D (The Statistician), 44 (2):290–291, ???? 1995. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348463.

Crowder:1995:BRB

[614] Martin Crowder. Book review: An Introduction to Stochastic Modelling, by H. M. Taylor; S. Karlin. Journal of the Royal Statistical Society. Series D (The Statistician), 44(2):291– 292, ???? 1995. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348464.

Brooks:1995:BRB

[615] Stephen Brooks. Book review: Markov Decision Processes, by D. J. White.

Journal of the Royal Statistical Society. Series D (The Statistician), 44(2): 292, ???? 1995. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348465.

Anonymous:1995:BMb

[616] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 44(2):??, ???? 1995. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2348466.

Anonymous:1995:FMb

[617] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 44(2):??, ???? 1995. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2348438.

Anonymous:1995:Fb

[618] Anonymous. Foreword. Journal of the Royal Statistical Society. Series D (The Statistician), 44(3):293, ???? 1995. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2348700.

Kmietowicz:1995:AII

[619] Z. W. Kmietowicz. Accuracy of indices of industrial production in developing countries. Journal of the Royal Statistical Society. Series D (The Statistician), 44(3):295–307, ???? 1995. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2348701.

Silver:1995:EIV

[620] Mick Silver and Christos Ioannidis. European inflation and its variability. Journal of the Royal Statistical Society. Series D (The Statistician), 44 (3):309–322, ???? 1995. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348702.

Mills:1995:MSK

[621] Terence C. Mills. Modelling skewness and kurtosis in the London Stock Exchange FT-SE Index return distributions. Journal of the Royal Statistical Society. Series D (The Statistician), 44(3):323–332, ???? 1995. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348703.

Li:1995:MAS

[622] W. K. Li and K. Lam. Modelling asymmetry in stock returns by a threshold autoregressive conditional heteroscedastic model. Journal of the Royal Statistical Society. Series D (The Statistician), 44(3):333–341, ???? 1995. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348704.

Amirkhalkhali:1995:FPE

[623] Saleh Amirkhalkhali, Tomson Ogwang, Samad Amirkhalkhali, and U. L. Gouranga Rao. On the forecasting performance of estimators for a structural equation in a large system: Some Monte Carlo results. Journal of the Royal Statistical Society. Series D (The Statistician), 44(3):343– 351, ???? 1995. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348705.

Watier:1995:MET

[624] Laurence Watier and Sylvia Richardson. Modelling of an epidemiological time series by a threshold autoregressive model. Journal of the Royal Statistical Society. Series D (The Statistician), 44(3):353–364, ???? 1995. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2348706.

Mackenzie:1995:RMR

[625] Gilbert Mackenzie. Regression models for repeated medical random counts. Journal of the Royal Statistical Society. Series D (The Statistician), 44 (3):365–372, ???? 1995. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2348707.

Collins:1995:BCL

[626] Alan J. Collins. Bootstrap confidence limits on process capability indices. Journal of the Royal Statistical Society. Series D (The Statistician), 44 (3):373–378, ???? 1995. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2348708.

Wang:1995:DTL

[627] P. C. Wang and H. W. Jan. Designing two-level factorial experiments using orthogonal arrays when the run order is important. Journal of the Royal Statistical Society. Series D (The Statistician), 44(3):379–388, ???? 1995. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348709.

Mukhopadhyay:1995:ELB

[628] Saurabh Mukhopadhyay and Brani Vidakovic. Efficiency of linear Bayes rules for a normal mean: Skewed priors class. Journal of the Royal Statistical Society. Series D (The Statistician), 44(3):389–397, ???? 1995. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348710.

Draper:1995:CIV

[629] Norman R. Draper and Irwin Guttman. Confidence intervals versus regions. Journal of the Royal Statistical Society. Series D (The Statistician), 44 (3):399–403, ???? 1995. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348711.

Caudill:1995:C

[630] Steven B. Caudill. Correspondence. Journal of the Royal Statistical Society. Series D (The Statistician), 44 (3):405–406, ???? 1995. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348712.

Hingley:1995:BRB

[631] Peter J. Hingley. Book review: Inference and Asymptotics, by O. E. Barndorff-Nielsen; D. R. Cox. Journal of the Royal Statistical Society. Series D (The Statistician), 44(3):407– 408, ???? 1995. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348713.

Kimber:1995:BRB

[632] Alan Kimber. Book review: A Handbook of Statistical Analyses Using S-Plus, by B. S. Everitt. Journal of the Royal Statistical Society. Series D (The Statistician), 44(3):408–409, ???? 1995. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348714.

Gerontidis:1995:BRB

[633] Ioannis I. Gerontidis. Book review: Mathematical Methods in Queuing Theory, by V. V. Kalashnikov. Journal of the Royal Statistical Society. Series D (The Statistician), 44 (3):409–410, ???? 1995. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2348715.

Howel:1995:BRB

[634] Denise Howel. Book review: Logistic Regression: A Self Learning Text, by D. G. Kleinbaum. Journal of the Royal Statistical Society. Series D (The Statistician), 44(3):410–411, ???? 1995. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348716.

Glendinning:1995:BRBb

[635] R. H. Glendinning. Book review: Theory of U-Statistics, by V. S. Koroljuk; Yu. V. Borovskich. Journal of the Royal Statistical Society. Series D (The Statistician), 44(3):411, 1995. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2348717.

Bullard:1995:BRB

[636] Peter R. Bullard. Book review: Understanding Probability and Statistics — A Book of Problems, by A. K. Peters. Journal of the Royal Statistical Society. Series D (The Statistician), 44(3):412, ???? 1995. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348718.

Phillips:1995:BRB

[637] Michael J. Phillips. Book review: The Reliability, Availability and Productiveness of Systems, by D. J. Sherwin; A. Bossche. Journal of the Royal Statistical Society. Series D (The Statistician), 44(3):412–413, ???? 1995. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348719.

Csenki:1995:BRB

[638] A. Csenki. Book review: Applied Methods of Structural Reliability, by M. Tichy. Journal of the Royal Statistical Society. Series D (The Statistician), 44(3):414–415, ???? 1995. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348720.

Anonymous:1995:BMc

[639] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 44(3):??, ???? 1995. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348721.

Anonymous:1995:FMc

[640] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 44(3):??, ???? 1995. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348699.

Berry:1995:MCI

[641] G. Berry and P. Armitage. Mid-P confidence intervals: A brief review. Journal of the Royal Statistical Society. Series D (The Statistician), 44 (4):417–423, ???? 1995. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348891.

Tomizawa:1995:MDM

[642] Sadao Tomizawa. Measures of departure from marginal homogeneity for contingency tables with nominal categories. Journal of the Royal Statistical Society. Series D (The Statistician), 44(4):425–439, ???? 1995. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348892.

Andres:1995:WSF

[643] A. Martín Andrés, I. Herranz Tejedor, and A. Silva Mato. The Wilcoxon, Spearman, Fisher, χ^{2}-, Student and Pearson tests and 2 × 2 tables. Journal of the Royal Statistical Society. Series D (The Statistician), 44(4):441– 450, ???? 1995. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348893.

Byar:1995:IWS

[644] David P. Byar, Laurence S. Freedman, and Agnes M. Herzberg. Identifying which sets of parameters are simultaneously estimable in an incomplete factorial design. Journal of the Royal Statistical Society. Series D (The Statistician), 44(4):451–456, ???? 1995. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348894.

Kmietowicz:1995:EDS

[645] Z. W. Kmietowicz. The exact distribution of the sample lead: Further results. Journal of the Royal Statistical Society. Series D (The Statistician), 44(4):457–467, ???? 1995. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348895.

Robinson:1995:SWS

[646] Stephen Robinson and Sally McClean. Simulation of wind speed for the modelling of pine saw-fly mortality. Journal of the Royal Statistical Society. Series D (The Statistician), 44(4):469– 477, ???? 1995. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348896.

Box:1995:DPI

[647] George E. P. Box and Alberto Luceño. Discrete proportional-integral control with constrained adjustment. Journal of the Royal Statistical Society. Series D (The Statistician), 44(4):479– 495, ???? 1995. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348897.

Graf:1995:IME

[648] Erika Graf and Martin Schumacher. An investigation on measures of explained variation in survival analysis.

Journal of the Royal Statistical Society. Series D (The Statistician), 44 (4):497–507, ???? 1995. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2348898.

Clarke:1995:HGA

[649] Stephen R. Clarke and John M. Norman. Home ground advantage of individual clubs in English soccer. Journal of the Royal Statistical Society. Series D (The Statistician), 44(4):509– 521, ???? 1995. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348899.

Kuk:1995:MPC

[650] Anthony Y. C. Kuk. Modelling paired comparison data with large numbers of draws and large variability of draw percentages among players. Journal of the Royal Statistical Society. Series D (The Statistician), 44(4):523– 528, ???? 1995. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348900.

Appleton:1995:MBM

[651] David R. Appleton. May the best man win? Journal of the Royal Statistical Society. Series D (The Statistician), 44(4):529–538, ???? 1995. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2348901.

Goda:1995:BRB

[652] David Goda. Book review: Understanding Social Science Statistics: A Spreadsheet Approach, by R. Bakeman.

Journal of the Royal Statistical Society. Series D (The Statistician), 44 (4):539–540, ???? 1995. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348902.

Smith:1995:BRB

[653] Murray D. Smith. Book review: Computational Techniques for Econometrics and Economic Analysis, by D. A. Belsley. Journal of the Royal Statistical Society. Series D (The Statistician), 44(4):540, ???? 1995. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348903.

Shade:1995:BRB

[654] John Shade. Book review: Using Statistics in Industry, by G. Betteley; N. Mettrick; E. Sweeney; D. Wilson. Journal of the Royal Statistical Society. Series D (The Statistician), 44(4):541–542, ???? 1995. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348904.

Burrell:1995:BRB

[655] Quentin Burrell. Book review: Probability with a View Toward Statistics, by J. Hoffmann-Jörgensen. Journal of the Royal Statistical Society. Series D (The Statistician), 44(4):542– 543, ???? 1995. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348905.

Sherwin:1995:BRB

[656] David Sherwin. Book review: System Reliability Theory-Models and Statistical Methods, by A. Høyland; M. Rausand. Journal of the Royal Statistical Society. Series D (The Statistician), 44(4):543–544, ???? 1995. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2348906.

Chatfield:1995:BRB

[657] Chris Chatfield. Book review: Continuous Univariate Distributions, by N. L. Johnson; S. Kotz; N. Balakrishnan. Journal of the Royal Statistical Society. Series D (The Statistician), 44(4):544–545, ???? 1995. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2348907.

Day:1995:BRB

[658] Simon Day. Book review: Analysis of Two-Way Layouts, by J. Mandel. Journal of the Royal Statistical Society. Series D (The Statistician), 44(4):545–546, ???? 1995. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2348908.

Fielding:1995:BRBc

[659] Antony Fielding. Book review: Data Analysis: An Introduction, by B. Nolan. Journal of the Royal Statistical Society. Series D (The Statistician), 44(4):546–547, ???? 1995. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2348909.

Ipinyomi:1995:CNR

[660] R. A. Ipinyomi and B. L. Adeleke. Correction: Nested row-column designs. Journal of the Royal Statistical Society. Series D (The Statistician), 44(4):547, ???? 1995. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348910.

Anonymous:1995:VI

[661] Anonymous. Volume information. Journal of the Royal Statistical Society. Series D (The Statistician), 44 (4):548–550, ???? 1995. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348889.

Anonymous:1995:BMd

[662] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 44(4):??, ???? 1995. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348911.

Anonymous:1995:FMd

[663] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 44(4):??, ???? 1995. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348890.

Boland:1996:F

[664] Philip J. Boland and Yudi Pawitan. Foreword. Journal of the Royal Statistical Society. Series D (The Statistician), 45(1):1, ???? 1996. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348406.

Barry:1996:EBA

[665] Daniel Barry. An empirical Bayes approach to growth curve analysis. Journal of the Royal Statistical Society. Series D (The Statistician), 45(1):3–19, ???? 1996. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348407.

Mackenzie:1996:RMS

[666] Gilbert Mackenzie. Regression models for survival data: The generalized time-dependent logistic family. Journal of the Royal Statistical Society. Series D (The Statistician), 45(1):21–34, ???? 1996. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348408.

Kelly:1996:ACT

[667] Gabrielle Kelly. Adaptive choice of tuning constant for robust regression estimators. Journal of the Royal Statistical Society. Series D (The Statistician), 45(1):35–40, ???? 1996. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348409.

Anonymous:1996:APPa

[668] Anonymous. [abstracts of papers presented at the Conference on Applied Statistics in Ireland, 1994]. Journal of the Royal Statistical Society. Series D (The Statistician), 45(1):41–48, ???? 1996. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348410.

Gilmour:1996:IMS

[669] Steven G. Gilmour. The interpretation of Mallows’s C_{p}-statistic. Journal of the Royal Statistical Society. Series D (The Statistician), 45(1):49–56, ???? 1996. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348411.

Duong:1996:BFS

[670] Quang P. Duong and Richard W. Shorrock. On Behrens–Fisher solutions. Journal of the Royal Statistical Society. Series D (The Statistician), 45(1):57–63, ???? 1996. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348412.

Lee:1996:EAT

[671] Kiseok Lee. An economic application of time domain Chebyshev filters: A test of the permanent income hypothesis. Journal of the Royal Statistical Society. Series D (The Statistician), 45(1):65–76, ???? 1996. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348413.

Henley:1996:NNC

[672] W. E. Henley and D. J. Hand. A knearest-neighbour classifier for assessing consumer credit risk. Journal of the Royal Statistical Society. Series D (The Statistician), 45(1):77–95, ???? 1996. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348414.

Flores:1996:QEM

[673] Renato G. Flôres, Jr. and Victor A. Ginsburgh. The Queen elisabeth musical competition: How fair is the final ranking? Journal of the Royal Statistical Society. Series D (The Statistician), 45(1):97–104, ???? 1996. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2348415.

Largey:1996:TMR

[674] Ann Largey and John E. Spencer. F - and t-tests in multiple regression: The possibility of ‘conflicting’ outcomes. Journal of the Royal Statistical Society. Series D (The Statistician), 45 (1):105–109, ???? 1996. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2348416.

George:1996:MSP

[675] Edward I. George and Samuel D. Oman. Multiple-shrinkage principal component regression. Journal of the Royal Statistical Society. Series D (The Statistician), 45(1):111–124, ???? 1996. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348417.

Weight:1996:C

[676] Terry Weight and Alan J. Collins. Correspondence. Journal of the Royal Statistical Society. Series D (The Statistician), 45(1):125–126, ???? 1996. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2348418.

Smith:1996:BRBa

[677] R. I. Smith. Book review: Statistics for the Environment 2-Water Related Issues, by V. Barnett; K. F. Turkman. Journal of the Royal Statistical Society. Series D (The Statistician), 45(1):127–128, ???? 1996. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348419.

Disney:1996:BRB

[678] John Disney. Book review: Achieving Quality Improvement — A Practical Guide, by R. Caulcutt. Journal of the Royal Statistical Society. Series D (The Statistician), 45(1):128, ???? 1996. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348420.

Brooks:1996:BRBa

[679] Stephen Brooks. Book review: Problem Solving: A Statistician’s Guide, by C. Chatfield. Journal of the Royal Statistical Society. Series D (The Statistician), 45(1):129, ???? 1996. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348421.

Wates:1996:BRB

[680] R. T. Wates. Book review: Statistics: Concepts and Applications, by H. Frank; S. Althoen; Instructor’s Manual for Statistics: Concepts and Application, by H. Frank; S. Althoen; User-Friendly-Workbook for ASP Tutorial for Statistics: Concepts and Applications, by H. Frank; S. Althoen. Journal of the Royal Statistical Society. Series D (The Statistician), 45 (1):129–130, ???? 1996. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348422.

Glendinning:1996:BRB

[681] R. H. Glendinning. Book review: Runs and Patterns in Probability: Selected Papers, by A. P. Godbole; S. G. Papastavridis. Journal of the Royal Statistical Society. Series D (The Statistician), 45(1):130, ???? 1996. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348423.

Oliver:1996:BRB

[682] F. R. Oliver. Book review: Teach Yourself Statistics, by A. Graham. Journal of the Royal Statistical Society. Series D (The Statistician), 45(1):131, ???? 1996. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348424.

Howarth:1996:BRB

[683] Richard J. Howarth. Book review: Computerized Basin Analysis: The Prognosis of Energy and Mineral Resources, by J. Harff; D. F. Merriam. Journal of the Royal Statistical Society. Series D (The Statistician), 45 (1):131–132, ???? 1996. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348425.

Robinson:1996:BRB

[684] David Robinson and Bao Lie Robinson. Book review: Bilingual (English/Chinese) Edition of Essentials of Statistical Methods, by T. P. Hutchinson; J. Z. Wang. Journal of the Royal Statistical Society. Series D (The Statistician), 45(1):132–133, ???? 1996. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348426.

Shoesmith:1996:BRB

[685] Eddie Shoesmith. Book review: Statistics for Accountants, by S. Letchford. Journal of the Royal Statistical Society. Series D (The Statistician), 45 (1):133–134, ???? 1996. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348427.

Brooks:1996:BRBb

[686] Stephen Brooks. Book review: Distribution-Free Statistical Methods, by J. S. Maritz. Journal of the Royal Statistical Society. Series D (The Statistician), 45(1):134, ???? 1996. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348428.

Gibson:1996:BRB

[687] Martin Gibson. Book review: Statistics in Engineering — A Practical Approach, by A. V. Metcalfe. Journal of the Royal Statistical Society. Series D (The Statistician), 45(1):134– 135, ???? 1996. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348429.

Metcalfe:1996:BRBa

[688] A. V. Metcalfe. Book review: Applied Bayesian Forecasting and Time Series Analysis, by A. Pole; M. West; J. Harrison. Journal of the Royal Statistical Society. Series D (The Statistician), 45(1):136–137, ???? 1996. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348430.

Cheal:1996:BRB

[689] Ryan Cheal and J. Warren. Book review: Diffusions, Markov Processes and Martingales, Vol. 1, Foundations, by L. C. G. Rogers; D. Williams. Journal of the Royal Statistical Society. Series D (The Statistician), 45(1):137, ???? 1996. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348431.

Swires-Hennessy:1996:BRB

[690] E. Swires-Hennessy. Book review: Getting into Print: A Guide for Scientists and Technologists, by P. Sprent. Journal of the Royal Statistical Society. Series D (The Statistician), 45 (1):137–138, ???? 1996. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348432.

Woodward:1996:BRB

[691] Phil Woodward. Book review: Statistics for Engineering Problem Solving, by S. B. Vardeman. Journal of the Royal Statistical Society. Series D (The Statistician), 45(1):138–139, ???? 1996. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348433.

Nunez-Anton:1996:BRB

[692] Vicente Núñez-Antón. Book review: Exact Statistical Methods for Data Analysis, by S. Weerahandi. Journal of the Royal Statistical Society. Series D (The Statistician), 45(1):139– 141, ???? 1996. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348434.

French:1996:BRB

[693] Simon French. Book review: Subjective Probability, by G. Wright; P. Ayton. Journal of the Royal Statistical Society. Series D (The Statistician), 45(1):141, ???? 1996. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2348435.

Li:1996:CMA

[694] W. K. Li and K. Lam. Correction: Modelling asymmetry in stock returns by a threshold autoregressive conditional heteroscedastic model. Journal of the Royal Statistical Society. Series D (The Statistician), 45(1):142, ???? 1996. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348436.

Anonymous:1996:BMa

[695] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 45(1):??, ???? 1996. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348437.

Anonymous:1996:FMa

[696] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 45(1):??, ???? 1996. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348405.

Ball:1996:TGP

[697] Michael Ball and Andrew Wood. Trend growth in post-1850 British economic history: The Kalman filter and historical judgment. Journal of the Royal Statistical Society. Series D (The Statistician), 45(2):143–152, ???? 1996. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2988403.

Mills:1996:MTE

[698] Terence C. Mills and N. F. R. Crafts. Modelling trends in economic history. Journal of the Royal Statistical Society. Series D (The Statistician), 45 (2):153–159, ???? 1996. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2988404.

Ball:1996:SCM

[699] Michael Ball and Andrew Wood. Some comments on modelling economic trends. Journal of the Royal Statistical Society. Series D (The Statistician), 45(2):161–162, ???? 1996. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2988405.

Gray:1996:ESN

[700] Rebecca Gray, Pamela Campanelli, Kavita Deepchand, and Patricia Prescott-Clarke. Exploring survey non-response: The effect of attrition on a follow-up of the 1984–85 Health and Life Style Survey. Journal of the Royal Statistical Society. Series D (The Statistician), 45(2):163– 183, ???? 1996. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988406.

Walls:1996:GAI

[701] Lesley Walls. A graphical approach to identification of dependent failures. Journal of the Royal Statistical Society. Series D (The Statistician), 45 (2):185–196, ???? 1996. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2988407.

Pettitt:1996:IEF

[702] A. N. Pettitt. Infinite estimates with fractional factorial experiments. Journal of the Royal Statistical Society. Series D (The Statistician), 45 (2):197–206, ???? 1996. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2988408.

Saporu:1996:ULM

[703] F. W. O. Saporu. On the use of the Leslie matrix in the analysis of prevalence data for general epidemic processes. Journal of the Royal Statistical Society. Series D (The Statistician), 45(2):207–221, ???? 1996. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2988409.

Lindsey:1996:EPL

[704] J. K. Lindsey and C. Laurent. Estimating the proportion of lymphoblastoid cells affected by exposure to ethylene oxide through micronuclei counts. Journal of the Royal Statistical Society. Series D (The Statistician), 45 (2):223–229, ???? 1996. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2988410.

Gillett:1996:RPS

[705] Raphael Gillett. Retrospective power surveys. Journal of the Royal Statistical Society. Series D (The Statistician), 45(2):231–236, ???? 1996. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2988411.

Ghazali:1996:MWC

[706] Syed Shakir Ali Ghazali. Modification of Waterton’s controlled selection method. Journal of the Royal Statistical Society. Series D (The Statistician), 45(2):237–242, ???? 1996. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2988412.

Blest:1996:LBA

[707] David C. Blest. Lower bounds for athletic performance. Journal of the Royal Statistical Society. Series D (The Statistician), 45(2):243–253, ???? 1996. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2988413.

Crowe:1996:CLB

[708] S. M. Crowe and Jennifer Middeldorp. A comparison of leg before wicket rates between Australians and their visiting teams for test cricket series played in Australia, 1977–94. Journal of the Royal Statistical Society. Series D (The Statistician), 45 (2):255–262, ???? 1996. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2988414.

Sprent:1996:C

[709] P. Sprent, A. Martin, I. Herranz, and A. Silva. Correspondence. Journal of the Royal Statistical Society. Series D (The Statistician), 45(2):263– 264, ???? 1996. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988415.

Moorthy:1996:BRB

[710] Uma Moorthy. Book review: A Casebook for a First Course in Statistics and Data Analysis, by S. Chatterjee; M. Handcock; J. S. Simonoff. Journal of the Royal Statistical Society. Series D (The Statistician), 45(2):265, ???? 1996. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988416.

Brooks:1996:BRBc

[711] Stephen Brooks. Book review: Bayesian Data Analysis, by A. Gelman; J. B. Carlin; H. S. Stern; D. B. Rubin. Journal of the Royal Statistical Society. Series D (The Statistician), 45(2):266–267, ???? 1996. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2988417.

Smith:1996:BRBb

[712] Murray D. Smith. Book review: Structural Equation Modeling: Concepts, Issues, and Applications, by R. H. Hoyle. Journal of the Royal Statistical Society. Series D (The Statistician), 45(2):267, ???? 1996. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2988418.

Greenblatt:1996:BRB

[713] Seth A. Greenblatt. Book review: Elementary Probability, by D. Stirzaker. Journal of the Royal Statistical Society. Series D (The Statistician), 45 (2):267–268, ???? 1996. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2988419.

Jolliffe:1996:BRB

[714] Flavia Jolliffe. Book review: An Introduction to Epidemiology, by T. C. Timmreck. Journal of the Royal Statistical Society. Series D (The Statistician), 45(2):268, ???? 1996. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2988420.

Anonymous:1996:BMb

[715] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 45(2):??, ???? 1996. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2988421.

Anonymous:1996:FMb

[716] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 45(2):??, ???? 1996. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2988402.

Alanko:1996:CBD

[717] Timo Alanko and John C. Duffy. Compound binomial distributions for modelling consumption data. Journal of the Royal Statistical Society. Series D (The Statistician), 45(3):269–286, ???? 1996. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988465.

Vandenbroeck:1996:SNB

[718] P. Vandenbroeck and P. Vandevyvere. Statistics in a new business environment: An example. Journal of the Royal Statistical Society. Series D (The Statistician), 45(3):287– 292, ???? 1996. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988466.

Walker:1996:BMP

[719] Stephen Walker. A Bayesian maximum a posteriori algorithm for categorical data under informative general censoring. Journal of the Royal Statistical Society. Series D (The Statistician), 45(3):293–298, ???? 1996. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2988467.

Gillett:1996:SFM

[720] Raphael Gillett. Sampling with fallible memory: An occupancy problem. Journal of the Royal Statistical Society. Series D (The Statistician), 45(3):299–305, ???? 1996. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2988468.

Chen:1996:MCM

[721] H. H. Chen, S. W. Duffy, and Laszlo Tabar. A Markov chain method to estimate the tumour progression rate from preclinical to clinical phase, sensitivity and positive predictive value for mammography in breast cancer screening. Journal of the Royal Statistical Society. Series D (The Statistician), 45(3):307–317, ???? 1996. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2988469.

Kumbhakar:1996:ECE

[722] Subal C. Kumbhakar. Estimation of cost efficiency with heteroscedasticity: An application to electricity utilities. Journal of the Royal Statistical Society. Series D (The Statistician), 45 (3):319–335, ???? 1996. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2988470.

Lyon:1996:CTH

[723] John D. Lyon and Chih-Ling Tsai. A comparison of tests for heteroscedasticity. Journal of the Royal Statistical Society. Series D (The Statistician), 45(3):337–349, ???? 1996. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2988471.

Chakraborti:1996:PTC

[724] S. Chakraborti and P. van der Laan. Precedence tests and confidence bounds for complete data: An overview and some results. Journal of the Royal Statistical Society. Series D (The Statistician), 45(3):351– 369, ???? 1996. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2988472.

Schmid:1996:TFO

[725] Friedrich Schmid and Mark Trede. Testing for first-order stochastic dominance: A new distribution-free test. Journal of the Royal Statistical Society. Series D (The Statistician), 45 (3):371–380, ???? 1996. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2988473.

Bland:1996:C

[726] Nicholas D. Bland, J. Martin Bland, Stephen R. Clarke, and John M. Norman. Correspondence. Journal of the Royal Statistical Society. Series D (The Statistician), 45(3):381–383, ???? 1996. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988474.

Coolen:1996:BRB

[727] Frank Coolen. Book review: Statistical Models Based on Counting Processes, by P. K. Andersen; Ø. Borgan; R. D. Gill; N. Keiding. Journal of the Royal Statistical Society. Series D (The Statistician), 45(3):384, ???? 1996. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2988475.

Lindsey:1996:BRB

[728] J. K. Lindsey. Book review: The Statistical Analysis of Time Series, by T. W. Anderson. Journal of the Royal Statistical Society. Series D (The Statistician), 45(3):385, ???? 1996. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2988476.

Wheeler:1996:BRB

[729] Frederick P. Wheeler. Book review: Chaos and Catastrophe Theory, by C. Brown. Journal of the Royal Statistical Society. Series D (The Statistician), 45(3):385–386, ???? 1996. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2988477.

Crowder:1996:BRB

[730] Martin Crowder. Book review: Statistical Inference, by P. H. Garthwaite; I. T. Jolliffe; B. Jones. Journal of the Royal Statistical Society. Series D (The Statistician), 45(3):386– 387, ???? 1996. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988478.

White:1996:BRB

[731] Allan White. Book review: Ecology of Infectious Diseases in Natural Populations, by B. T. Grenfell; A. P. Dobson. Journal of the Royal Statistical Society. Series D (The Statistician), 45(3):387–388, ???? 1996. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2988479.

Bartels:1996:BRB

[732] Robert Bartels. Book review: XploRe: An Interactive Statistical Computing Environment, by W. Härdle; S. Klinke; B. A. Turlach. Journal of the Royal Statistical Society. Series D (The Statistician), 45(3):388–389, ???? 1996. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988480.

Byrne:1996:BRB

[733] Patrick Byrne. Book review: Mathematical Models in Finance, by S. D. Howison; F. P. Kelly; P. Wilmott. Journal of the Royal Statistical Society. Series D (The Statistician), 45(3): 389, ???? 1996. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988481.

Porter:1996:BRB

[734] Max Porter. Book review: Multiobjective Decision Support for Environmental Management, by R. Janssen. Journal of the Royal Statistical Society. Series D (The Statistician), 45 (3):389–390, ???? 1996. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2988482.

Gordon:1996:BRB

[735] A. D. Gordon. Book review: Recent Advances in Descriptive Multivariate Analysis, by W. J. Krzanowski. Journal of the Royal Statistical Society. Series D (The Statistician), 45 (3):390–391, ???? 1996. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2988483.

Watkins:1996:BRBa

[736] James F. Watkins. Book review: Analysing Survival Data from Clinical Trials and Observational Studies, by E. Marubini; M.-G. Valsecchi. Journal of the Royal Statistical Society. Series D (The Statistician), 45(3):391–392, ???? 1996. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2988484.

Sedgwick:1996:BRB

[737] Philip Sedgwick. Book review: Statistical Methods — An Introductory Text, by J. Medhi. Journal of the Royal Statistical Society. Series D (The Statistician), 45(3):392–393, ???? 1996. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2988485.

Csenki:1996:BRB

[738] A. Csenki. Book review: Stochastic Processes, by J. Medhi. Journal of the Royal Statistical Society. Series D (The Statistician), 45(3):393– 394, ???? 1996. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2988486.

Anonymous:1996:BMc

[739] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 45(3):??, ???? 1996. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2988487.

Anonymous:1996:FMc

[740] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 45(3):??, ???? 1996. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2988464.

Barry:1996:F

[741] Don Barry and Tom O’Donovan. Foreword. Journal of the Royal Statistical Society. Series D (The Statistician), 45(4):395, ???? 1996. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2988541.

Dunne:1996:TSV

[742] Adrian Dunne, Yudi Pawitan, and Liam Doody. Two-sided P -values from discrete asymmetric distributions based on uniformly most powerful unbiased tests. Journal of the Royal Statistical Society. Series D (The Statistician), 45(4):397–405, ???? 1996. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2988542.

Haberman:1996:GLM

[743] Steven Haberman and Arthur E. Renshaw. Generalized linear models and actuarial science. Journal of the Royal Statistical Society. Series D (The Statistician), 45(4):407–436, ???? 1996. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988543.

Boland:1996:SPC

[744] Philip J. Boland and James Nicholson. The statistics and probability curriculum at the secondary school level in the USA, Ireland and the UK. Journal of the Royal Statistical Society. Series D (The Statistician), 45 (4):437–446, ???? 1996. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2988544.

Anonymous:1996:APPb

[745] Anonymous. [abstracts of papers presented at the Conference on Applied Statistics in Ireland, 1995]. Journal of the Royal Statistical Society. Series D (The Statistician), 45(4):447– 455, ???? 1996. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2988545.

Walley:1996:ACD

[746] Peter Walley, Lyle Gurrin, and Paul Burton. Analysis of clinical data using imprecise prior probabilities. Journal of the Royal Statistical Society. Series D (The Statistician), 45(4):457– 485, ???? 1996. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2988546.

Bernardo:1996:GMB

[747] José M. Bernardo and Telba Z. Irony. A general multivariate Bayesian process capability index. Journal of the Royal Statistical Society. Series D (The Statistician), 45(4):487–502, ???? 1996. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2988547.

Makov:1996:BMA

[748] U. E. Makov, A. F. M. Smith, and Y.- H. Liu. Bayesian methods in actuarial science. Journal of the Royal Statistical Society. Series D (The Statistician), 45(4):503–515, ???? 1996. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2988548.

Smeeton:1996:OWT

[749] Nigel Smeeton. Open wide: Teaching statistics to dental students. Journal of the Royal Statistical Society. Series D (The Statistician), 45(4):517– 520, ???? 1996. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988549.

Campbell:1996:NTU

[750] Bryan Campbell and John W. Galbraith. Nonparametric tests of the unbiasedness of Olympic figure-skating judgments. Journal of the Royal Statistical Society. Series D (The Statistician), 45(4):521–526, ???? 1996. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2988550.

Anonymous:1996:EA

[751] Anonymous. Editorial acknowledgement. Journal of the Royal Statistical Society. Series D (The Statistician), 45(4):527, ???? 1996. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2988551.

Baxter:1996:BRB

[752] L. A. Baxter. Book review: The Elements of Integration and Lebesgue Measure, by R. G. Bartle. Journal of the Royal Statistical Society. Series D (The Statistician), 45(4):528– 529, ???? 1996. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988552.

Prys-Williams:1996:BRB

[753] Allan G. Prys-Williams. Book review: Testing Regression Models Based on Sample Survey Data, by M. I. Bhatti. Journal of the Royal Statistical Society. Series D (The Statistician), 45(4):529, ???? 1996. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2988553.

Watkins:1996:BRBb

[754] James F. Watkins. Book review: Nonlinear Models for Repeated Data Measurement, by M. Davidian; D. M. Giltinan. Journal of the Royal Statistical Society. Series D (The Statistician), 45(4):529–531, ???? 1996. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2988554.

OBrien:1996:BRB

[755] C. M. O’Brien. Book review: Introduction to Graphical Modelling, by D. Edwards. Journal of the Royal Statistical Society. Series D (The Statistician), 45(4):531–532, ???? 1996. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2988555.

Stein:1996:BRB

[756] A. Stein. Book review: The Bootstrap and Edgeworth Expansion, by P. Hall. Journal of the Royal Statistical Society. Series D (The Statistician), 45(4):532, ???? 1996. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2988556.

Metcalfe:1996:BRBb

[757] A. V. Metcalfe. Book review: Introduction to Stochastic Processes, by G. F. Lawler. Journal of the Royal Statistical Society. Series D (The Statistician), 45(4):533, ???? 1996. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2988557.

Clark:1996:BRB

[758] Stephen D. Clark. Book review: Introductory Statistics — A Modelling Approach, by J. K. Lindsey. Journal of the Royal Statistical Society. Series D (The Statistician), 45(4):533– 534, ???? 1996. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988558.

Jessen:1996:BRB

[759] Hans Christian Jessen. Book review: Applied Logistic Regression Analysis, by S. Menard. Journal of the Royal Statistical Society. Series D (The Statistician), 45(4):534–535, ???? 1996. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2988559.

Everitt:1996:BRB

[760] Brian Everitt. Book review: Methods of Multivariate Analysis, by A. C. Rencher. Journal of the Royal Statistical Society. Series D (The Statistician), 45(4):535–536, ???? 1996. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2988560.

Lindley:1996:BRB

[761] D. V. Lindley. Book review: Theory of Statistics, by M. J. Schervish. Journal of the Royal Statistical Society. Series D (The Statistician), 45 (4):536–538, ???? 1996. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2988561.

Thomas:1996:BRB

[762] Ray Thomas. Book review: Interpreting Crime Statistics, by A. Walker. Journal of the Royal Statistical Society. Series D (The Statistician), 45 (4):538–539, ???? 1996. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2988562.

Kadane:1996:CSB

[763] Joseph B. Kadane. Correction: Subjective Bayesian analysis for surveys with missing data. Journal of the Royal Statistical Society. Series D (The Statistician), 45(4):539, ???? 1996. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2988563.

Anonymous:1996:VI

[764] Anonymous. Volume information. Journal of the Royal Statistical Society. Series D (The Statistician), 45 (4):540–542, ???? 1996. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2988539.

Anonymous:1996:BMd

[765] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 45(4):??, ???? 1996. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2988564.

Anonymous:1996:FMd

[766] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 45(4):??, ???? 1996. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2988540.

Mannion:1997:AAC

[767] David Mannion and Peter Dixon. Authorship attribution: The case of Oliver Goldsmith. Journal of the Royal Statistical Society. Series D (The Statistician), 46(1):1–18, ???? 1997. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2988489.

Janacek:1997:CCB

[768] G. J. Janacek and S. E. Meikle. Control charts based on medians. Journal of the Royal Statistical Society. Series D (The Statistician), 46(1):19–31, ???? 1997. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2988490.

Garvin:1997:CST

[769] J. S. Garvin and S. I. McClean. Convolution and sampling theory of the binormal distribution as a prerequisite to its application in statistical process control. Journal of the Royal Statistical Society. Series D (The Statistician), 46(1):33–47, ???? 1997. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2988491.

Wang:1997:ORE

[770] Yuedong Wang. Odds ratio estimation in Bernoulli smoothing spline analysisof-variance models. Journal of the Royal Statistical Society. Series D (The Statistician), 46(1):49–56, ???? 1997. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2988492.

Coutts:1997:PSM

[771] J. Andrew Coutts, Jennifer Roberts, and Terence C. Mills. Parameter stability in the market model: Tests and time varying parameter estimation with UK data. Journal of the Royal Statistical Society. Series D (The Statistician), 46(1):57–70, ???? 1997. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2988493.

Galea:1997:LIE

[772] Manuel Galea, Gilberto A. Paula, and Heleno Bolfarine. Local influence in elliptical linear regression models. Journal of the Royal Statistical Society. Series D (The Statistician), 46(1):71–79, ???? 1997. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2988494.

Fitzmaurice:1997:MSO

[773] Garrett M. Fitzmaurice. Model selection with overdispersed data. Journal of the Royal Statistical Society. Series D (The Statistician), 46(1):81–91, ???? 1997. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2988495.

Jeevanand:1997:BEB

[774] E. S. Jeevanand. Bayes estimation of P (X_{2} < X_{1}) for a bivariate Pareto distribution. Journal of the Royal Statistical Society. Series D (The Statistician), 46(1):93–99, ???? 1997. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2988496.

Morton:1997:SEL

[775] R. Hugh Morton. Statistical effects of lane allocation on times in running races. Journal of the Royal Statistical Society. Series D (The Statistician), 46(1):101–104, ???? 1997. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2988497.

Wright:1997:FSM

[776] Daniel B. Wright. Football standings and measurement levels. Journal of the Royal Statistical Society. Series D (The Statistician), 46(1):105– 110, ???? 1997. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988498.

Morton:1997:C

[777] R. Hugh Morton and David Best. Correspondence. Journal of the Royal Statistical Society. Series D (The Statistician), 46(1):111–113, ???? 1997. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2988499.

Everitt:1997:BRBa

[778] B. Everitt. Book review: An Introduction to Categorical Data Analysis, by A. Agresti. Journal of the Royal Statistical Society. Series D (The Statistician), 46(1):115, ???? 1997. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2988500.

Upendra:1997:BRB

[779] Dave Upendra. Book review: Measurement Error in Nonlinear Models, by R. J. Carroll; D. Ruppert; L. A. Stefanski. Journal of the Royal Statistical Society. Series D (The Statistician), 46(1):115–116, ???? 1997. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2988501.

Congdon:1997:BRB

[780] Peter Congdon. Book review: Clinical Biostatistics: An Introduction to Evidence-Based Medicine, by G. Dunn; B. Everitt. Journal of the Royal Statistical Society. Series D (The Statistician), 46(1):116–117, ???? 1997. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2988502.

Brooks:1997:BRB

[781] Stephen Brooks. Book review: Markov Chain Monte Carlo in Practice, by W. R. Gilks; S. Richardson; D. J. Spiegelhalter. Journal of the Royal Statistical Society. Series D (The Statistician), 46(1):117–118, ???? 1997. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2988503.

Grubb:1997:BRB

[782] Howard Grubb. Book review: Stochastic Modelling of Scientific Data, by P. Guttorp. Journal of the Royal Statistical Society. Series D (The Statistician), 46(1):118–119, ???? 1997. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2988504.

Martin:1997:BRBa

[783] R. J. Martin. Book review: Cyclic and Computer Generated Designs, by J. A. John; E. R. Williams. Journal of the Royal Statistical Society. Series D (The Statistician), 46(1):119, ???? 1997. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2988505.

Lintott:1997:BRB

[784] John Lintott. Book review: Measuring the Economy: A Guide to Understanding Official Statistics, by C. Johnson; S. Briscoe. Journal of the Royal Statistical Society. Series D (The Statistician), 46(1):119–120, ???? 1997. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2988506.

Currall:1997:BRB

[785] James Currall. Book review: Statistics: An Interactive Text for the Health and Life Sciences, by G. B. Krishnamurty; P. Kasovia-Schmitt; D. J. Ostroff. Journal of the Royal Statistical Society. Series D (The Statistician), 46(1):120–121, ???? 1997. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2988507.

Starkings:1997:BRBa

[786] Susan Starkings. Book review: Models for Uncertainty in Educational Testing, by N. T. Longford. Journal of the Royal Statistical Society. Series D (The Statistician), 46(1):121– 122, ???? 1997. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988508.

Wheatley:1997:BRB

[787] Alison Wheatley. Book review: Basic Epidemiological Methods and Biostatistics: A Practical Guidebook, by R. M. Page; G. E. Cole; T. C. Timmreck. Journal of the Royal Statistical Society. Series D (The Statistician), 46(1):122–123, ???? 1997. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2988509.

Everitt:1997:BRBb

[788] B. Everitt. Book review: Workshop Statistics Discovery with Data, by A. J. Rossman. Journal of the Royal Statistical Society. Series D (The Statistician), 46(1):123, ???? 1997. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2988510.

Jolliffe:1997:BRB

[789] Ian Jolliffe. Book review: Multivariate Geostatistics: An Introduction with Applications, by H. Wackernagel. Journal of the Royal Statistical Society. Series D (The Statistician), 46(1):124, ???? 1997. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2988511.

Castagliola:1997:BRB

[790] Philippe Castagliola. Book review: Kernel Smoothing, by M. P. Wand; M. C. Jones. Journal of the Royal Statistical Society. Series D (The Statistician), 46(1):124–125, ???? 1997. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2988512.

Anonymous:1997:BMa

[791] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 46(1):126, ???? 1997. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988513.

Anonymous:1997:FMa

[792] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 46(1):??, ???? 1997. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2988488.

Smeeton:1997:F

[793] N. C. Smeeton and C. J. Adcock. Foreword. Journal of the Royal Statistical Society. Series D (The Statistician), 46(2):127, ???? 1997. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2988515.

Lindley:1997:CSSa

[794] Dennis V. Lindley. The choice of sample size. Journal of the Royal Statistical Society. Series D (The Statistician), 46(2):129–138, ???? 1997. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2988516. See com-

ments [798].

Pham-Gia:1997:BAB

[795] T. Pham-Gia. On Bayesian analysis, Bayesian decision theory and the sample size problem. Journal of the Royal Statistical Society. Series D (The Statistician), 46(2):139– 144, ???? 1997. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988517.

Joseph:1997:IBV

[796] Lawrence Joseph and David B. Wolfson. Interval-based versus decision theoretic criteria for the choice of sample size. Journal of the Royal Statistical Society. Series D (The Statistician), 46(2):145–149, ???? 1997. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2988518.

Bernardo:1997:SID

[797] José M. Bernardo. Statistical inference as a decision problem: The choice of sample size. Journal of the Royal Statistical Society. Series D (The Statistician), 46(2):151–153, ???? 1997. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2988519.

Adcock:1997:CSS

[798] C. J. Adcock. The choice of sample size and the method of maximum expected utility — comments on the paper by Lindley. Journal of the Royal Statistical Society. Series D (The Statistician), 46(2):155–162, ???? 1997. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2988520. See

[794].

Lindley:1997:CSSb

[799] Dennis V. Lindley. The choice of sample size — a reply to the discussion. Journal of the Royal Statistical Society. Series D (The Statistician), 46 (2):163–166, ???? 1997. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2988521.

Goldstein:1997:CSS

[800] Michael Goldstein and David A. Wooff. Choosing sample sizes in balanced experimental designs: A Bayes linear approach. Journal of the Royal Statistical Society. Series D (The Statistician), 46(2):167–183, ???? 1997. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2988522.

Weiss:1997:BSS

[801] Robert Weiss. Bayesian sample size calculations for hypothesis testing. Journal of the Royal Statistical Society. Series D (The Statistician), 46(2):185– 191, ???? 1997. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988523.

Yeh:1997:BDS

[802] Lam Yeh and Lam Chi Van. Bayesian double-sampling plans with normal distributions. Journal of the Royal Statistical Society. Series D (The Statistician), 46(2):193–207, ???? 1997. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2988524.

Joseph:1997:BSS

[803] Lawrence Joseph and Patrick Bélisle. Bayesian sample size determination for normal means and differences between normal means. Journal of the Royal Statistical Society. Series D (The Statistician), 46(2):209–226, ???? 1997. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2988525.

Thode:1997:PSS

[804] Henry C. Thode, Jr. Power and sample size requirements for tests of differences between two Poisson rates. Journal of the Royal Statistical Society. Series D (The Statistician), 46 (2):227–230, ???? 1997. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2988526.

Lindsey:1997:ESS

[805] J. K. Lindsey. Exact sample size calculations for exponential family models. Journal of the Royal Statistical Society. Series D (The Statistician), 46(2):231–237, ???? 1997. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2988527.

Liu:1997:SSS

[806] W. Liu. On some sample size formulae for controlling both size and power in clinical trials. Journal of the Royal Statistical Society. Series D (The Statistician), 46(2):239–251, ???? 1997. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2988528.

Sparks:1997:SSE

[807] T. H. Sparks, J. O. Mountford, S. J. Manchester, P. Rothery, and J. R. Treweek. Sample size for estimating species lists in vegetation surveys. Journal of the Royal Statistical Society. Series D (The Statistician), 46 (2):253–260, ???? 1997. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2988529.

Adcock:1997:BRB

[808] C. J. Adcock. Book review: Sample Size Determination: A Review. Journal of the Royal Statistical Society. Series D (The Statistician), 46 (2):261–283, ???? 1997. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2988530.

Smith:1997:BRB

[809] Murray D. Smith. Book review: Regression Models: Censored, Sample Selected, or Truncated Data, by R. Breen. Journal of the Royal Statistical Society. Series D (The Statistician), 46(2):285, ???? 1997. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2988531.

Militino:1997:BRB

[810] Ana F. Militino. Book review: Multivariate Dependencies, Models, Analysis and Interpretation, by D. R. Cox; N. Wermuth. Journal of the Royal Statistical Society. Series D (The Statistician), 46(2):285–286, ???? 1997. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2988532.

Everitt:1997:BRBc

[811] B. Everitt. Book review: Challenges Ahead for Improving Sports Statistics, by International Statistical Institute Sports Statistics Committee and International Statistical Institute. Journal of the Royal Statistical Society. Series D (The Statistician), 46(2):287, ???? 1997. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2988534.

Lindley:1997:BRBa

[812] D. V. Lindley. Book review: Life’s Grandeur: The Spread of Excellence from Plato to Darwin, by S. J. Gould. Journal of the Royal Statistical Society. Series D (The Statistician), 46(2): 287, ???? 1997. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2988533.

Coolen:1997:BRBa

[813] Frank Coolen. Book review: Survival Analysis: A Self-Learning Text, by D. G. Kleinbaum. Journal of the Royal Statistical Society. Series D (The Statistician), 46(2):288–289, ???? 1997. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2988535.

Scott:1997:BRBa

[814] Marian Scott. Book review: VARI-OWIN: Software for Spatial Data Analysis in 2D, by Y. Pannatier. Journal of the Royal Statistical Society. Series D (The Statistician), 46(2):289– 290, ???? 1997. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2988536.

Barbour:1997:BRB

[815] A. D. Barbour. Book review: Introduction to Computational Biology: Maps, Sequences and Genomes, by M. S. Waterman. Journal of the Royal Statistical Society. Series D (The Statistician), 46(2):290–291, ???? 1997. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2988537.

Anonymous:1997:BMb

[816] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 46(2):292, ???? 1997. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988538.

Anonymous:1997:FMb

[817] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 46(2):??, ???? 1997. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2988514.

Schluter:1997:RSM

[818] P. J. Schlüter, J. J. Deely, and A. J. Nicholson. Ranking and selecting motor vehicle accident sites by using a hierarchical Bayesian model. Journal of the Royal Statistical Society. Series D (The Statistician), 46(3):293– 316, ???? 1997. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988566.

Sheldon:1997:VPR

[819] Aaron Sheldon and Robert Sproule. Variance as a proxy for risk: The case of the binomial distribution. Journal of the Royal Statistical Society. Series D (The Statistician), 46(3):317– 322, ???? 1997. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2988567.

Hawkins:1997:IGC

[820] Douglas M. Hawkins and David H. Olwell. Inverse Gaussian cumulative sum control charts for location and shape. Journal of the Royal Statistical Society. Series D (The Statistician), 46(3):323–335, ???? 1997. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2988568.

Rayner:1997:AOT

[821] J. C. W. Rayner. The asymptotically optimal tests. Journal of the Royal Statistical Society. Series D (The Statistician), 46(3):337–346, ???? 1997. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2988569.

Best:1997:PMR

[822] D. J. Best and J. C. W. Rayner. Product maps for ranked preference data. Journal of the Royal Statistical Society. Series D (The Statistician), 46 (3):347–354, ???? 1997. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2988570.

Castillo:1997:FCB

[823] Enrique Castillo, José María Sarabia, and Ali S. Hadi. Fitting continuous bivariate distributions to data. Journal of the Royal Statistical Society. Series D (The Statistician), 46 (3):355–369, ???? 1997. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2988571.

Woodward:1997:STE

[824] Mark Woodward, Neil Dourmashkin, Etienne Twagirumukiza, Maurice C. Y. Mbago, and Adrião S. Ferreira da Cunha. Statistics training in Eastern and Southern Africa: A study of supply and demand. Journal of the Royal Statistical Society. Series D (The Statistician), 46(3):371–386, ???? 1997. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988572.

Johnson:1997:TDP

[825] David Johnson. The triangular distribution as a proxy for the beta distribution in risk analysis. Journal of the Royal Statistical Society. Series D (The Statistician), 46(3):387– 398, ???? 1997. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988573.

Ogwang:1997:SAE

[826] Tomson Ogwang and U. L. Gouranga Rao. A simple algorithm for estimating Box–Cox models. Journal of the Royal Statistical Society. Series D (The Statistician), 46(3):399–409, ???? 1997. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988574.

Royston:1997:ASF

[827] Patrick Royston and Douglas G. Altman. Approximating statistical functions by using fractional polynomial regression. Journal of the Royal Statistical Society. Series D (The Statistician), 46(3):411–422, ???? 1997. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2988575.

Smith:1997:PSS

[828] Philip J. Smith. Power and sample size considerations for detecting deviations from secular trends in surveillance surveys. Journal of the Royal Statistical Society. Series D (The Statistician), 46(3):423–432, ???? 1997. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2988576.

Walters:1997:CLL

[829] D. E. Walters. Confidence limits and log-odds ratios. Journal of the Royal Statistical Society. Series D (The Statistician), 46(3):433–438, ???? 1997. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2988577.

Csenki:1997:OAS

[830] Attila Csenki. Optimum allocation in stratified random sampling via Hölder’s inequality. Journal of the Royal Statistical Society. Series D (The Statistician), 46(3):439–441, ???? 1997. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2988578.

Longford:1997:C

[831] Nick Longford, S. R. Clarke, J. M. Norman, J. M. Bland, and Gerard J. P. Van Breukelen. Correspondence.

Journal of the Royal Statistical Society. Series D (The Statistician), 46 (3):443–447, ???? 1997. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2988579.

Seuc:1997:BRB

[832] Armando Seuc. Book review: Statistics as Principled Argument, by R. P. Abelson. Journal of the Royal Statistical Society. Series D (The Statistician), 46(3):448–449, ???? 1997. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2988580.

Scott:1997:BRBb

[833] Marian Scott. Book review: Bayesian Approach to Interpreting Archaeological Data, by C. E. Buck; W. G. Cavanagh; C. D. Litton. Journal of the Royal Statistical Society. Series D (The Statistician), 46(3):449– 450, ???? 1997. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988581.

Martin:1997:BRBb

[834] R. J. Martin. Book review: Introduction to Statistical Time Series, by W. A. Fuller. Journal of the Royal Statistical Society. Series D (The Statistician), 46(3):450–451, ???? 1997. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2988582.

Nunez-Anton:1997:BRB

[835] Vicente Núñez-Antón. Book review: Practical Longitudinal Data Analysis, by D. Hand; M. Crowder. Journal of the Royal Statistical Society. Series D (The Statistician), 46(3):451– 452, ???? 1997. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988583.

Gani:1997:BRBa

[836] J. Gani. Book review: Models for Infectious Human Diseases: Their Structure and Relation to Data, by V. Isham; G. Medley. Journal of the Royal Statistical Society. Series D (The Statistician), 46(3):453, ???? 1997. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2988584.

Gani:1997:BRBb

[837] J. Gani. Book review: Lifetime Data: Models in Reliability and Survival Analysis, by N. P. Jewell; A. C. Kimber; M.-L. T. Lee; G. A. Whitmore. Journal of the Royal Statistical Society. Series D (The Statistician), 46(3):454, ???? 1997. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2988585.

Lindley:1997:BRBb

[838] D. V. Lindley. Book review: Bayesian Methods and Ethics in a Clinical Trial Design, by J. B. Kadane. Journal of the Royal Statistical Society. Series D (The Statistician), 46(3):454– 455, ???? 1997. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988586.

Branco:1997:BRB

[839] Joao A. Branco. Book review: Multivariate Analysis,Part 2: Classification, Covariance Structures and Repeated Measurements, by W. J. Krzanowski; F. H. C. Marriott. Journal of the Royal Statistical Society. Series D (The Statistician), 46(3):455–456, ???? 1997. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988587.

Firmin:1997:BRB

[840] Paul E. Firmin. Book review: Fundamentals of Applied Statistics and Surveys, by D. B. Orr. Journal of the Royal Statistical Society. Series D (The Statistician), 46(3):457, ???? 1997. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2988588.

Everitt:1997:BRBd

[841] B. Everitt. Book review: Activity Based Statistics, by R. L. Scheaffer; M. Gnanadesikan; A. Watkins; J. A. Witmer. Journal of the Royal Statistical Society. Series D (The Statistician), 46(3):457–458, ???? 1997. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2988589.

Everitt:1997:BRBe

[842] B. Everitt. Book review: Graphing Statistics and Data, by A. Wallgren; B. Wallgren; R. Persson; U. Jorner; J. A. Haaland. Journal of the Royal Statistical Society. Series D (The Statistician), 46(3):458, ???? 1997. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2988590.

Anonymous:1997:BMc

[843] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 46(3):??, ???? 1997. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2988591.

Anonymous:1997:FMc

[844] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 46(3):??, ???? 1997. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2988565.

Beattie:1997:F

[845] Robert Beattie, Edward Goodall, and Gil Stevenson. Foreword. Journal of the Royal Statistical Society. Series D (The Statistician), 46(4):459, ???? 1997. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2988594.

Chatfield:1997:F

[846] Chris Chatfield. Forecasting in the 1990s. Journal of the Royal Statistical Society. Series D (The Statistician), 46(4):461–473, ???? 1997. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2988595.

Haslett:1997:SVS

[847] John Haslett. On the sample variogram and the sample autocovariance for non-stationary time series. Journal of the Royal Statistical Society. Series D (The Statistician), 46 (4):475–485, ???? 1997. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2988596.

Conniffe:1997:ILR

[848] Denis Conniffe. Improving a linear regression through joint estimation with a probit model. Journal of the Royal Statistical Society. Series D (The Statistician), 46(4):487– 493, ???? 1997. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988597.

Watson:1997:ESE

[849] Sally Watson. Evaluation of semivariance estimators under normal conditions. Journal of the Royal Statistical Society. Series D (The Statistician), 46(4):495–503, ???? 1997. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2988598.

Boniface:1997:ASE

[850] David R. Boniface and Margaret E. Tefft. The application of structural equation modelling to the construction of an index for the measurement of health-related behaviours. Journal of the Royal Statistical Society. Series D (The Statistician), 46(4):505– 514, ???? 1997. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988599.

Hay:1997:SMD

[851] Gordon Hay. The selection from multiple data sources in epidemiological capture-recapture studies. Journal of the Royal Statistical Society. Series D (The Statistician), 46(4):515–520, ???? 1997. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2988600.

Smeeton:1997:SEM

[852] Nigel Smeeton. Statistical education in medicine and dentistry. Journal of the Royal Statistical Society. Series D (The Statistician), 46(4):521– 527, ???? 1997. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2988601.

Chedzoy:1997:EUE

[853] Olaf B. Chedzoy. The effect of umpiring errors in Cricket. Journal of the Royal Statistical Society. Series D (The Statistician), 46(4):529–540, ???? 1997. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2988602.

Pollard:1997:MEP

[854] Richard Pollard and Charles Reep. Measuring the effectiveness of playing strategies at soccer. Journal of the Royal Statistical Society. Series D (The Statistician), 46(4):541–550, ???? 1997. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2988603.

Holder:1997:MPI

[855] Roger L. Holder and Alan M. Nevill. Modelling performance at international tennis and golf tournaments: Is there a home advantage? Journal of the Royal Statistical Society. Series D (The Statistician), 46(4):551–559, ???? 1997. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988604.

Ugarte:1997:BRB

[856] Lola Ugarte. Book review: Statistical Inference-Based on the Likelihood, by A. Azzalini. Journal of the Royal Statistical Society. Series D (The Statistician), 46(4):561–562, ???? 1997. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2988605.

Starkings:1997:BRBb

[857] Susan Starkings. Book review: Statistical Tests — An Introduction with Minitab Commentary, by G. P. Beaumont; J. D. Knowles. Journal of the Royal Statistical Society. Series D (The Statistician), 46(4):562, ???? 1997. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2988606.

Scott:1997:BRBc

[858] Marion Scott. Book review: Introduction to Time Series and Forecasting, by P. J. Brockwell; R. A. Davis. Journal of the Royal Statistical Society. Series D (The Statistician), 46 (4):562–563, ???? 1997. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2988607.

Starkings:1997:BRBc

[859] Susan Starkings. Book review: An Introduction to Queueing Theory, by B. D. Bunday. Journal of the Royal Statistical Society. Series D (The Statistician), 46(4):563–564, ???? 1997. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2988608.

Clark:1997:BRB

[860] Stephen D. Clark. Book review: The Analysis of Time Series: An Introduction, by C. Chatfield. Journal of the Royal Statistical Society. Series D (The Statistician), 46(4):564– 565, ???? 1997. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2988609.

Foulkes:1997:BRB

[861] Mary A. Foulkes. Book review: Introduction to the Design and Analysis of Experiments, by G. M. Clarke; R. E. Kempson. Journal of the Royal Statistical Society. Series D (The Statistician), 46(4):565–566, ???? 1997. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2988610.

Woodward:1997:BRB

[862] Phil Woodward. Book review: The Pocket Statistician: A Practical Guide to Quality Improvement, by S. Coleman; T. Greenfield; R. Jones; C. Morris; I. Puzey. Journal of the Royal Statistical Society. Series D (The Statistician), 46(4):566–567, ???? 1997. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2988611.

Watkins:1997:BRB

[863] James Watkins. Book review: Biplots, by J. C. Gower; D. J. Hand.

Journal of the Royal Statistical Society. Series D (The Statistician), 46 (4):567–568, ???? 1997. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2988612.

Wheeler:1997:BRB

[864] Frederick P. Wheeler. Book review: Research Methods: Guidance for Postgraduates, by T. Greenfield. Journal of the Royal Statistical Society. Series D (The Statistician), 46(4):569, ???? 1997. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988613.

Coolen:1997:BRBb

[865] Frank Coolen. Book review: Maximum Entropy and Bayesian Methods, by G. R. Heidbreder. Journal of the Royal Statistical Society. Series D (The Statistician), 46(4):569– 570, ???? 1997. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988614.

Bartholomew:1997:BRB

[866] D. J. Bartholomew. Book review: Latent Class and Discrete Latent Trait Models: Similarities and Differences, by T. Heinen. Journal of the Royal Statistical Society. Series D (The Statistician), 46(4):570–571, ???? 1997. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2988615.

Gani:1997:BRBc

[867] J. Gani. Book review: Discrete-Time Markov Control Processes: Basic Optimality Criteria, by O. Hernandez-Lerma; J. B. Lasserre. Journal of the Royal Statistical Society. Series D (The Statistician), 46(4):571–572, ???? 1997. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2988616.

Lindley:1997:BRBc

[868] D. V. Lindley. Book review: Multiple Comparisons: Theory and Methods, by J. C. Hsu. Journal of the Royal Statistical Society. Series D (The Statistician), 46(4):572–573, ???? 1997. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2988617.

Thomas:1997:BRB

[869] Ray Thomas. Book review: Interpreting Official Statistics, by R. Levitas; W. Guy. Journal of the Royal Statistical Society. Series D (The Statistician), 46(4):573–576, ???? 1997. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2988618.

Haigh:1997:BRB

[870] John Haigh. Book review: Discrete Gambling and Stochastic Games, by A. P. Maitra; W. D. Sudderth. Journal of the Royal Statistical Society. Series D (The Statistician), 46 (4):576–577, ???? 1997. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2988619.

Anonymous:1997:VI

[871] Anonymous. Volume information. Journal of the Royal Statistical Society. Series D (The Statistician), 46 (4):579–582, ???? 1997. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2988592.

Anonymous:1997:BMd

[872] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 46(4):??, ???? 1997. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2988620.

Anonymous:1997:FMd

[873] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 46(4):??, ???? 1997. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2988593.

Anonymous:1998:E

[874] Anonymous. Editorial. Journal of the Royal Statistical Society. Series D (The Statistician), 47(1):1–2, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2988423.

Kadane:1998:EE

[875] Joseph B. Kadane and Lara J. Wolfson. Experiences in elicitation. Journal of the Royal Statistical Society. Series D (The Statistician), 47(1):3–19, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988424.

OHagan:1998:EEB

[876] A. O’Hagan. Eliciting expert beliefs in substantial practical applications. Journal of the Royal Statistical Society. Series D (The Statistician), 47(1):21–35, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2988425.

Craig:1998:CPP

[877] Peter S. Craig, Michael Goldstein, Allan H. Seheult, and James A. Smith. Constructing partial prior specifications for models of complex physical systems. Journal of the Royal Statistical Society. Series D (The Statistician), 47(1):37–53, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2988426.

Anonymous:1998:DPE

[878] Anonymous. Discussion on the papers on ‘elicitation’. Journal of the Royal Statistical Society. Series D (The Statistician), 47(1):55–68, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2988427.

Brooks:1998:MCM

[879] Stephen P. Brooks. Markov chain Monte Carlo method and its application. Journal of the Royal Statistical Society. Series D (The Statistician), 47(1):69–100, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2988428.

Bernardo:1998:IBR

[880] José M. Bernardo and José M. Ramón. An introduction to Bayesian reference analysis: Inference on the ratio of multinomial parameters. Journal of the Royal Statistical Society. Series D (The Statistician), 47(1):101–135, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988429.

Farebrother:1998:NLC

[881] Richard William Farebrother. Nonlinear curve fitting and the true method of least squares. Journal of the Royal Statistical Society. Series D (The Statistician), 47(1):137–147, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988430.

Rahme:1998:EPR

[882] Elham Rahme and Lawrence Joseph. Estimating the prevalence of a rare disease: Adjusted maximum likelihood. Journal of the Royal Statistical Society. Series D (The Statistician), 47 (1):149–158, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2988431.

Congdon:1998:MMI

[883] Peter Congdon. A multilevel model for infant health outcomes: Maternal risk factors and geographic variation. Journal of the Royal Statistical Society. Series D (The Statistician), 47(1):159–182, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2988432.

Joanes:1998:CMS

[884] D. N. Joanes and C. A. Gill. Comparing measures of sample skewness and kurtosis. Journal of the Royal Statistical Society. Series D (The Statistician), 47(1):183–189, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2988433.

Sutradhar:1998:BTM

[885] Brajendra C. Sutradhar and Estela Bee Dagum. Bartlett-type modified test for moving seasonality with applications. Journal of the Royal Statistical Society. Series D (The Statistician), 47(1):191–206, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2988434.

Loughin:1998:TER

[886] Thomas M. Loughin. On Type I error rates in the analysis of unreplicated 2^{3}-factorials. Journal of the Royal Statistical Society. Series D (The Statistician), 47(1):207–210, ???? 1998. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2988435.

Cox:1998:SRS

[887] D. R. Cox. Some remarks on statistical education. Journal of the Royal Statistical Society. Series D (The Statistician), 47(1):211–213, ???? 1998. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2988436.

Halley:1998:BRB

[888] John M. Halley. Book review: Probability and Information, by D. Applebaum. Journal of the Royal Statistical Society. Series D (The Statistician), 47(1):215–216, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2988437.

Sapatinas:1998:BRBa

[889] Theofanis Sapatinas. Book review: Unimodality of Probability Measures, by E. M. J. Bertin; I. Cuculescu; R. Theodorescu. Journal of the Royal Statistical Society. Series D (The Statistician), 47(1):216–217, ???? 1998. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2988438.

Ridges:1998:BRB

[890] Peter Ridges. Book review: A Course in Stochastic Processes, by D. Bosq; H. T. Nguyen. Journal of the Royal Statistical Society. Series D (The Statistician), 47(1):217–218, ???? 1998. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2988439.

Argyle:1998:BRBa

[891] Jennifer Argyle. Book review: A Conceptual Introduction to Modeling: Qualitative and Quantitative Perspectives, by D. W. Britt. Journal of the Royal Statistical Society. Series D (The Statistician), 47(1):218, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2988440.

Castagliola:1998:BRBa

[892] Philippe Castagliola. Book review: General Statistics, by W. Chase; F. Bown. Journal of the Royal Statistical Society. Series D (The Statistician), 47(1):218–219, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2988441.

Kimber:1998:BRBa

[893] Alan Kimber. Book review: The Pocket Statistician: A Practical Guide to Quality Improvement, by S. Coleman; T. Greenfield; R. Jones; C. Morris; I. Puzey. Journal of the Royal Statistical Society. Series D (The Statistician), 47(1):219, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2988442.

Gordon:1998:BRBa

[894] A. D. Gordon. Book review: Tree Models of Similarity and Association, by J. E. Corter. Journal of the Royal Statistical Society. Series D (The Statistician), 47(1):220, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2988443.

Bartels:1998:BRBa

[895] Robert Bartels. Book review: Principles of Statistical Data Handling, by F. Davidson. Journal of the Royal Statistical Society. Series D (The Statistician), 47(1):220–221, ???? 1998. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2988444.

White:1998:BRB

[896] Allan White. Book review: Discovering Advanced Mathematics: Statistics, by R. Fentem. Journal of the Royal Statistical Society. Series D (The Statistician), 47(1):221–222, ???? 1998. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2988445.

Gani:1998:BRBa

[897] Joe Gani. Book review: An Intermediate Course in Probability, by A. Gut. Journal of the Royal Statistical Society. Series D (The Statistician), 47(1):222, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2988446.

LeVey:1998:BRB

[898] Georges Le Vey. Book review: Robust Statistical Procedures: Asymptotics and Interrelations, by J. Jureckova; P. K. Sen. Journal of the Royal Statistical Society. Series D (The Statistician), 47(1):222–224, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2988447.

French:1998:BRB

[899] Simon French. Book review: Operational Subjective Statistical Methods: A Mathematical, Philosophical and Historical Introduction, by F. Lad. Journal of the Royal Statistical Society. Series D (The Statistician), 47(1):224, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988448.

Bartels:1998:BRBb

[900] Robert Bartels. Book review: Handbook of Matrices, by H. Lütkepohl. Journal of the Royal Statistical Society. Series D (The Statistician), 47(1): 225, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988449.

Janacek:1998:BRB

[901] G. Janacek. Book review: The Biostatistics Cookbook: The Most User Friendly Guide for the Bio/ Medical Scientist, by S. Michelson; T. Schofield. Journal of the Royal Statistical Society. Series D (The Statistician), 47 (1):225–226, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2988450.

Ridout:1998:BRB

[902] Martin Ridout. Book review: Beyond ANOVA: Basics of Applied Statistics, by R. G. Miller. Journal of the Royal Statistical Society. Series D (The Statistician), 47(1):226–227, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988451.

Northrop:1998:BRB

[903] Paul Northrop. Book review: Optimization of Stochastic Models: The Interface Between Simulation and Optimization, by G. Ch. Pflug. Journal of the Royal Statistical Society. Series D (The Statistician), 47(1):227, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988452.

Dunn:1998:BRB

[904] Graham Dunn. Book review: Applied Factor Analysis in the Natural Sciences, by R. Reyment; K. G. Jöreskog.

Journal of the Royal Statistical Society. Series D (The Statistician), 47 (1):227–228, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2988453.

Young:1998:BRB

[905] Philip Young. Book review: Matrix Analysis for Statistics, by J. R. Schott. Journal of the Royal Statistical Society. Series D (The Statistician), 47 (1):228–229, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2988454.

Kiesel:1998:BRBa

[906] Rüdiger Kiesel. Book review: Probability, by A. N. Shiryaev. Journal of the Royal Statistical Society. Series D (The Statistician), 47(1):229– 230, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988455.

Dialsingh:1998:BRBa

[907] Isaac Dialsingh. Book review: Elements of Survey Sampling, by R. Singh; N. S. Mangat. Journal of the Royal Statistical Society. Series D (The Statistician), 47(1):230–231, ???? 1998. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2988456.

Sapatinas:1998:BRBb

[908] Theofanis Sapatinas. Book review: Counterexamples in Probability, by J. Stoyanov. Journal of the Royal Statistical Society. Series D (The Statistician), 47(1):231, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2988457.

Boshnakov:1998:BRBa

[909] Georgi Boshnakov. Book review: Applications of Time Series Analysis in Astronomy and Meteorology, by T. Subba Rao; M. B. Priestley; O. Lessi. Journal of the Royal Statistical Society. Series D (The Statistician), 47 (1):231–233, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2988458.

Kiesel:1998:BRBb

[910] Rüdiger Kiesel. Book review: Weak Convergence and Empirical Processes with Applications to Statistics, by A. W. van der Vaart; J. A. Wellner. Journal of the Royal Statistical Society. Series D (The Statistician), 47(1):233, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2988459.

Venables:1998:BRB

[911] W. N. Venables. Book review: Aspects of Statistical Inference, by A. H. Welsh. Journal of the Royal Statistical Society. Series D (The Statistician), 47(1):233–234, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2988460.

Huhtala:1998:BRB

[912] Kyösti Huhtala. Book review: Understanding Statistics: An Introduction for the Social Sciences, by D. B.

Wright. Journal of the Royal Statistical Society. Series D (The Statistician), 47(1):234–235, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2988461.

Coleman:1998:BRB

[913] Shirley Coleman. Book review: Elements of Statistics. Journal of the Royal Statistical Society. Series D (The Statistician), 47(1):237–238, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988462.

Anonymous:1998:BMa

[914] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 47(1):??, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2988463.

Anonymous:1998:FMa

[915] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 47(1):??, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2988422.

Sprent:1998:SMT

[916] P. Sprent. Statistics and mathematicstrouble at the interface? Journal of the Royal Statistical Society. Series D (The Statistician), 47(2):239– 244, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988664.

Hand:1998:BMS

[917] David J. Hand. Breaking misconceptionsstatistics and its relationship to mathematics. Journal of the Royal Statistical Society. Series D (The Statistician), 47(2):245–250, ???? 1998. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2988665.

Senn:1998:MGH

[918] Stephen Senn. Mathematics: Governess or handmaiden? Journal of the Royal Statistical Society. Series D (The Statistician), 47(2):251–259, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2988666.

Bailey:1998:SMA

[919] R. A. Bailey. Statistics and mathematics: The appropriate use of mathematics within statistics. Journal of the Royal Statistical Society. Series D (The Statistician), 47(2):261–271, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2988667.

Anonymous:1998:DPS

[920] Anonymous. Discussion on the papers on ‘statistics and mathematics’. Journal of the Royal Statistical Society. Series D (The Statistician), 47 (2):273–290, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2988668.

Boyle:1998:OSP

[921] Conall Boyle. Organizations selecting people: How the process could be made fairer by the appropriate use of lotteries. Journal of the Royal Statistical Society. Series D (The Statistician), 47(2):291–321, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2988669.

Nadal:1998:ABP

[922] Neidha Nadal and Luis Raúl Pericchi. An automatic Bayesian procedure for likelihoods with shifted origin. Journal of the Royal Statistical Society. Series D (The Statistician), 47(2):323–332, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2988670.

Raghunathan:1998:CEI

[923] Trivellore E. Raghunathan and David S. Siscovick. Combining exposure information from various sources in an analysis of a case-control study. Journal of the Royal Statistical Society. Series D (The Statistician), 47(2):333– 347, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988671.

Bowman:1998:CBL

[924] Adrian Bowman, W. Harper Gilmour, Gillian Constable, Neville Davies, Steven G. Gilmour, and Edwin J. Redfern. Computer-based learning in statistics: A problem solving approach. Journal of the Royal Statistical Society. Series D (The Statistician), 47 (2):349–364, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2988672.

Saei:1998:LTM

[925] A. Saei and C. A. McGilchrist. Longitudinal threshold models with random components. Journal of the Royal Statistical Society. Series D (The Statistician), 47(2):365–375, ???? 1998. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2988673.

Bertrand:1998:CEE

[926] Philip V. Bertrand. Constructing explained and explanatory variables with strange statistical analysis results. Journal of the Royal Statistical Society. Series D (The Statistician), 47 (2):377–383, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2988674.

Zumbo:1998:NMC

[927] Bruno D. Zumbo and Anita M. Hubley. A note on misconceptions concerning prospective and retrospective power. Journal of the Royal Statistical Society. Series D (The Statistician), 47(2):385–388, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2988675.

Sukhatme:1998:OPV

[928] Suhas P. Sukhatme. Obituary: Pandurang Vasudeo Sukhatme. Journal of the Royal Statistical Society. Series D (The Statistician), 47(2):389– 390, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988676.

Hinde:1998:BRB

[929] John Hinde. Book review: Introduction to the Statistical Analysis of Categorical Data, by E. B. Andersen. Journal of the Royal Statistical Society. Series D (The Statistician), 47(2):391, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2988677.

Stephens:1998:BRBa

[930] D. A. Stephens. Book review: Probability and Information: An Integrated Approach, by D. Applebaum. Journal of the Royal Statistical Society. Series D (The Statistician), 47(2):392, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988678.

Scott:1998:BRBa

[931] E. Marian Scott. Book review: Statistics for the Environment 3: Pollution Assessment and Control, by V. Barnett; K. F. Turkman. Journal of the Royal Statistical Society. Series D (The Statistician), 47(2):392– 393, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988679.

Stephens:1998:BRBb

[932] D. A. Stephens. Book review: Bayesian Statistics 5, by J. M. Bernardo; J. O. Berger; A. P. Dawid; A. F. M. Smith. Journal of the Royal Statistical Society. Series D (The Statistician), 47(2):393–394, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2988680.

Currall:1998:BRB

[933] James Currall. Book review: Statistical Significance: Rationale, Validity and Utility, by S. L. Chow. Journal of the Royal Statistical Society. Series D (The Statistician), 47 (2):394–395, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2988681.

McCollin:1998:BRB

[934] Chris McCollin. Book review: Statistical Methods for Industrial Process Control, by D. Drain. Journal of the Royal Statistical Society. Series D (The Statistician), 47(2):395, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2988682.

Prys-Williams:1998:BRB

[935] Allan G. Prys-Williams. Book review: A Guide to Chi-Squared Testing, by P. E. Greenwood; M. S. Nikulin. Journal of the Royal Statistical Society. Series D (The Statistician), 47(2):396, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2988683.

Collop:1998:BRB

[936] M. J. Collop. Book review: Elements of Pattern Theory, by U. Grenander. Journal of the Royal Statistical Society. Series D (The Statistician), 47 (2):396–397, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2988684.

Brewer:1998:BRB

[937] Mark Brewer. Book review: An Introduction to Bayesian Networks, by F. V. Jensen. Journal of the Royal Statistical Society. Series D (The Statistician), 47(2):397–398, ???? 1998. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2988685.

Militino:1998:BRB

[938] Ana F. Militino. Book review: Statistics, Principles and Methods, by R. A. Johnson; G. K. Bhattacharyya. Journal of the Royal Statistical Society. Series D (The Statistician), 47 (2):398–399, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2988686.

Coolen:1998:BRBa

[939] Frank Coolen. Book review: Business Statistics: Decision Making with Data, by R. A. Johnson; D. W. Wichern. Journal of the Royal Statistical Society. Series D (The Statistician), 47(2):399, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2988687.

Bell:1998:BRBa

[940] John F. Bell. Book review: Statistical Methods for Plant Variety Evaluation, by R. A. Kempton; P. N. Fox. Journal of the Royal Statistical Society. Series D (The Statistician), 47 (2):399–400, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2988688.

Kemp:1998:BRB

[941] A. W. Kemp. Book review: Markov Processes for Stochastic Modelling, by M. Kijima. Journal of the Royal Statistical Society. Series D (The Statistician), 47(2):400–401, ???? 1998. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2988689.

Machado:1998:BRBa

[942] Orlando Machado. Book review: Probability: An Introduction with Statistical Applications, by J. J. Kinney. Journal of the Royal Statistical Society. Series D (The Statistician), 47(2):401, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2988690.

Gani:1998:BRBb

[943] Joe Gani. Book review: Probability: A Survey of the Mathematical Theory, by J. W. Lamperti. Journal of the Royal Statistical Society. Series D (The Statistician), 47(2):401– 402, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2988691.

Mayo:1998:BRB

[944] O. Mayo. Book review: Error and the Growth of Experimental Knowledge, by D. G. Mayo. Journal of the Royal Statistical Society. Series D (The Statistician), 47(2):402–404, ???? 1998. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2988692.

Craig:1998:BRB

[945] Peter Craig. Book review: Statistics of the Boolean Model for Practitioners and Mathematicians, by I. Molchanov. Journal of the Royal Statistical Society. Series D (The Statistician), 47(2):404, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2988693.

Haigh:1998:BRB

[946] John Haigh. Book review: Markov Chains, by J. R. Norris. Journal of the Royal Statistical Society. Series D (The Statistician), 47(2):404– 405, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988694.

Holden:1998:BRB

[947] Ken Holden. Book review: A Beginner’s Guide to Structural Equation Modelling, by R. E. Schumacker; R. G. Lomax. Journal of the Royal Statistical Society. Series D (The Statistician), 47(2):405–406, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2988695.

Molchanov:1998:BRB

[948] Ilya Molchanov. Book review: The Statistical Theory of Shape, by C. G. Small. Journal of the Royal Statistical Society. Series D (The Statistician), 47(2):406, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2988696.

Metcalfe:1998:BRB

[949] A. V. Metcalfe. Book review: An Introduction to Statistics for Managers, by H. A. Spooner; C. D. Lewis. Journal of the Royal Statistical Society. Series D (The Statistician), 47(2):407, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988697.

Morris:1998:BRB

[950] Clare Morris. Book review: Multivariate Analysis Techniques in Social Science Research, by J. Tacq. Journal of the Royal Statistical Society. Series D (The Statistician), 47 (2):407–408, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2988698.

Glendinning:1998:BRB

[951] R. H. Glendinning. Book review: Nonlinear Filters, Estimation and Applications, by H. Tanizaki. Journal of the Royal Statistical Society. Series D (The Statistician), 47(2):409, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2988699.

Morgan:1998:BRB

[952] Geoffrey Morgan. Book review: Computational Economics and Finance: Modelling and Analysis with Mathematica, by H. R. Varian. Journal of the Royal Statistical Society. Series D (The Statistician), 47(2):409– 410, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988700.

Borja:1998:BRBa

[953] Mario Cortina Borja. Book review: Linear and Nonlinear Models for the Analysis of Repeated Measurements, by E. F. Vonesh; V. M. Chinchilli. Journal of the Royal Statistical Society. Series D (The Statistician), 47 (2):410–411, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2988701.

Shoesmith:1998:BRBa

[954] Eddie Shoesmith. Book review: Handbook of Demonstrations and Activities in the Teaching of Psychology: Volume 1, Introductory, Statistics, Research Methods, and History, by M. E. Ware; D. E. Johnson. Journal of the Royal Statistical Society. Series D (The Statistician), 47(2):411, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2988702.

Anonymous:1998:BMb

[955] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 47(2):412, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988703.

Anonymous:1998:FMb

[956] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 47(2):??, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2988663.

Unwin:1998:F

[957] David Unwin and Antony Unwin. Foreword. Journal of the Royal Statistical Society. Series D (The Statistician), 47(3):413, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2988622.

Unwin:1998:ESD

[958] A. Unwin and D. Unwin. Exploratory spatial data analysis with local statistics. Journal of the Royal Statistical Society. Series D (The Statistician), 47(3):415–421, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2988623.

Wilhelm:1998:ESD

[959] Adalbert Wilhelm and Robert Steck. Exploring spatial data by using interactive graphics and local statistics. Journal of the Royal Statistical Society. Series D (The Statistician), 47 (3):423–430, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2988624.

Brunsdon:1998:GWR

[960] Chris Brunsdon, Stewart Fotheringham, and Martin Charlton. Geographically weighted regression-modelling spatial non-stationarity. Journal of the Royal Statistical Society. Series D (The Statistician), 47(3):431–443, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988625.

Wilhelm:1998:ISA

[961] Adalbert Wilhelm and Meike Sander. Interactive statistical analysis of dialect features. Journal of the Royal Statistical Society. Series D (The Statistician), 47(3):445–455, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2988626.

Haining:1998:ESD

[962] Robert Haining, Stephen Wise, and Jingsheng Ma. Exploratory spatial data analysis in a geographic information system environment. Journal of the Royal Statistical Society. Series D (The Statistician), 47(3):457– 469, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988627.

Brunsdon:1998:ESD

[963] Chris Brunsdon. Exploratory spatial data analysis and local indicators of spatial association with XLISP–STAT. Journal of the Royal Statistical Society. Series D (The Statistician), 47 (3):471–484, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2988628.

Dykes:1998:CVE

[964] Jason Dykes. Cartographic visualization: Exploratory spatial data analysis with local indicators of spatial association using Tcl/Tk and cdv. Journal of the Royal Statistical Society. Series D (The Statistician), 47 (3):485–497, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2988629.

Bivand:1998:SSD

[965] Roger S. Bivand. Software and software design issues in the exploration of local dependence. Journal of the Royal Statistical Society. Series D (The Statistician), 47(3):499–508, ???? 1998. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2988630.

Grubb:1998:MCA

[966] H. J. Grubb. Models for comparing athletic performances. Journal of the Royal Statistical Society. Series D (The Statistician), 47(3):509– 521, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2988631.

Dixon:1998:BPM

[967] Mark J. Dixon and Michael E. Robinson. A birth process model for Association Football matches. Journal of the Royal Statistical Society. Series D (The Statistician), 47(3):523– 538, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2988632.

Edwards:1998:BRB

[968] A. W. F. Edwards. Book review: Advances in Biometry: 50 Years of the International Biometric Society, by P. Armitage; H. A. David. Journal of the Royal Statistical Society. Series D (The Statistician), 47(3):539, ???? 1998.

CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2988633.

Denison:1998:BRB

[969] David G. T. Denison. Book review: Applied Smoothing Techniques for Data Analysis — The Kernel Approach with S-Plus Illustrations, by A. W. Bowman; A. Azzalini. Journal of the Royal Statistical Society. Series D (The Statistician), 47(3):539–540, ???? 1998. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2988634.

Shabbir:1998:BRB

[970] Javid Shabbir. Book review: Applied Statistical Methods: For Business, Economics and the Social Sciences, by W. L. Carlson; B. Thorne. Journal of the Royal Statistical Society. Series D (The Statistician), 47(3):540, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2988635.

Brooks:1998:BRBa

[971] Stephen Brooks. Book review: Bootstrap Methods and Their Applications, by A. C. Davidson; D. V. Hinkley. Journal of the Royal Statistical Society. Series D (The Statistician), 47(3):541, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2988636.

Zumbo:1998:BRB

[972] Bruno D. Zumbo. Book review: Design and Analysis of Single-Case Research, by R. D. Franklin; D. B. Allison; B.

S. Gorman. Journal of the Royal Statistical Society. Series D (The Statistician), 47(3):541–543, ???? 1998. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2988637.

Knott:1998:BRB

[973] Martin Knott. Book review: Bonferroni-Type Inequalities with Applications, by J. Galambos; I. Simonelli. Journal of the Royal Statistical Society. Series D (The Statistician), 47(3):543– 544, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988638.

Dialsingh:1998:BRBb

[974] Isaac Dialsingh. Book review: The SPSS Guide to the New Statistical Analysis of Data, by S. Gerber; K. E. Vogel. Journal of the Royal Statistical Society. Series D (The Statistician), 47(3):544, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2988639.

Wooff:1998:BRB

[975] David Wooff. Book review: Statistical Tools for Nonlinear Regression: A Practical Guide with S-PLUS Examples, by S. Huet; A. Bouvier; M.- A. Gruet; E. Jolivet. Journal of the Royal Statistical Society. Series D (The Statistician), 47(3):545, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2988640.

Kumar:1998:BRB

[976] Kuldeep Kumar. Book review: Asymptotic Theory of Nonlinear Regression, by A. V. Ivanov. Journal of the Royal Statistical Society. Series D (The Statistician), 47(3):545–546, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988641.

Tobias:1998:BRB

[977] Aurelio Tobías. Book review: Statistical Graphics for Univariate and Bivariate Data, by W. G. Jacoby. Journal of the Royal Statistical Society. Series D (The Statistician), 47 (3):546–547, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2988642.

Nadarajah:1998:BRBa

[978] S. Nadarajah. Book review: Multivariate Models and Dependence Concepts, by H. Joe. Journal of the Royal Statistical Society. Series D (The Statistician), 47(3):547, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2988643.

Dialsingh:1998:BRBc

[979] Isaac Dialsingh. Book review: Response Surfaces: Designs and Analyses, by A. I. Khuri; J. A. Cornell. Journal of the Royal Statistical Society. Series D (The Statistician), 47 (3):547–548, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2988644.

Gani:1998:BRBc

[980] Joe Gani. Book review: Markov Processes for Stochastic Modeling, by M. Kijima. Journal of the Royal Statistical Society. Series D (The Statistician), 47(3):548–549, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2988645.

Gasparini:1998:BRB

[981] Mauro Gasparini. Book review: Survival Analysis: Techniques for Censored and Truncated Data, by J. P. Klein; M. L. Moeschberger. Journal of the Royal Statistical Society. Series D (The Statistician), 47(3):549, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988646.

Brooks:1998:BRBb

[982] Stephen Brooks. Book review: Breakthroughs in Statistics, Volume III, by S. Kotz; N. L. Johnson. Journal of the Royal Statistical Society. Series D (The Statistician), 47(3):549– 550, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988647.

Machado:1998:BRBb

[983] Orlando Machado. Book review: The Basics of S and S-PLUS, by A. Krause; M. Olson. Journal of the Royal Statistical Society. Series D (The Statistician), 47(3):550–551, ???? 1998. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2988648.

Scott:1998:BRBb

[984] Marian Scott. Book review: Applying Generalised Linear Models, by J.

K. Lindsey. Journal of the Royal Statistical Society. Series D (The Statistician), 47(3):551–552, ???? 1998. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2988649.

Bashir:1998:BRB

[985] Saghir A. Bashir. Book review: Parametric Statistical Inference, by J. K. Lindsey. Journal of the Royal Statistical Society. Series D (The Statistician), 47(3):552–553, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2988650.

Rees:1998:BRB

[986] Hedley Rees. Book review: Regression Models for Categorical and Limited Dependent Variables, by J. S. Long. Journal of the Royal Statistical Society. Series D (The Statistician), 47(3):553–554, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2988651.

Quinn:1998:BRB

[987] Barry Quinn. Book review: Hidden Markov and Other Models for Discrete-Valued Time Series, by L. MacDonald; W. Zucchini. Journal of the Royal Statistical Society. Series D (The Statistician), 47(3):554, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2988652.

Thompson:1998:BRB

[988] John Thompson. Book review: The EM Algorithm and Extensions, by G.

J. McLachlan; T. Krishnan. Journal of the Royal Statistical Society. Series D (The Statistician), 47(3):554– 555, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988653.

Nadarajah:1998:BRBb

[989] S. Nadarajah. Book review: Statistical Analysis of Extreme Values with Applications to Insurance, Finance, Hydrology and Other Fields, by R.-D. Reiss; M. Thomas. Journal of the Royal Statistical Society. Series D (The Statistician), 47(3):555–556, ???? 1998. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2988654.

Rodrigues:1998:BRB

[990] Paulo Rodrigues. Book review: Modern Regression Methods, by T. R. Ryan. Journal of the Royal Statistical Society. Series D (The Statistician), 47 (3):556–557, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2988655.

West:1998:BRB

[991] Robert West. Book review: Statistical Methods — A Geometric Primer, by D. J. Saville; G. R. Wood. Journal of the Royal Statistical Society. Series D (The Statistician), 47(3):557– 558, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988656.

Bowman:1998:BRB

[992] Adrian Bowman. Book review: Smoothing Methods in Statistics, by J.

S. Simonoff. Journal of the Royal Statistical Society. Series D (The Statistician), 47(3):558, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2988657.

Rao:1998:BRB

[993] T. Subba Rao. Book review: Computational Economics and Finance-Modelling and Analysis with Mathematica, by H. R. Varian. Journal of the Royal Statistical Society. Series D (The Statistician), 47(3):558– 559, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988658.

Pettit:1998:BRB

[994] Lawrence Pettit. Book review: Bayesian Forecasting and Dynamic Models, by M. West; J. Harrison. Journal of the Royal Statistical Society. Series D (The Statistician), 47(3):559– 560, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988659.

Kolev:1998:BRBa

[995] Nikolai Kolev. Book review: Econometric Analysis of Count Data, by R. Winkelmann. Journal of the Royal Statistical Society. Series D (The Statistician), 47(3):560–561, ???? 1998. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2988660.

Minkova:1998:BRB

[996] Leda Minkova. Book review: Some Aspects of Brownian Motion, Part II, Some Recent Martingale Problems, by M. Yor. Journal of the Royal Statistical Society. Series D (The Statistician), 47(3):561–562, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2988661.

Anonymous:1998:BMc

[997] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 47(3):??, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2988662.

Anonymous:1998:FMc

[998] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 47(3):??, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2988621.

Anonymous:1998:BML

[999] Anonymous. The Beveridge Memorial Lectures: Foreword. Journal of the Royal Statistical Society. Series D (The Statistician), 47(4):563, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2988359.

Field:1998:BML

[1000] Frank Field. The Beveridge Memorial Lecture, 1998. Journal of the Royal Statistical Society. Series D (The Statistician), 47(4):565–571, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2988360.

Barnes:1998:CTT

[1001] Paul Barnes. Can takeover targets be identified by statistical techniques?: Some UK evidence. Journal of the Royal Statistical Society. Series D (The Statistician), 47(4):573– 591, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988361.

Billah:1998:CBP

[1002] Md. Baki Billah and A. K. Md. E.

Saleh. Conflict between pretest estimators induced by three large sample tests under a regression model with Student t-error. Journal of the Royal Statistical Society. Series D (The Statistician), 47(4):593–606, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2988362.

Reeves:1998:BRM

[1003] John Reeves and Gilbert MacKenzie. A bivariate regression model with serial correlation. Journal of the Royal Statistical Society. Series D (The Statistician), 47(4):607–615, ???? 1998. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2988363.

Barnett:1998:FFD

[1004] Vic Barnett and John Haworth. A fractional factorial design for benchmark testing of a Bayesian method for multilocation audits. Journal of the Royal Statistical Society. Series D (The Statistician), 47(4):617–628, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988364.

Cox:1998:UME

[1005] Simon J. Cox, Geoffrey J. Daniell, and Denis A. Nicole. Using maximum entropy to double One’s expected winnings in the UK national lottery. Journal of the Royal Statistical Society. Series D (The Statistician), 47 (4):629–641, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2988365.

Champion:1998:DDW

[1006] R. Champion, C. T. Lenard, and T. M. Mills. Demonstrating the Durbin–Watson statistic. Journal of the Royal Statistical Society. Series D (The Statistician), 47(4):643–644, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988366.

Holt:1998:NSM

[1007] Tim Holt. National statistics — meeting new challenges. Journal of the Royal Statistical Society. Series D (The Statistician), 47(4):645–646, 1998. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2988367.

Anonymous:1998:RCS

[1008] Anonymous. Report of the Council for the session 1997–1998. Journal of the Royal Statistical Society. Series D (The Statistician), 47(4):647– 699, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988368.

Mardia:1998:OGS

[1009] K. V. Mardia. Obituary: Geoffrey Stuart Watson. Journal of the Royal Statistical Society. Series D (The Statistician), 47(4):701–702, ???? 1998. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2988369.

Assuncao:1998:BRB

[1010] Renato Assuncao. Book review: Methods for Investigating Localized Clustering of Disease, by F. E. Alexander; P. Boyle. Journal of the Royal Statistical Society. Series D (The Statistician), 47(4):703–704, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2988370.

Dialsingh:1998:BRBd

[1011] Isaac Dialsingh. Book review: The Statistical Approach to Social Measurement, by D. J. Bartholomew. Journal of the Royal Statistical Society. Series D (The Statistician), 47(4):704, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988371.

Saez:1998:BRB

[1012] Marc Saez. Book review: Modern Multidimensional Scaling: Theory and Applications, by I. Borg; P. Groenen. Journal of the Royal Statistical Society. Series D (The Statistician), 47(4):705, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2988372.

Scott:1998:BRBc

[1013] Marian Scott. Book review: Log-Linear Models and Logistic Regression, by R. Christensen. Journal of the Royal Statistical Society. Series D (The Statistician), 47(4):705–706, ???? 1998. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2988373.

Wilhelm:1998:BRB

[1014] Adalbert Wilhelm. Book review: Plane Answers to Complex Questions: The Theory of Linear Models, by R. Christensen. Journal of the Royal Statistical Society. Series D (The Statistician), 47(4):706–708, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2988374.

Kimber:1998:BRBb

[1015] Alan Kimber. Book review: Statistical Case Studies for Industrial Process Improvement, by V. Czitrom; P. D. Spagon. Journal of the Royal Statistical Society. Series D (The Statistician), 47(4):708, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2988375.

Kolev:1998:BRBb

[1016] Nikolai Kolev. Book review: Random Iterative Models, by M. Duflo. Journal of the Royal Statistical Society. Series D (The Statistician), 47 (4):708–709, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2988376.

Scotto:1998:BRB

[1017] M. G. Scotto. Book review: Modelling Extremal Events for Insurance and Finance, by P. Embrechts; C. Klüppelberg; T. Mikosch. Journal of the Royal Statistical Society. Series D (The Statistician), 47(4):709– 710, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988377.

Gordon:1998:BRBb

[1018] A. D. Gordon. Book review: The Analysis of Proximity Data, by B. S. Everitt; S. Rabe-Hesketh. Journal of the Royal Statistical Society. Series D (The Statistician), 47(4):710– 711, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988378.

Bell:1998:BRBb

[1019] John F. Bell. Book review: Social Statistics for a Diverse Society, by C. Franfort-Nachmias. Journal of the Royal Statistical Society. Series D (The Statistician), 47(4):711, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2988379.

Bartkowiak:1998:BRB

[1020] Anna Bartkowiak. Book review: The Practice of Bayesian Analysis, by S. French; J. Q. Smith. Journal of the Royal Statistical Society. Series D (The Statistician), 47(4):712–713, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988380.

Fan:1998:BRB

[1021] Yanan Fan. Book review: Markov Chain Monte Carlo-Stochastic Simulation for Bayesian Inference, by D. Gamerman. Journal of the Royal Statistical Society. Series D (The Statistician), 47(4):713, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2988381.

Coolen:1998:BRBb

[1022] Frank Coolen. Book review: Bayesian Methods for Finite Population Sampling, by M. Ghosh; G. Meeden. Journal of the Royal Statistical Society. Series D (The Statistician), 47 (4):713–714, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2988382.

Castagliola:1998:BRBb

[1023] Philippe Castagliola. Book review: Sequential Estimation, by M. Ghosh; N. Mukhopadhyay; P. K. Sen. Journal of the Royal Statistical Society. Series D (The Statistician), 47(4):714– 715, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2988383.

Reed:1998:BRB

[1024] R. J. Reed. Book review: The Beginner’s Guide to Mathematica Version 3, by J. Glynn; T. W. Gray. Journal of the Royal Statistical Society. Series D (The Statistician), 47(4):715, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2988384.

Boshnakov:1998:BRBb

[1025] Georgi Boshnakov. Book review:

ARCH Models and Financial Applications, by C. Gourieroux. Journal of the Royal Statistical Society. Series D (The Statistician), 47(4):715– 716, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988385.

Penenberg:1998:BRB

[1026] Darryl Neil Penenberg. Book review:

Applied Statistics: A First Course in Inference, by F. Graybill; H. Iyer; R. Burdick. Journal of the Royal Statistical Society. Series D (The Statistician), 47(4):716–717, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2988386.

Wright:1998:BRB

[1027] Lesley F. Wright. Book review: Understanding Data: Principles and Practice of Statistics, by D. Griffiths; W. D. Stirling; K. L. Weldon. Journal of the Royal Statistical Society. Series D (The Statistician), 47(4):717– 718, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988387.

Gray:1998:BRB

[1028] Alison Gray. Book review: Construction and Assessment of Classification Rules, by D. J. Hand. Journal of the Royal Statistical Society. Series D (The Statistician), 47(4):718– 720, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2988388.

Neykov:1998:BRB

[1029] Neyko M. Neykov. Book review: Regional Frequency Analysis: An Approach Based on L-Moments, by J. R. M. Hosking; J. R. Wallis. Journal of the Royal Statistical Society. Series D (The Statistician), 47(4):720–721, 1998. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2988389.

Sapatinas:1998:BRBc

[1030] Theofanis Sapatinas. Book review: Advances in the Theory and Practice of Statistics, by N. L. Johnson; N. Balakrishnan. Journal of the Royal Statistical Society. Series D (The Statistician), 47(4):721–722, ???? 1998. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2988390.

Sprent:1998:BRB

[1031] P. Sprent. Book review: Leading Personalities in Statistical Sciences — From the Seventeenth Century to the Present, by N. L. Johnson; S. Kotz. Journal of the Royal Statistical Society. Series D (The Statistician), 47 (4):722–723, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2988391.

Stein:1998:BRB

[1032] A. Stein. Book review: A Solution to the Ecological Inference Problem, by G. King. Journal of the Royal Statistical Society. Series D (The Statistician), 47(4):723, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2988392.

Stoimenova:1998:BRB

[1033] Eugenia Stoimenova. Book review:

Testing Statistical Hypotheses, by E. L. Lehmann. Journal of the Royal Statistical Society. Series D (The Statistician), 47(4):723–724, ???? 1998. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2988393.

Borja:1998:BRBb

[1034] Mario Cortina Borja. Book review: Multivariate Statistical Methods: A First Course, by G. A. Marcoulides; S. L. Hershberger. Journal of the Royal Statistical Society. Series D (The Statistician), 47(4):724– 725, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988394.

Stone:1998:BRB

[1035] M. Stone. Book review: Actualité et Universalité de la Pensée Scientifique d’Adolphe Quetelet, by Mémoire de la Classe des Sciences. Journal of the Royal Statistical Society. Series D (The Statistician), 47(4):725– 726, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988395.

Lampard:1998:BRB

[1036] Richard Lampard. Book review: Constructing Classes: Towards a New Social Classification for the UK, by D. Rose; K. O’Reilly. Journal of the Royal Statistical Society. Series D (The Statistician), 47(4):726–727, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2988396.

Woodward:1998:BRB

[1037] P. W. Woodward. Book review: Statistical Issues in Drug Development, by S. Senn. Journal of the Royal Statistical Society. Series D (The Statistician), 47(4):727–728, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2988397.

Shoesmith:1998:BRBb

[1038] Eddie Shoesmith. Book review: The Art of Causal Conjecture, by G. Shafer. Journal of the Royal Statistical Society. Series D (The Statistician), 47 (4):728–729, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2988398.

Balding:1998:BRB

[1039] David Balding. Book review: Statistics in Human Genetics, by P. Sham. Journal of the Royal Statistical Society. Series D (The Statistician), 47 (4):729–730, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2988399.

Argyle:1998:BRBb

[1040] Jennifer Argyle. Book review: Modern Applied Statistics with S-PLUS, by W. N. Venables; B. D. Ripley. Journal of the Royal Statistical Society. Series D (The Statistician), 47(4):730, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988400.

Anonymous:1998:VI

[1041] Anonymous. Volume information.

Journal of the Royal Statistical Society. Series D (The Statistician), 47 (4):731–740, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2988357.

Anonymous:1998:BMd

[1042] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 47(4):??, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2988401.

Anonymous:1998:FMd

[1043] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 47(4):??, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2988358.

Lindsey:1999:SSH

[1044] J. K. Lindsey. Some statistical heresies. Journal of the Royal Statistical Society. Series D (The Statistician), 48(1):1–40, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2680893.

Brien:1999:TSF

[1045] C. J. Brien and R. W. Payne. Tiers, structure formulae and the analysis of complicated experiments. Journal of the Royal Statistical Society. Series D (The Statistician), 48(1):41–52, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2680894.

OQuigley:1999:LST

[1046] John O’Quigley, Philippe Flandre, and Ethan Reiner. Large sample theory for Schemper’s measures of explained variation in the Cox regression model. Journal of the Royal Statistical Society. Series D (The Statistician), 48(1):53–62, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2680895.

Forsberg:1999:TEA

[1047] Lars Forsberg, Bo Jonsson, and Ulf Kristiansson. On Type I errors after a preliminary test for heteroscedasticity. Journal of the Royal Statistical Society. Series D (The Statistician), 48(1):63–72, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2680896.

Fung:1999:ODS

[1048] Wing-Kam Fung. Outlier diagnostics in several multivariate samples. Journal of the Royal Statistical Society. Series D (The Statistician), 48 (1):73–84, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2680897.

Cramer:1999:PPB

[1049] J. S. Cramer. Predictive performance of the binary logit model in unbalanced samples. Journal of the Royal Statistical Society. Series D (The Statistician), 48(1):85–94, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2680898.

Lim:1999:SAI

[1050] Dong Hoon Lim, Robert A. Lordo, and Douglas A. Wolfe. A screening analysis for incomplete ranking data. Journal of the Royal Statistical Society. Series D (The Statistician), 48 (1):95–107, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2680899.

Buckland:1999:EMW

[1051] S. T. Buckland, D. C. Macmillan, E. I. Duff, and N. Hanley. Estimating mean willingness to pay from dichotomous choice contingent valuation studies. Journal of the Royal Statistical Society. Series D (The Statistician), 48(1):109–124, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2680900.

Dunne:1999:RBV

[1052] Adrian Dunne, John Devane, and Thomas O’Hara. The relationship between in vitro drug dissolution and in vivo absorption. Journal of the Royal Statistical Society. Series D (The Statistician), 48(1):125–133, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2680901.

Anonymous:1999:FMa

[1053] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 48(1):135, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2680892.

Militino:1999:BRBa

[1054] Ana F. Militino. Book review: Mathematical Tools for Applied Multivariate Analysis, by J. D. Carroll; P. E. Green. Journal of the Royal Statistical Society. Series D (The Statistician), 48(1):136, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2680902.

Sofroniou:1999:BRBa

[1055] Nick Sofroniou. Book review: SPSS: Analysis Without Anguish, by S. J. Coakes; L. G. Steed. Journal of the Royal Statistical Society. Series D (The Statistician), 48(1):136–137, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2680903.

Collop:1999:BRB

[1056] Michael Collop. Book review: An Electronic Companion to Statistics, by G. W. Cobb; J. D. Cryer. Journal of the Royal Statistical Society. Series D (The Statistician), 48(1):137– 138, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2680904.

Glendinning:1999:BRB

[1057] R. H. Glendinning. Book review: Limit Theorems in Change-Point Analysis, by M. Csörgő; L. Horváth. Journal of the Royal Statistical Society. Series D (The Statistician), 48(1):138, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2680905.

Kunst:1999:BRB

[1058] Robert M. Kunst. Book review:

Time Series, Unit Roots, and Cointegration, by P. Dhrymes. Journal of the Royal Statistical Society. Series D (The Statistician), 48(1):139– 140, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2680906.

Sprent:1999:BRBa

[1059] P. Sprent. Book review: Applied Regression Analysis, by N. R. Draper; H. Smith. Journal of the Royal Statistical Society. Series D (The Statistician), 48(1):140–141, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2680907.

Upton:1999:BRB

[1060] Graham Upton. Book review: Applied Regression Analysis, Linear Models, and Related Methods, by J. Fox. Journal of the Royal Statistical Society. Series D (The Statistician), 48(1): 141, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2680908.

Lark:1999:BRB

[1061] R. M. Lark. Book review: Geostatistics for Natural Resources Evaluation, by P. Goovaerts. Journal of the Royal Statistical Society. Series D (The Statistician), 48(1):142–143, ???? 1999. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2680909.

Kimber:1999:BRBa

[1062] Alan Kimber. Book review: Statistical Graphics for Visualizing Multivariate Data, by W. G. Jacoby. Journal of the Royal Statistical Society. Series D (The Statistician), 48 (1):143–144, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2680910.

Kiesel:1999:BRB

[1063] Rüdiger Kiesel. Book review: Foundations of Modern Probability: Probability and Its Applications, by O. Kallenberg. Journal of the Royal Statistical Society. Series D (The Statistician), 48(1):144, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2680911.

Apostolides:1999:BRB

[1064] Michael G. Apostolides. Book review: A Self-Teaching Guide: Statistics, by D. J. Koosis. Journal of the Royal Statistical Society. Series D (The Statistician), 48(1):144–145, ???? 1999. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2680912.

Gani:1999:BRBa

[1065] Joe Gani. Book review: Exponential Families of Stochastic Processes, by U. Kuchler; M. Sorensen. Journal of the Royal Statistical Society. Series D (The Statistician), 48(1):145– 146, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2680913.

Bashir:1999:BRB

[1066] Saghir A. Bashir. Book review: Mathematical and Statistical Methods for Genetic Analysis, by K. Lange. Journal of the Royal Statistical Society. Series D (The Statistician), 48(1):146, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2680914.

Moreno:1999:BRB

[1067] Víctor Moreno. Book review: Applied Survival Analysis, by C. T. Le. Journal of the Royal Statistical Society. Series D (The Statistician), 48 (1):146–147, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2680915.

Lumley:1999:BRB

[1068] Thomas Lumley. Book review: Limit Theory for Mixing Dependent Random Variables, by Z. Lin; C. Lu. Journal of the Royal Statistical Society. Series D (The Statistician), 48 (1):147–148, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2680916.

Thas:1999:BRB

[1069] Olivier Thas. Book review: Basics of Structural Equation Modeling, by G. M. Maruyama. Journal of the Royal Statistical Society. Series D (The Statistician), 48(1):148–149, ???? 1999. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2680917.

Downs:1999:BRB

[1070] Matthew Downs and Mary A. Foulkes. Book review: Management of Data in Clinical Trials, by E. McFadden. Journal of the Royal Statistical Society. Series D (The Statistician), 48 (1):149–150, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2680918.

French:1999:BRB

[1071] Simon French. Book review: Advances in Statistical Decision Theory and Applications, by S. Panchapakesan; N. Balakrishnan. Journal of the Royal Statistical Society. Series D (The Statistician), 48(1):150–151, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2680919.

Price:1999:BRB

[1072] James Price. Book review: Functional Data Analysis, by J. O. Ramsay; B. W. Silverman. Journal of the Royal Statistical Society. Series D (The Statistician), 48(1):151–152, ???? 1999. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2680920.

Boshnakov:1999:BRB

[1073] Georgi Boshnakov. Book review: Elements of Multivariate Time Series Analysis, by G. C. Reinsel. Journal of the Royal Statistical Society. Series D (The Statistician), 48(1):152– 153, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2680921.

Ugarte:1999:BRB

[1074] M. Dolores Ugarte. Book review: Multivariate Statistical Inference and Applications, by A. C. Rencher. Journal of the Royal Statistical Society. Series D (The Statistician), 48(1):153– 154, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2680922.

Day:1999:BRBa

[1075] Simon Day. Book review: Statistical Evidence: A Likelihood Paradigm, by R. Royall. Journal of the Royal Statistical Society. Series D (The Statistician), 48(1):154–155, ???? 1999. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2680923.

Day:1999:BRBb

[1076] Simon Day. Book review: Odds Ratios in the Analysis of Contingency Tables, by T. Rudas. Journal of the Royal Statistical Society. Series D (The Statistician), 48(1):155–156, ???? 1999. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2680924.

Longford:1999:BRB

[1077] Nick Longford. Book review: Theory of Sample Surveys, by M. E. Thompson. Journal of the Royal Statistical Society. Series D (The Statistician), 48(1):156–157, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2680925.

Anonymous:1999:CTT

[1078] Anonymous. Corrigendum: A toolkit for testing for non-normality in complete and censored samples. Journal of the Royal Statistical Society. Series D (The Statistician), 48(1):158, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2680926. See [308].

Anonymous:1999:BMa

[1079] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 48(1):??, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2680927.

Montgomery:1999:EDP

[1080] Douglas C. Montgomery. Experimental design for product and process design and development. Journal of the Royal Statistical Society. Series D (The Statistician), 48(2):159–177, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2681183.

Johnson:1999:NSS

[1081] Norman L. Johnson and Samuel Kotz. Non-smooth sailing or triangular distributions revisited after some 50 years. Journal of the Royal Statistical Society. Series D (The Statistician), 48 (2):179–187, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2681184.

Stovring:1999:RHM

[1082] H. Støvring. On radicke and his method for testing mean differences. Journal of the Royal Statistical Society. Series D (The Statistician), 48 (2):189–201, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2681185.

Silver:1999:MVP

[1083] Mick Silver and Sharifah Sakinah Aidid. Modelling the volatility of purchase behaviour. Journal of the Royal Statistical Society. Series D (The Statistician), 48(2):203–213, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2681186.

Leung:1999:DFR

[1084] C. S. Leung, M. S. Patel, and C. A.

McGilchrist. A distribution-free regional cumulative sum for identifying hyperendemic periods of disease incidence. Journal of the Royal Statistical Society. Series D (The Statistician), 48(2):215–225, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2681187.

Blest:1999:COE

[1085] David C. Blest. Choice and order: An extension to Kendall’s τ. Journal of the Royal Statistical Society. Series D (The Statistician), 48 (2):227–237, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2681188.

Magnus:1999:ENB

[1086] Jan R. Magnus and Franc J. G. M. Klaassen. The effect of new balls in tennis: Four years at Wimbledon. Journal of the Royal Statistical Society. Series D (The Statistician), 48 (2):239–246, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2681189.

Magnus:1999:ASF

[1087] Jan R. Magnus and Franc J. G. M. Klaassen. On the advantage of serving first in a tennis set: Four years at Wimbledon. Journal of the Royal Statistical Society. Series D (The Statistician), 48(2):247–256, ???? 1999. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2681190.

Nelder:1999:SSS

[1088] John A. Nelder. From statistics to statistical science. Journal of the Royal Statistical Society. Series D (The Statistician), 48(2):257–269, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2681191.

Bollobas:1999:PE

[1089] Bela Bollobas. Paul Erdős, 1913–1996. Journal of the Royal Statistical Society. Series D (The Statistician), 48 (2):271–273, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2681192.

Bartlett:1999:JEM

[1090] M. S. Bartlett. Jose Enrique Moyal, 1910–1998. Journal of the Royal Statistical Society. Series D (The Statistician), 48(2):273–274, ???? 1999. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2681193.

Bartholomew:1999:AS

[1091] D. J. Bartholomew and M. Knott.

Alan Stuart, 1922–1998. Journal of the Royal Statistical Society. Series D (The Statistician), 48(2):274–276, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2681194.

Anonymous:1999:FMb

[1092] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 48(2):277, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2681182.

Omran:1999:BRB

[1093] M. F. Omran. Book review: Econometrics, by B. H. Baltagi. Journal of the Royal Statistical Society. Series D (The Statistician), 48(2):278, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2681195.

Turlach:1999:BRBa

[1094] B. A. Turlach. Book review: Morphometric Tools for Landmark Data: Geometry and Biology, by F. L. Bookstein. Journal of the Royal Statistical Society. Series D (The Statistician), 48(2):279, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2681196.

Castagliola:1999:BRB

[1095] Philippe Castagliola. Book review: Statistical Control, by Monitoring and Feedback Adjustment, by G. Box; A. Luceño. Journal of the Royal Statistical Society. Series D (The Statistician), 48(2):279–280, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2681197.

Al-Bayatti:1999:BRB

[1096] M. F. Al-Bayatti. Book review: The Practice of Data Analysis: Essays in Honor of John W. Tukey, by D. R. Brillinger; L. T. Fernholz; S. Morgenthaler. Journal of the Royal Statistical Society. Series D (The Statistician), 48(2):280, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2681198.

Militino:1999:BRBb

[1097] Ana F. Militino. Book review: Introduction to Design and Analysis of Experiments, by G. W. Cobb. Journal of the Royal Statistical Society. Series D (The Statistician), 48(2):280– 281, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2681199.

Harkness:1999:BRB

[1098] Miles Harkness. Book review: Statistical Shape Analysis, by I. L. Dryden; K. V. Mardia. Journal of the Royal Statistical Society. Series D (The Statistician), 48(2):281–282, ???? 1999. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2681200.

Machado:1999:BRBa

[1099] Orlando Machado. Book review: The Cambridge Dictionary of Statistics, by B. S. Everitt. Journal of the Royal Statistical Society. Series D (The Statistician), 48(2):282, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2681201.

Gani:1999:BRBb

[1100] Joe Gani. Book review: Retrial Queues, by G. I. Falin; J. G. C. Templeton. Journal of the Royal Statistical Society. Series D (The Statistician), 48(2):282–283, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2681202.

Scott:1999:BRBa

[1101] E. Marian Scott. Book review: Statistics in Ecology and Environmental Monitoring 2, by D. J. Fletcher; L. Kavalieris; B. F. J. Manly. Journal of the Royal Statistical Society. Series D (The Statistician), 48(2):283– 284, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2681203.

Wright:1999:BRBa

[1102] L. F. Wright. Book review: Chaos, Scattering and Statistical Mechanics, by P. Gaspard. Journal of the Royal Statistical Society. Series D (The Statistician), 48(2):284, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2681204.

Kumar:1999:BRBa

[1103] Kuldeep Kumar. Book review: Data Science, Classification and Related Methods, by C. Hayashi; N. Ottsumi; K. Yajima; Y. Tanaka; H. H. Bock; Y. Baba. Journal of the Royal Statistical Society. Series D (The Statistician), 48(2):284–285, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2681205.

Haigh:1999:BRB

[1104] John Haigh. Book review: How to Win More: Strategies for Increasing a Lottery Win, by N. Henze; H. Riedwyl. Journal of the Royal Statistical Society. Series D (The Statistician), 48(2):285–286, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2681206.

Borja:1999:BRBa

[1105] Mario Cortina Borja. Book review: Robust Nonparametric Statistical Methods, by T. P. Hettmansperger; J. W. McKean. Journal of the Royal Statistical Society. Series D (The Statistician), 48(2):286–287, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2681207.

Gani:1999:BRBc

[1106] Joe Gani. Book review: Collected Works of Jaroslav Hajek-with Commentary, by Jaroslav Hajek; M. Huskova; R. Beran; V. Dupac. Journal of the Royal Statistical Society. Series D (The Statistician), 48(2):287– 288, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2681208.

Harris:1999:BRB

[1107] Ray Harris. Book review: Statistical Reasoning and Methods, by R. A. Johnson; K.-W. Tsui. Journal of the Royal Statistical Society. Series D (The Statistician), 48(2):288, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2681209.

Metcalfe:1999:BRB

[1108] Andrew Metcalfe. Book review: A Friendly Guide to Wavelets, by G. Kaiser. Journal of the Royal Statistical Society. Series D (The Statistician), 48(2):288–289, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2681210.

Borja:1999:BRBb

[1109] Mario Cortina Borja. Book review:

S+SpatialStats User’s Manual, by S. P. Kaluzny; S. C. Vega; T. P. Cardoso; A. A. Shelly. Journal of the Royal Statistical Society. Series D (The Statistician), 48(2):289–290, ???? 1999. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2681211.

Reed:1999:BRB

[1110] R. J. Reed. Book review: Loss Models: From Data to Decisions, by S. A. Klugman; H. H. Panjer; G. E. Willmot. Journal of the Royal Statistical Society. Series D (The Statistician), 48(2):290–291, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2681212.

Gani:1999:BRBd

[1111] Joe Gani. Book review: Encyclopedia of Statistical Sciences Update Volume 2, by S. Kotz; C. B. Read; D. L. Banks. Journal of the Royal Statistical Society. Series D (The Statistician), 48(2):291–292, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2681213.

Bell:1999:BRB

[1112] John F. Bell. Book review: Exploring Textual Data, by L. Lebart; A. Salem; L. Berry. Journal of the Royal Statistical Society. Series D (The Statistician), 48(2):292–293, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2681214.

Scott:1999:BRBb

[1113] E. Marian Scott. Book review: Statistical Concepts: A Second Course for Education and the Behavioural Sciences, by R. G. Lomax. Journal of the Royal Statistical Society. Series D (The Statistician), 48(2):293– 294, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2681215.

Sofroniou:1999:BRBb

[1114] Nick Sofroniou. Book review: Easy Use and Interpretation of SPSS for Windows: Answering Research Questions with Statistics, by G. A. Morgan; O. V. Griego. Journal of the Royal Statistical Society. Series D (The Statistician), 48(2):294, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2681216.

Puig:1999:BRB

[1115] Pedro Puig. Book review: Principles of Statistical Inference from a Neo-Fisherian Perspective, by L. Pace; A. Salvan. Journal of the Royal Statistical Society. Series D (The Statistician), 48(2):294–296, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2681217.

Wright:1999:BRBb

[1116] L. F. Wright. Book review: Foundations of Queueing Theory, by N. U. Prabhu. Journal of the Royal Statistical Society. Series D (The Statistician), 48(2):296, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2681218.

Machado:1999:BRBb

[1117] Orlando Machado. Book review: Modern Simulation and Modeling, by R. Y. Rubenstein; B. Melamed. Journal of the Royal Statistical Society. Series D (The Statistician), 48(2):296– 297, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2681219.

Anonymous:1999:BMb

[1118] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 48(2):??, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2681220.

Bendell:1999:FRS

[1119] A. Bendell, J. Disney, and C. Mc-Collin. The future role of statistics in quality engineering and management. Journal of the Royal Statistical Society. Series D (The Statistician), 48 (3):299–326, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2680994.

Nicholson:1999:ISS

[1120] Mike Nicholson and Jon Barry. Inferences from spatial surveys about the distribution of patch size of an unobserved species. Journal of the Royal Statistical Society. Series D (The Statistician), 48(3):327–337, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2680995.

Basu:1999:MAB

[1121] Srabashi Basu, Ayanendranath Basu, and Aparna Raychaudhuri. Measuring agreement between two raters for ordinal response: A model-based approach. Journal of the Royal Statistical Society. Series D (The Statistician), 48(3):339–348, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2680996.

Avery:1999:ADA

[1122] Peter J. Avery and Gillian A. Masters. Advice on the design and analysis of sensory evaluation experiments using a case-study with cooked pork. Journal of the Royal Statistical Society. Series D (The Statistician), 48 (3):349–359, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2680997.

Omran:1999:TCS

[1123] M. F. Omran and E. McKenzie. Testing for covariance stationarity in the UK all-equity returns. Journal of the Royal Statistical Society. Series D (The Statistician), 48(3):361–369, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2680998.

Tofallis:1999:MBM

[1124] Chris Tofallis. Model building with multiple dependent variables and constraints. Journal of the Royal Statistical Society. Series D (The Statistician), 48(3):371–378, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2680999.

Canty:1999:RBV

[1125] Angelo J. Canty and A. C. Davison. Resampling-based variance estimation for labour force surveys. Journal of the Royal Statistical Society. Series D (The Statistician), 48(3):379– 391, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2681000.

Horgan:1999:TSB

[1126] G. W. Horgan, D. A. Elston, M. F. Franklin, C. A. Glasbey, E. A. Hunter, M. Talbot, R. A. Kempton, J. W. McNicol, and F. Wright. Teaching statistics to biological research scientists. Journal of the Royal Statistical Society. Series D (The Statistician), 48(3):393–400, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2681001.

Lindsey:1999:RAS

[1127] J. K. Lindsey. Relationships among sample size, model selection and likelihood regions, and scientifically important differences. Journal of the Royal Statistical Society. Series D (The Statistician), 48(3):401–411, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2681002.

Reiser:1999:CIO

[1128] Benjamin Reiser and David Faraggi. Confidence intervals for the overlapping coefficient: The normal equal variance case. Journal of the Royal Statistical Society. Series D (The Statistician), 48(3):413–418, ???? 1999. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2681003.

LaMotte:1999:PIC

[1129] Lynn Roy LaMotte and Júlia Volaufová. Prediction intervals via consonance intervals. Journal of the Royal Statistical Society. Series D (The Statistician), 48(3):419–424, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2681004.

Haigh:1999:PIB

[1130] John Haigh. (performance) index betting and fixed odds. Journal of the Royal Statistical Society. Series D (The Statistician), 48(3):425–434, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2681005.

Henery:1999:MRP

[1131] R. J. Henery. Measures of overround in performance index betting. Journal of the Royal Statistical Society. Series D (The Statistician), 48 (3):435–439, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2681006.

Ryan:1999:C

[1132] Tom Ryan, Peter P. Howley, G. Janacek, and S. Meikle. Correspondence. Journal of the Royal Statistical Society. Series D (The Statistician), 48(3):441– 444, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2681007.

Anonymous:1999:FMc

[1133] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 48(3):445, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2680993.

Scott:1999:BRBc

[1134] Marian Scott. Book review: Basic Linear Geostatistics, by M. Armstrong. Journal of the Royal Statistical Society. Series D (The Statistician), 48(3):446, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2681008.

Gerontidis:1999:BRB

[1135] Ioannis I. Gerontidis. Book review: Nonnegative Matrices and Applications, by R. B. Bapat; T. E. S. Raghavan. Journal of the Royal Statistical Society. Series D (The Statistician), 48(3):447–448, ???? 1999. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2681009.

Gani:1999:BRBe

[1136] Joe Gani. Book review: Nonlinear System Techniques and Applications, by J. S. Bendat. Journal of the Royal Statistical Society. Series D (The Statistician), 48(3):448–449, ???? 1999. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2681010.

Gani:1999:BRBf

[1137] Joe Gani. Book review: Efficient and Adaptive Estimation for Semiparametric Models, by P. J. Bickel; C. A. J. Klaassen; Y. Ritov; J. A. Wellner. Journal of the Royal Statistical Society. Series D (The Statistician), 48(3):449–450, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2681011.

Stoyanov:1999:BRB

[1138] Jordan Stoyanov. Book review: Ergodicity and Stability of Stochastic Processes, by A. A. Borovkov. Journal of the Royal Statistical Society. Series D (The Statistician), 48(3):450– 451, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2681012.

Turlach:1999:BRBb

[1139] B. A. Turlach. Book review: Programming with Data: A Guide to the S Language, by J. M. Chambers. Journal of the Royal Statistical Society. Series D (The Statistician), 48 (3):451–452, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2681013.

Rohrl:1999:BRB

[1140] Armin Röhrl. Book review: The Mathematica Primer, by K. R. Coombes; B. R. Hunt; R. L. Lipsman; J. E. Osborn; G. K. Stuck. Journal of the Royal Statistical Society. Series D (The Statistician), 48(3):452–453, ???? 1999. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2681014.

Allin:1999:BRB

[1141] Paul Allin. Book review: Statistics in Society: The Arithmetic of Politics, by D. Dorling; S. Simpson. Journal of the Royal Statistical Society. Series D (The Statistician), 48(3):454, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2681015.

Thompson:1999:BRB

[1142] Robin Thompson. Book review: A First Course in Multivariate Statistics, by B. Flury. Journal of the Royal Statistical Society. Series D (The Statistician), 48(3):455, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2681016.

Wright:1999:BRBc

[1143] L. F. Wright. Book review: Cooperation: Game-Theoretic Approaches, by S. Hart; A. Mas-Colell. Journal of the Royal Statistical Society. Series D (The Statistician), 48(3):455– 456, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2681017.

Hinde:1999:BRB

[1144] John Hinde. Book review: Quasi-Likelihood and Its Application: A General Approach to Optimal Parameter Estimation, by C. C. Heyde. Journal of the Royal Statistical Society. Series D (The Statistician), 48(3):456, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2681018.

Coolen:1999:BRB

[1145] Frank Coolen. Book review: Statistical Methods for Reliability Data, by W. Q. Meeker; L. A. Escobar. Journal of the Royal Statistical Society. Series D (The Statistician), 48 (3):456–458, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2681019.

Hemmett:1999:BRB

[1146] Lindsay Hemmett. Book review: Statistical Power Analysis, by K. R. Murphy; B. Myors. Journal of the Royal Statistical Society. Series D (The Statistician), 48(3):458, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2681020.

Wright:1999:BRBd

[1147] L. F. Wright. Book review: Game Theoretical Applications to Economics and Operations Research, by T. Parthasarathy. Journal of the Royal Statistical Society. Series D (The Statistician), 48(3):458–459, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2681021.

Sapatinas:1999:BRB

[1148] Theofanis Sapatinas. Book review:

Matrix Algebra and Its Applications to Statistics and Econometrics, by C. R. Rao; M. B. Rao. Journal of the Royal Statistical Society. Series D (The Statistician), 48(3):459–460, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2681022.

Kimber:1999:BRBb

[1149] Alan Kimber. Book review: Applied Regression Analysis — A Research Tool, by J. O. Rawlings; S. G. Pantula; D. A. Dickey. Journal of the Royal Statistical Society. Series D (The Statistician), 48(3):460–461, ???? 1999. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2681023.

Denison:1999:BRB

[1150] David G. T. Denison. Book review: Introduction to Probability and Statistics for Scientists and Engineers, by W. A. Rosenkrantz. Journal of the Royal Statistical Society. Series D (The Statistician), 48(3):461, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2681024.

Myklebust:1999:BRB

[1151] Thor Myklebust. Book review: Into Statistics, by P. J. Smith. Journal of the Royal Statistical Society. Series D (The Statistician), 48(3):461– 462, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2681025.

Kumar:1999:BRBb

[1152] Kuldeep Kumar. Book review: Image Processing and Data Analysis, by J.-L. Starck; F. Murtagh; A. Bijaoui. Journal of the Royal Statistical Society. Series D (The Statistician), 48(3):462, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2681026.

Anonymous:1999:BMc

[1153] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 48(3):??, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2681027.

Curnow:1999:UNG

[1154] Robert N. Curnow. Unfathomable nature and government policy. Journal of the Royal Statistical Society. Series D (The Statistician), 48(4):463– 476, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2681124.

Hood:1999:AIV

[1155] Kerenza Hood, Barry A. J. Nix, and Terence C. Iles. Asymptotic information and variance-covariance matrices for the linear structural model. Journal of the Royal Statistical Society. Series D (The Statistician), 48 (4):477–493, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2681125.

Lipsitz:1999:DFA

[1156] Stuart R. Lipsitz, Joseph G. Ibrahim, and Michael Parzen. A degrees-of-freedom approximation for a t-statistic with heterogeneous variance. Journal of the Royal Statistical Society. Series D (The Statistician), 48(4):495– 506, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2681126.

vandeWiel:1999:SCE

[1157] M. A. van de Wiel, A. Di Bucchianico, and P. van der Laan. Symbolic computation and exact distributions of nonparametric test statistics. Journal of the Royal Statistical Society. Series D (The Statistician), 48 (4):507–516, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2681127.

Giraudeau:1999:TDT

[1158] Bruno Giraudeau and Claude Chastang. Two discordancy tests for location slippage and dispersion slippage outlier detection in agreement studies. Journal of the Royal Statistical Society. Series D (The Statistician), 48(4):517–527, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2681128.

Paula:1999:LIC

[1159] Gilberto A. Paula. Leverage in inequality-constrained regression models. Journal of the Royal Statistical Society. Series D (The Statistician), 48(4):529–538, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2681129.

Anonymous:1999:RCS

[1160] Anonymous. Report of the Council for the session 1998–1999. Journal of the Royal Statistical Society. Series D (The Statistician), 48(4):539– 603, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2681130.

Plackett:1999:JPB

[1161] Robin Plackett. J. P. Burman, 1924– 1998. Journal of the Royal Statistical Society. Series D (The Statistician), 48(4):605–606, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2681131.

Anonymous:1999:FMd

[1162] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 48(4):607, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2681123.

Kemp:1999:BRBa

[1163] David Kemp. Book review: Records, by B. C. Arnold; N. Balakrishnan; H. N. Nagaraja. Journal of the Royal Statistical Society. Series D (The Statistician), 48(4):608, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2681132.

Bartkowiak:1999:BRB

[1164] Anna Bartkowiak. Book review: Statistics for the Environment 4: Statistical Aspects of Health and the Environment, by V. Barnett; A. Stein; K. F. Turkman. Journal of the Royal Statistical Society. Series D (The Statistician), 48(4):608–610, ???? 1999. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2681133.

Kemp:1999:BRBb

[1165] Adrienne (Freda) Kemp. Book review: Probability and Its Applications, by M. C. Bhattacharjee; S. K. Basu. Journal of the Royal Statistical Society. Series D (The Statistician), 48 (4):610–611, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2681134.

Scott:1999:BRBd

[1166] Marian Scott. Book review: Statistical Regression with Measurement Error, by C.-L. Cheng; J. W. Van Ness. Journal of the Royal Statistical Society. Series D (The Statistician), 48 (4):611–612, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2681135.

Woodward:1999:BRB

[1167] P. W. Woodward. Book review: Regression Graphics: Ideas for Studying Regression through Graphics, by R. D. Cook. Journal of the Royal Statistical Society. Series D (The Statistician), 48(4):612–613, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2681136.

Foulkes:1999:BRB

[1168] Mary A. Foulkes. Book review: Design and Analysis of Experiments, by A. M. Dean; D. T. Voss. Journal of the Royal Statistical Society. Series D (The Statistician), 48(4):613– 614, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2681137.

Sprent:1999:BRBb

[1169] P. Sprent. Book review: Fitting Linear Relationships — A History of the Calculus of Observations, 1750–1900, by R. W. Farebrother. Journal of the Royal Statistical Society. Series D (The Statistician), 48(4):614–615, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2681138.

Lopatatzidis:1999:BRB

[1170] Thanasis Lopatatzidis. Book review:

Policy Simulations in the European Union, by A. Fossati; J. Hutton. Journal of the Royal Statistical Society. Series D (The Statistician), 48 (4):615–616, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2681139.

Wright:1999:BRBe

[1171] L. F. Wright. Book review: Data Analysis Using SPSS for Windows, by J. J. Foster. Journal of the Royal Statistical Society. Series D (The Statistician), 48(4):616, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2681140.

Fouskakis:1999:BRB

[1172] Dimitris Fouskakis and David Draper.

Book review: Tabu Search, by F. Glover; M. Laguna. Journal of the Royal Statistical Society. Series D (The Statistician), 48(4):616–619, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2681141.

Rigby:1999:BRB

[1173] Alan S. Rigby. Book review: Nonresponse in Household Interview Surveys, by R. M. Groves; M. P. Couper. Journal of the Royal Statistical Society. Series D (The Statistician), 48(4): 619, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2681142.

Newell:1999:BRB

[1174] John Newell. Book review: Applied Survival Analysis: Regression Modelling of Time to Event Data, by D. W. Hosmer; S. Lemeshow. Journal of the Royal Statistical Society. Series D (The Statistician), 48(4):620– 621, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2681143.

Mizera:1999:BRB

[1175] Ivan Mizera. Book review: Theory of Point Estimation, by E. L. Lehmann; G. Casella. Journal of the Royal Statistical Society. Series D (The Statistician), 48(4):621–622, ???? 1999. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2681144.

Roddam:1999:BRB

[1176] Andrew W. Roddam. Book review: Statistics Explained, by J. P. Lewis; A. Traill. Journal of the Royal Statistical Society. Series D (The Statistician), 48(4):622, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2681145.

Espejo:1999:BRB

[1177] Mariano Ruiz Espejo. Book review:

Decoupling: From Dependence to Independence, by V. H. de la Peña; E. Giné. Journal of the Royal Statistical Society. Series D (The Statistician), 48(4):622–623, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2681146.

Kemp:1999:BRBc

[1178] Adrienne (Freda) Kemp. Book review: Probability Theory III, by Yu. V. Prokhorov; A. N. Shiryaev. Journal of the Royal Statistical Society. Series D (The Statistician), 48(4):623– 624, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2681147.

Militino:1999:BRBc

[1179] Ana F. Militino. Book review: Multivariate Geostatistics: An Introduction with Applications, by H. Wackernagel. Journal of the Royal Statistical Society. Series D (The Statistician), 48(4):624–625, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2681148.

Everitt:1999:BRB

[1180] B. S. Everitt. Book review: Multivariate Taxometric Procedures: Distinguishing Types from Continua, by N. G. Waller; P. E. Meel. Journal of the Royal Statistical Society. Series D (The Statistician), 48(4):625, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2681149.

Ho:1999:BRB

[1181] Tsung wu Ho. Book review: Advances in Econometric Theory — The Selected Works of Halbert White, by H. White. Journal of the Royal Statistical Society. Series D (The Statistician), 48(4):625–626, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2681150.

Anonymous:1999:VI

[1182] Anonymous. Volume information. Journal of the Royal Statistical Society. Series D (The Statistician), 48 (4):627–634, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/2681122.

Anonymous:1999:BMd

[1183] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 48(4):??, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2681151.